### v3

# Testing Conditional Independence via the Spectral Generalized Covariance Measure: Beyond Euclidean Data

Abstract
We propose a conditional independence (CI) test based on a new measure, the spectral generalized covariance measure (SGCM). The SGCM is constructed by approximating the basis expansion of the squared norm of the conditional cross-covariance operator, using data-dependent bases obtained via spectral decompositions of empirical covariance operators. This construction avoids direct estimation of conditional mean embeddings and reduces the problem to scalar-valued regressions, resulting in robust finite-sample size control. Theoretically, we derive the limiting distribution of the SGCM statistic, establish the validity of a wild bootstrap for inference, and obtain uniform asymptotic size control under doubly robust conditions. As an additional contribution, we show that exponential kernels induced by continuous semimetrics of negative type are characteristic on general Polish spaces—with extensions to finite tensor products—thereby providing a foundation for applying our test and other kernel methods to complex objects such as distribution-valued data and curves on metric spaces. Extensive simulations indicate that the SGCM-based CI test attains near-nominal size and exhibits power competitive with or superior to state-of-the-art alternatives across a range of challenging scenarios.
Spectral decomposition,
keywords:
[…]
and
1 Introduction
Conditional independence (CI) is a cornerstone of statistics and machine learning, shaping modern approaches to inference and causality. In causal inference, Pearl2009Overview shows how CI clarifies causal relationships and supports the construction of causal models, while Dawid1979 demonstrates that many core ideas in statistical inference can be characterized in terms of CI. A prototypical application of CI is variable selection. Building on Candesetal2018, it is known that, in parametric generalized linear models, CI corresponds to the vanishing of regression coefficients, making CI tests a genuinely model-free route to variable selection within that framework. While foundational, these classical approaches implicitly assume that observations can be represented as vectors.
Modern data pipelines routinely generate complex observations that are naturally represented as (time-varying) random objects—elements of general metric spaces, potentially observed over time as evolving trajectories—rather than Euclidean vectors (PetersenandMuller2019; DubeyandMuller2020). Applications range from the macroscopic to the molecular: hurricane tracks are modeled as random paths on a sphere (Kendall2015), and protein backbones as sequences of points on a torus in probabilistic structure prediction (Boomsmaetal2008). This random-object framework also naturally accommodates representations such as probability distributions and compositional data. For example, entire income distributions can be compared directly, and cause-specific mortality shares can be analyzed as time-varying compositions to reveal evolving public-health trends (BritoandDias2022; StefanucciandMazzuco2022).
Motivated by these applications, we introduce a nonparametric methodology for testing
[…]
based on i.i.d. samples $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{n}$ drawn from $P_{XYZ}$ on a product of Polish spaces. Here, “nonparametric CI testing” means that validity and power do not rely on any finite-dimensional (semi)parametric specification of $P_{X\mid Z}$ or $P_{Y\mid Z}$, nor on strong structural or distributional restrictions such as Gaussianity (KalischandBuhlmann2007), isotonicity (Horeetal2025), or the Model-X requirement of a correctly specified $P_{X\mid Z}$ (Berrettetal2020; Niuetal2024). Our framework advances kernel-based methodology with a spectral, regression-friendly construction that ensures rigorous size control and high power, even in high-dimensional and non-Euclidean settings.
1.1 Related Work
Before presenting our contributions, we review the related literature. Following the comprehensive taxonomy of Lietal2020survey_CI, recent nonparametric CI tests can be broadly grouped into four categories:
  1. 1.
    Regression-based methods exploit residuals from nonparametric regressions of $X$ on $Z$ and $Y$ on $Z$ to construct test statistics (e.g., Shah and Peters, 2020; Scheidegger, Hörrmann and Bühlmann, 2022; Lundborg, Shah and Peters, 2022; Shi et al., 2021; Fernández and Rivera, 2024).
  2. 2.
    Kernel-based methods rely on kernel mean embeddings (KMEs) to characterize distributional quantities (e.g., Zhang2011; Scetbon2022; Pogodin et al., 2025; Zhangetal2025).
  3. 3.
    Metric-based methods employ distances between distributions or associated variants (e.g., Wangetal2015; Caietal2022; Banerjeeetal2024).
  4. 4.
    Other methods include permutation-based tests (e.g., Doranetal2014; Kimetal2022), a classification-based test (Senetal2017) and tests driven by generative models (e.g., BellotandVanDerSchaar2019; DuongandNguyen2024; Yangetal2024).
We compare these categories along four axes: (i) uniform asymptotic size control, (ii) power against $H_{1}$, (iii) robustness to high-dimensional data, and (iv) applicability to non-Euclidean data. Our high-level comparison across these axes is summarized in Table 1; detailed discussion follows in Section 1.1.11.1.4. Our comparisons focus on tests with stable null calibration; generative model approaches, which often require substantial data and tuning, are therefore not included here.
Table 1: Qualitative comparison of CI tests across four groups (✓= favorable, - = limited).
Method Uniform Size Power High-Dimensional Non-Euclid
1. Regression-based
Shah and Peters (2020) - -
Shi et al. (2021) - - -
Lundborg, Shah and Peters (2022) -
Scheidegger, Hörrmann and Bühlmann (2022) - -
Fernández and Rivera (2024) - - -
2. Kernel-based
Zhang2011 - -
Scetbon2022 - - -
Pogodin et al. (2025) - -
Zhangetal2025 - -
3. Metric-based
Wangetal2015 - - -
Caietal2022 - - -
Banerjeeetal2024 - - -
4. Others
Senetal2017 - -
Kimetal2022 - -
1.1.1 Uniform Asymptotic Size Control
It is desirable to have uniform size control over the entire set of null distributions $\mathcal{P}_{0}$ under which CI holds; see Lundborg2022 for a detailed discussion. For a sequence of test functions $\{\phi_{n}\}_{n=1}^{\infty}$, the uniform and pointwise asymptotic levels are defined as
$\limsup_{n\to\infty}\sup_{P\in\mathcal{P}_{0}}\mathbb{P}_{P}(\phi_{n}=1)~~~\text{and}~~~\sup_{P\in\mathcal{P}_{0}}\limsup_{n\to\infty}\mathbb{P}_{P}(\phi_{n}=1),$
respectively. Control of the former is stricter than that of the latter. In the context of CI testing, although pointwise asymptotic size control is commonly considered, achieving uniform asymptotic size control poses a significant challenge. In fact, a body of recent work, such as Shah and Peters (2020); Neykovetal2021; Kimetal2022, has established a fundamental impossibility result: any CI test that satisfies the uniform level criterion necessarily has “trivial” power; namely, this is formalized as
$\sup_{Q\in\mathcal{P}_{1}}\limsup_{n\to\infty}\mathbb{P}_{Q}(\phi_{n}=1)\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}_{0}}\mathbb{P}_{P}(\phi_{n}=1),$
where $\mathcal{P}_{1}$ is the complementary set of $\mathcal{P}_{0}$, representing the alternative hypothesis. To bypass this “no-free-lunch” barrier, Kimetal2022 impose restrictive smoothness conditions on the conditional distributions. Another prominent approach, particularly for regression-based methods, makes use of more tractable surrogate dependence measures for CI, such as $\operatorname{\mathbb{E}}[\operatorname{Cov}(X,Y\mid Z)]$, which vanish under CI yet are not equivalent to it (e.g., Shah and Peters, 2020; Scheidegger, Hörrmann and Bühlmann, 2022; Lundborg, Shah and Peters, 2022). This choice allows uniform asymptotic size to be established under relatively mild assumptions, but entails that rigorous power guarantees can be stated only for a restricted subclass of alternatives for which vanishing of the surrogate coincides with genuine CI. In effect, the existing CI tests that offer uniform asymptotic size either trade power guarantees for restrictions on the alternative class via tractable surrogate measures, or else rely on local permutation schemes that do not scale beyond low-dimensional settings.
1.1.2 Power Against $H_{1}$
Many tests, most of which are regression-based, restrict the class of alternatives for which they can guarantee asymptotic power (e.g., Shah and Peters, 2020; Shi et al., 2021; Scheidegger, Hörrmann and Bühlmann, 2022; Lundborg, Shah and Peters, 2022; Fernández and Rivera, 2024); we discuss the resulting power properties in detail in the Supplementary Material (Appendix F). Other methods often achieve asymptotic power by employing measures that are equivalent to CI.
1.1.3 Robustness to High-Dimensional Data
Regression-based tests enjoy double robustness, which renders them particularly effective in high-dimensional settings. They can achieve $\sqrt{n}$-consistency and normal approximation as long as the nuisance errors converge sufficiently fast. By contrast, while kernel methods have demonstrated a certain degree of empirical robustness in high dimensions (e.g., Lietal2020survey_CI; Pogodin et al., 2025), their theoretical guarantees remain weak, often requiring stringent assumptions on the bias and convergence rates of Conditional Mean Embeddings (CMEs) (Pogodin et al., 2025). Crucially, the estimators utilized in Zhang2011; Scetbon2022; Pogodin et al. (2025) are based on kernel ridge regression; however, this approach encounters fundamental limitations in high-dimensional settings, as the associated regression bias becomes non-negligible. On the other hand, Zhangetal2025 estimate the CME via a generative neural network, aiming to mitigate finite-sample bias and to yield an estimator that is more resilient in high-dimensional regimes; they report improved performance under high-dimensional settings. Metric-based approaches, which rely on explicit estimation of conditional laws (e.g., conditional characteristic functions or densities), typically suffer from severe dimensionality problems (Wangetal2015; Caietal2022; Banerjeeetal2024). Other alternatives include the classification-based strategy of Senetal2017, who implement their method with XGBoost (ChenandGuestrin2016), and the local-permutation method of Kimetal2022, for which upper bounds on the test size grow with the dimension of $Z$, so the procedure is vulnerable to the curse of dimensionality.
1.1.4 Applicability to Non-Euclidean Data
In the context of CI testing, Lundborg, Shah and Peters (2022) extended the Generalized Covariance Measure of Shah and Peters (2020) to separable Hilbert-space-valued data. Extensions of distance covariance and Ball divergence (or Ball covariance)—methods widely used for testing (unconditional) independence in general metric spaces—have also been proposed for CI testing (Wangetal2015; Banerjeeetal2024). However, in the CI setting these approaches require explicit estimation of conditional laws, and for non-Euclidean data their theoretical guarantees are currently limited. Kernel-based CI methods remain underdeveloped. While the results of Scetbon2022 and Zhangetal2025 rely on Euclidean structure, kernel-based CI can, in principle, operate on non-Euclidean data by employing characteristic kernels tailored to the domain (Sriperumbudur, Fukumizu and Lanckriet, 2011), an approach that avoids the need to estimate distributional quantities directly. For example, Wynne and Duncan (2022) construct characteristic kernels for separable Hilbert-space-valued data, and Ziegeletal2024 extend this framework to metric spaces of strongly negative type (Lyons2013). These constructions, however, impose substantially stronger requirements than those needed for positive definiteness alone, and general necessary-and-sufficient criteria remain scarce.
1.2 Our Contributions
We develop a CI testing procedure based on the Spectral Generalized Covariance Measure (SGCM), which seeks to leverage the strengths of both regression-based and kernel-based methods. Specifically, it is grounded in the theoretical framework of the Conditional Cross-Covariance Operator (CCCO), but circumvents the direct and often unstable estimation of conditional mean embeddings (CMEs). Instead, we approximate the squared norm of the CCCO by a finite expansion using suitably chosen data-dependent basis functions. This reduces the direct operator estimation to Euclidean-valued regressions for the expansion coefficients. This construction helps control the approximation bias and ensures doubly robust behavior of the resulting test, leading to robust finite-sample size control. Moreover, and importantly, the SGCM applies seamlessly beyond Euclidean settings, including distribution-valued and time-varying metric-valued data, by exploiting characteristic kernels. Simulation studies confirm strong finite-sample performance, with accurate size control and competitive power across a wide range of scenarios.
Regarding the theory for our SGCM, we derive the asymptotic distribution of the statistic and establish the validity of its bootstrap version, ensuring both pointwise and uniform asymptotic size control under suitable conditions such as double robustness, while retaining power guarantees. As a by-product, in the Supplementary Material, we show that analogous doubly robust guarantees carry over to existing kernel-based methods, in turn sharpening their convergence rates. We further establish general results on characteristic kernels, showing that exponential kernels induced by continuous conditionally negative definite semimetrics on Polish spaces and their tensor products are characteristic. These findings not only justify the extension of kernel methods to non-Euclidean domains but also supply new tools for inference on probability densities and random objects, thereby substantially broadening the scope of kernel-based testing.
Collectively, our results satisfy all four desiderata in Section 1.1, establish double robustness for our test and for existing kernel-based methods, and develop general constructions of characteristic kernels on Polish spaces, enabling RKHS-based inference for non-Euclidean data.
1.3 Organization
The remainder of this paper is organized as follows. Section 2 introduces our CI measure, grounded in KMEs. Section 3 describes its practical implementation as the SGCM, including the test statistic and wild-bootstrap calibration. Section 4 establishes asymptotic properties—consistency, bootstrap validity, and uniform size control. Section 5 develops a general theory for constructing characteristic kernels on non-Euclidean Polish spaces. Section 6 presents extensive finite-sample evidence. All proofs and additional results appear in the Supplementary Material.
2 Definitions, Background, and Our CI Measure
Let $X,Y,Z$ be random variables with joint distribution $P_{XYZ}$ taking values in Polish spaces $\mathcal{X},\mathcal{Y},\mathcal{Z}$ (i.e., separable and completely metrizable topological spaces). When a metric is explicit, e.g., $(\mathcal{M},d_{\mathcal{M}})$, “Polish” is always understood with respect to the topology induced by $d_{\mathcal{M}}$. We equip $\mathcal{X},\mathcal{Y},\mathcal{Z}$ with positive definite kernels $k_{\mathcal{X}},k_{\mathcal{Y}},k_{\mathcal{Z}}$ and corresponding reproducing kernel Hilbert spaces (RKHSs) $\mathcal{H}_{\mathcal{X}},\mathcal{H}_{\mathcal{Y}},\mathcal{H}_{\mathcal{Z}}$. Denote by $\mathcal{H}\otimes\mathcal{H}^{\prime}$ the Hilbert tensor product, obtained by completing the algebraic tensor product under the inner product $\langle f\otimes f^{\prime},\,g\otimes g^{\prime}\rangle_{\mathcal{H}\otimes\mathcal{H}^{\prime}}\coloneqq\langle f,g\rangle_{\mathcal{H}}\,\langle f^{\prime},g^{\prime}\rangle_{\mathcal{H}^{\prime}}$. Unless stated otherwise, $\left\|\cdot\right\|$ and $\left\langle\cdot,\cdot\right\rangle$ denote, respectively, the norm and the inner product of the Hilbert space to which the arguments belong. $\|\cdot\|_{\mathrm{HS}}$, $\|\cdot\|_{\mathrm{tr}}$, and $\|\cdot\|_{\textrm{op}}$ denote the Hilbert–Schmidt, trace, and operator norms, respectively.
Let $\mathcal{S}$ be a Polish space with Borel $\sigma$-algebra $\mathcal{B}(\mathcal{S})$; we denote by $\mathcal{P}(\mathcal{S})$ the set of all Borel probability measures on $\mathcal{S}$. When it is not necessary to specify $\mathcal{S}$, we simply write $\mathcal{P}$. We use standard notations for convergence in probability ($\overset{p}{\to}$), in distribution ($\overset{d}{\to}$), and for probabilistic orders ($o_{p}(1),O_{p}(1)$). The notation $\overset{d^{\ast}}{\to}$ denotes conditional weak convergence of bootstrap laws.
We further introduce uniform versions of these notations over a family of probability measures $\mathcal{P}$. Specifically, $o_{\mathcal{P}}(1)$ denotes a sequence of random variables that converges uniformly in probability to zero over $\mathcal{P}$; that is, $X_{n}=o_{\mathcal{P}}(1)$ means that for every $\varepsilon>0$, we have $\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{n}|>\varepsilon)\to 0$ as $n\to\infty$. Similarly, $O_{\mathcal{P}}(1)$ denotes a sequence that is uniformly bounded in probability over $\mathcal{P}$.
2.1 Conditional Cross-Covariance Operator
To construct our CI test, we first build the core component, CCCO. Although this was already proposed in Pogodin et al. (2025), we reconstruct it here under weaker assumptions—the key ingredient is the CME. Hereafter, we assume that the kernels we work with satisfy the following conditions:
Assumption 2.1.
The kernels $k_{\mathcal{X}},k_{\mathcal{Y}},k_{\mathcal{Z}}$ are bounded; that is, $\sup_{x\in\mathcal{X}}k_{\mathcal{X}}(x,x)<\infty$, $\sup_{y\in\mathcal{Y}}k_{\mathcal{Y}}(y,y)<\infty$, and $\sup_{z\in\mathcal{Z}}k_{\mathcal{Z}}(z,z)<\infty$. Moreover, we assume that the feature maps $\mathcal{X}\ni x\mapsto k_{\mathcal{X}}(x,\cdot)\in\mathcal{H}_{\mathcal{X}}$, $\mathcal{Y}\ni y\mapsto k_{\mathcal{Y}}(y,\cdot)\in\mathcal{H}_{\mathcal{Y}}$, and $\mathcal{Z}\ni z\mapsto k_{\mathcal{Z}}(z,\cdot)\in\mathcal{H}_{\mathcal{Z}}$ are Borel measurable.
Since the input spaces are Polish, the measurability of the feature maps in Assumption 2.1 ensures separability of the associated RKHS (Owhadi and Scovel, 2017, Theorem 2.4). Furthermore, every RKHS function is measurable (Steinwart and Christmann, 2008, Lemma 4.24).
To minimize assumptions, we adopt the measure-theoretic definitions of the KME and CME following Park and Muandet (2020). Under Assumption 2.1, for any probability measure $P$ on the corresponding space, all Bochner integrals below are well defined.
Definition 2.1.
  1. 1.
    For any probability measure $P_{X}$ on $\mathcal{X}$, the KME is
    $\mu_{P_{X}}^{(k_{\mathcal{X}})}\coloneqq\int_{\mathcal{X}}k_{\mathcal{X}}(x,\cdot)P_{X}(\mathop{}\!\mathrm{d}x).$
    When $P_{X}$ and $k_{\mathcal{X}}$ are clear, we simply write $\mu_{X}$.
  2. 2.
    Given a regular conditional distribution $P_{X\mid Z=z}$, the CME is
    $\mu_{X\mid Z}(z)\coloneqq\int_{\mathcal{X}}k_{\mathcal{X}}(x,\cdot)P_{X\mid Z=z}(\mathop{}\!\mathrm{d}x),\qquad z\in\mathcal{Z}.$
  3. 3.
    A bounded kernel $k_{\mathcal{X}}$ is characteristic if and only if the map $\mathcal{P}(\mathcal{X})\to\mathcal{H}_{k_{\mathcal{X}}}$, $P_{X}\mapsto\mu_{P_{X}}^{(k_{\mathcal{X}})}$, is injective.
We impose the following characteristicity condition on the kernels.
Assumption 2.2.
At least one of the following conditions holds: (i) $k_{\mathcal{X}}\otimes k_{\mathcal{Z}}$ and $k_{\mathcal{Y}}$ are characteristic; (ii) $k_{\mathcal{Y}}\otimes k_{\mathcal{Z}}$ and $k_{\mathcal{X}}$ are characteristic; (iii) $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ are characteristic while $k_{\mathcal{Z}}$ is $L^{2}$-universal.
On general Polish spaces, tensor products of characteristic exponential-type kernels remain characteristic (Corollary 5.3). If one restricts attention to Euclidean spaces, more general conditions for when product kernels are characteristic are available; see SzaboandSriperumbudur2018. Alternatively, rather than imposing the characteristic property on the tensor product kernel, one can assume $L^{2}$-universality (Sriperumbudur, Fukumizu and Lanckriet, 2011).
We define the random element $\Psi\in\mathcal{H}_{XYZ}\coloneqq\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Y}}\otimes\mathcal{H}_{\mathcal{Z}}$ as
$\Psi$ $\coloneqq\left(k_{\mathcal{X}}(X,\cdot)-\mu_{X\mid Z}(Z)\right)\otimes\left(k_{\mathcal{Y}}(Y,\cdot)-\mu_{Y\mid Z}(Z)\right)\otimes k_{\mathcal{Z}}(Z,\cdot)$
$\eqqcolon\epsilon_{X\mid Z}\otimes\epsilon_{Y\mid Z}\otimes k_{\mathcal{Z}}(Z,\cdot).$ (1)
Following Pogodin et al. (2025), we then define the CCCO as its expectation:
$\Sigma_{XYZ\mid Z}\coloneqq\operatorname{\mathbb{E}}[\Psi].$ (2)
We now establish the connection between $\Sigma_{XYZ\mid Z}$ and CI.
Proposition 2.1.
For any Borel probability measure $P_{XYZ}\in\mathcal{P}(\mathcal{X}\times\mathcal{Y}\times\mathcal{Z})$, under Assumptions 2.1 and 2.2, $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$ holds if and only if $\Sigma_{XYZ\mid Z}=0$.
For the above equivalence, Assumption 2.2 is weaker than the $L^{2}$-universality assumed in Pogodin et al. (2025). The CCCO defined via CMEs in (1) coincides, under a suitable identification, with the classical CCCO introduced by Fukumizu, Bach and Jordan (2004), which was formulated without CMEs as an RKHS analog of the conditional covariance for Gaussian random variables. A formal proof is provided in the Supplementary Material (Proposition A.2).
Our CME-based construction offers three advantages: (i) it is less sensitive to high dimensionality because it avoids defining a kernel on the joint variable $\ddot{Y}=(Y,Z)$, instead yielding a separated representation in $\mathcal{H}_{XYZ}$; (ii) it relaxes the assumptions required for the kernels; and (iii) it delivers a doubly robust property established later.
3 Methodology: Spectral Generalized Covariance Measure
Building on the CCCO developed in (2), we introduce a practical CI test that remains applicable to complex data with non-Euclidean structures.
A conventional test statistic based on the CCCO is constructed by first estimating the CMEs $\mu_{X\mid Z}$ and $\mu_{Y\mid Z}$, plugging them into $\Psi$, and then computing the squared norm of the resulting empirical average. However, this plug-in approach has two drawbacks. First, it recasts the problem as an infinite-dimensional vector-valued nonparametric regression, an intrinsically hard setting in which some form of regularization is unavoidable; the resulting bias propagates to the test statistic and can distort the size of the tests. Second, to exploit the reproducing property, the estimator must be RKHS-valued, which rules out generic function regressors and, in practice, confines one to finite linear combinations of the feature maps of the training inputs, $\sum_{i=1}^{n}\alpha_{i}k_{\mathcal{X}}(X_{i},\cdot)$ (Li et al., 2022). Pogodin et al. (2025) and Zhangetal2025 partially address the symptoms, but the former does not entirely resolve the intrinsic difficulty of infinite-dimensional regression, and the latter requires direct estimation of conditional distributions via a generative model, which is inherently challenging.
To bypass these issues, we propose a testing procedure based on the SGCM. This methodology avoids any direct estimation of the CCCO or the underlying CMEs. Instead, the SGCM targets a basis function expansion of $\|\Sigma_{XYZ\mid Z}\|^{2}$, which simplifies the complex infinite dimensional regression problem into a series of standard, Euclidean-valued regressions.
3.1 Test Statistic
Let $\Psi_{i}$, $\epsilon_{X\mid Z}^{(i)}$, and $\epsilon_{Y\mid Z}^{(i)}$ be i.i.d. copies of the random elements defined in (1). To decompose the CCCO squared norm, $\|\Sigma_{XYZ\mid Z}\|^{2}$, using Parseval’s identity, we utilize the separability of the RKHSs under Assumption 2.1. Recall (1). For arbitrary orthonormal bases $\{e_{p}\}_{p=1}^{\infty}\subset\mathcal{H}_{\mathcal{X}}$ and $\{f_{q}\}_{q=1}^{\infty}\subset\mathcal{H}_{\mathcal{Y}}$, the coordinatewise residuals are expressed as
$\zeta_{X\mid Z,p}^{(i)}\coloneqq\langle\epsilon_{X\mid Z}^{(i)},e_{p}\rangle=e_{p}(X_{i})-\operatorname{\mathbb{E}}\left[e_{p}(X_{i})\mid Z_{i}\right]$
by the reproducing property, and analogously $\zeta_{Y\mid Z,q}^{(i)}\coloneqq f_{q}(Y_{i})-\operatorname{\mathbb{E}}\left[f_{q}(Y_{i})\mid Z_{i}\right]$. We then obtain
$\left\|\Sigma_{XYZ\mid Z}\right\|^{2}$ $=\langle\operatorname{\mathbb{E}}[\Psi_{1}],\operatorname{\mathbb{E}}[\Psi_{2}]\rangle=\operatorname{\mathbb{E}}\left[\langle\Psi_{1},\Psi_{2}\rangle\right]$
$=\operatorname{\mathbb{E}}\left[\langle\epsilon_{X\mid Z}^{(1)},\epsilon_{X\mid Z}^{(2)}\rangle\langle\epsilon_{Y\mid Z}^{(1)},\epsilon_{Y\mid Z}^{(2)}\rangle k_{\mathcal{Z}}(Z_{1},Z_{2})\right]$
$=\operatorname{\mathbb{E}}\left[\sum_{p,q=1}^{\infty}\zeta_{X\mid Z,p}^{(1)}\zeta_{X\mid Z,p}^{(2)}\zeta_{Y\mid Z,q}^{(1)}\zeta_{Y\mid Z,q}^{(2)}k_{\mathcal{Z}}(Z_{1},Z_{2})\right],$ (3)
where the last equality follows by Parseval’s identity to the inner products. Our SGCM statistic is constructed as an empirical analogue of (3) by replacing the infinite sums with finite ones.
Let $\left\{(X_{i},Y_{i},Z_{i})\right\}_{i=1}^{n}$ be i.i.d. samples from $P_{XYZ}$. To avoid some technical issues, we split the index set $\{1,\ldots,n\}$ into two disjoint parts, $I_{1}\subset\{1,\ldots,n\}$ and $I_{2}=\{1,\ldots,n\}\setminus I_{1}$, with their sizes $n_{1}=|I_{1}|$ and $n_{2}=|I_{2}|$. First, using $I_{2}$, we construct data-dependent basis functions that exploit the information in the data; in principle, any orthonormal basis could be used. Specifically, for the empirical (uncentered) covariance operators
$\hat{C}_{XX}\coloneqq\frac{1}{n_{2}}\sum_{\ell\in I_{2}}k_{\mathcal{X}}(X_{\ell},\cdot)\otimes k_{\mathcal{X}}(X_{\ell},\cdot),\quad\hat{C}_{YY}\coloneqq\frac{1}{n_{2}}\sum_{\ell\in I_{2}}k_{\mathcal{Y}}(Y_{\ell},\cdot)\otimes k_{\mathcal{Y}}(Y_{\ell},\cdot),$
we employ the principal components as the basis functions. Namely, by the spectral theorem for compact, self-adjoint operators, we obtain the orthonormal eigensystems $\{(\hat{\lambda}_{p},\hat{e}_{p})\}_{p=1}^{n_{2}}$ and $\{(\hat{\mu}_{q},\hat{f}_{q})\}_{q=1}^{n_{2}}$ such that
$\hat{C}_{XX}=\sum_{p=1}^{n_{2}}\hat{\lambda}_{p}\,\hat{e}_{p}\otimes\hat{e}_{p},\qquad\hat{C}_{YY}=\sum_{q=1}^{n_{2}}\hat{\mu}_{q}\,\hat{f}_{q}\otimes\hat{f}_{q},$
with the eigenvalues ordered decreasingly and $\|\hat{e}_{p}\|_{\mathcal{H}_{\mathcal{X}}}=\|\hat{f}_{q}\|_{\mathcal{H}_{\mathcal{Y}}}=1$. Using the eigenfunctions corresponding to the top $P$ and $Q$ eigenvalues, $\{\hat{e}_{p}\}_{p=1}^{P}$ and $\{\hat{f}_{q}\}_{q=1}^{Q}$, as the basis yields an efficient approximation to the infinite expansion in (3); see Section 3.3 for the selection scheme of truncation levels $P$ and $Q$. The explicit forms of the derived bases are detailed in the Supplementary Material (Appendix G).
Second, we construct the statistic with the subsample $I_{1}$ given the basis learned on $I_{2}$. For each $p\in[P]$ and $q\in[Q]$, we fit (one-dimensional) nonparametric regressions of the scores $\{\hat{e}_{p}(X_{i})\}_{i\in I_{1}}$ and $\{\hat{f}_{q}(Y_{i})\}_{i\in I_{1}}$ on $\{Z_{i}\}_{i\in I_{1}}$ to estimate the conditional expectations $\operatorname{\mathbb{E}}[\hat{e}_{p}(X)\mid\hat{e}_{p},Z=\cdot]$ and $\operatorname{\mathbb{E}}[\hat{f}_{q}(Y)\mid\hat{f}_{q},Z=\cdot]$ by any regression method, giving the estimated regression functions $\hat{\operatorname{\mathbb{E}}}[\hat{e}_{p}(X)\mid\hat{e}_{p},Z=\cdot]$ and $\hat{\operatorname{\mathbb{E}}}[\hat{f}_{q}(Y)\mid\hat{f}_{q},Z=\cdot]$. The residual scores for $i\in I_{1}$ are then defined by
$\hat{\zeta}^{(i)}_{X\mid Z,p}\coloneqq\hat{e}_{p}(X_{i})-\hat{\operatorname{\mathbb{E}}}[\hat{e}_{p}(X)\mid\hat{e}_{p},Z=Z_{i}],\qquad\hat{\zeta}^{(i)}_{Y\mid Z,q}\coloneqq\hat{f}_{q}(Y_{i})-\hat{\operatorname{\mathbb{E}}}[\hat{f}_{q}(Y)\mid\hat{f}_{q},Z=Z_{i}].$
Finally, with a positive definite kernel $k_{\mathcal{Z}}$, we define the test statistic, called the SGCM, as the empirical analogue of the truncated expansion in (3):
$n_{1}\,\hat{T}_{n_{1}}=\frac{1}{n_{1}}\sum_{i,j\in I_{1}}\sum_{p=1}^{P}\sum_{q=1}^{Q}\hat{\zeta}^{(i)}_{X\mid Z,p}\,\hat{\zeta}^{(i)}_{Y\mid Z,q}\,\hat{\zeta}^{(j)}_{X\mid Z,p}\,\hat{\zeta}^{(j)}_{Y\mid Z,q}\,k_{\mathcal{Z}}(Z_{i},Z_{j}).$
This is a kernel-weighted quadratic form in the residuals from the conditional expectations computed along the retained empirical spectral coordinates, aggregated over all observation pairs via $k_{\mathcal{Z}}(Z_{i},Z_{j})$. Each coordinatewise regression is $\mathbb{R}$-valued, so modern nonparametric or machine-learning methods (boosting, neural networks, kernel regression) can be employed directly without constructing RKHS-valued CMEs. Leveraging such flexible learners improves the fidelity of conditional mean removal and thereby reduces bias in the conditional expectations, which in turn curbs Type I error inflation under $H_{0}$.
3.2 Testing Procedure
Under $H_{0}$ with some assumptions, our statistic $n_{1}\hat{T}_{n_{1}}$ is asymptotically a degenerate $V$-statistic. The limiting null distribution consists of an infinite weighted sum of chi-square random variables, and is non-pivotal. Therefore, our procedure requires the bootstrap to construct the critical value. The details are as follows: {longlist}
Compute $n_{1}\hat{T}_{n_{1}}$.
For $b=1,\dotsc,B$, generate i.i.d. weights $\{W^{\ast(b)}_{i}\}_{i\in I_{1}}$ with mean zero, unit variance, and finite fourth moment, and compute
$n_{1}\,\hat{T}^{\ast(b)}_{n_{1}}\coloneqq\frac{1}{n_{1}}\sum_{i,j\in I_{1}}\sum_{p=1}^{P}\sum_{q=1}^{Q}\,\hat{\zeta}^{(i)}_{X\mid Z,p}\,\hat{\zeta}^{(i)}_{Y\mid Z,q}\,\hat{\zeta}^{(j)}_{X\mid Z,p}\,\hat{\zeta}^{(j)}_{Y\mid Z,q}\,k_{\mathcal{Z}}(Z_{i},Z_{j})\,W_{i}^{\ast(b)}W_{j}^{\ast(b)}.$
Let $Q_{n,1-\alpha}^{\ast(B)}$ be the empirical $(1-\alpha)$-quantile of $\{n_{1}\hat{T}^{\ast(b)}_{n_{1}}\}_{b=1}^{B}$.
Reject $H_{0}$ if $n_{1}\hat{T}_{n_{1}}>Q_{n,1-\alpha}^{\ast(B)}$.
In Step 2, examples include standard normal, Rademacher, and Mammen’s two-point multipliers.
3.3 Choice of the Truncation Levels
In principle, retaining sufficiently many eigenfunctions would make the statistic coincide with the squared Hilbert–Schmidt norm; however, the computational cost is prohibitive and the contribution of the highest-order components is practically negligible. Accordingly, we choose the truncation levels $P$ and $Q$ as the smallest integers such that the fraction of variance explained (FVE; PetersenandMuller2016) satisfies $\operatorname{FVE}_{X}(P)\geq\tau$ and $\operatorname{FVE}_{Y}(Q)\geq\tau$ for some $\tau>0$. Here, $\{\hat{\lambda}_{\ell}\}_{\ell=1}^{n_{2}}$ and $\{\hat{\mu}_{\ell}\}_{\ell=1}^{n_{2}}$ are the eigenvalues of the empirical operators $\hat{C}_{XX}$ and $\hat{C}_{YY}$, respectively, ordered in descending order, and the FVE for each variable is defined as
$\operatorname{FVE}_{X}(P)\coloneqq\frac{\sum_{\ell=1}^{P}\hat{\lambda}_{\ell}}{\sum_{\ell=1}^{n_{2}}\hat{\lambda}_{\ell}},\qquad\operatorname{FVE}_{Y}(Q)\coloneqq\frac{\sum_{\ell=1}^{Q}\hat{\mu}_{\ell}}{\sum_{\ell=1}^{n_{2}}\hat{\mu}_{\ell}}$.
While we fix $\tau=0.80$ in our simulations for a practical choice of $P$ and $Q$, our theory requires them to slowly diverge under suitable regularity conditions.
4 Theoretical Properties
Assume that the subsample sizes $n_{1},n_{2}$ and the truncation levels $P_{n},Q_{n}$ all diverge as $n\to\infty$. The discussion below is asymptotic in $n$ unless otherwise noted.
Assumption 4.1.
The eigenvalues $\{\lambda_{j}\}_{j\geq 1}$ of $C_{XX}$ and $\{\mu_{j}\}_{j\geq 1}$ of $C_{YY}$ satisfy the following: (i) They are strictly decreasing, $\lambda_{1}>\lambda_{2}>\cdots\downarrow 0$ and $\mu_{1}>\mu_{2}>\cdots\downarrow 0$; (ii) The spectral gaps $\gamma^{X}_{P_{n}}\coloneqq\lambda_{P_{n}}-\lambda_{P_{n}+1}$ and $\gamma^{Y}_{Q_{n}}\coloneqq\mu_{Q_{n}}-\mu_{Q_{n}+1}$ satisfy the growing-gap conditions, $\gamma^{X}_{P_{n}}\sqrt{n_{2}}\to\infty$ and $\gamma^{Y}_{Q_{n}}\sqrt{n_{2}}\to\infty$.
If the operators $C_{XX}$ and $C_{YY}$ are infinite-dimensional, the assumption of strictly decreasing eigenvalues can be relaxed to permit a finite number of repeated eigenvalues.
Define the coordinate-summed mean squared errors as
$B^{X}_{n_{1}}(P_{n})\coloneqq\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{p=1}^{P_{n}}\bigl(b^{(i)}_{X\mid Z,p}\bigr)^{2},\qquad B^{Y}_{n_{1}}(Q_{n})\coloneqq\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{q=1}^{Q_{n}}\bigl(b^{(i)}_{Y\mid Z,q}\bigr)^{2},$
where $b^{(i)}_{X\mid Z,p}\coloneqq\hat{\operatorname{\mathbb{E}}}[\hat{e}_{p}(X_{i})\mid Z_{i},\hat{e}_{p}]-\operatorname{\mathbb{E}}[\hat{e}_{p}(X_{i})\mid Z_{i},\hat{e}_{p}]$ and $b^{(i)}_{Y\mid Z,q}\coloneqq\hat{\operatorname{\mathbb{E}}}[\hat{f}_{q}(Y_{i})\mid Z_{i},\hat{f}_{q}]-\operatorname{\mathbb{E}}[\hat{f}_{q}(Y_{i})\mid Z_{i},\hat{f}_{q}]$ are the pointwise regression errors for the $p$-th and $q$-th empirical coordinates.
Assumption 4.2.
Assume $B^{X}_{n_{1}}(P_{n})=o_{p}(1)$ and $B^{Y}_{n_{1}}(Q_{n})=o_{p}(1)$.
Assumption 4.3.
Assume $n_{1}\,B^{X}_{n_{1}}(P_{n})\,B^{Y}_{n_{1}}(Q_{n})=o_{p}(1)$.
Assumption 4.4.
Assume $\|\operatorname{\mathbb{E}}[\Psi\otimes\Psi]\|_{\textrm{op}}>0$.
Assumptions 4.2 and 4.3 have two advantages. First, they are sources of double robustness: if the regression errors satisfy $B_{n_{1}}^{X}(P_{n})=O_{p}(n_{1}^{-\alpha})$ and $B_{n_{1}}^{Y}(Q_{n})=O_{p}(n_{1}^{-\beta})$ for some $\alpha,\beta>0$ with $\alpha+\beta>1$, then Assumption 4.3 holds. Thus, one side may converge slowly provided the other is sufficiently fast. Second, they relax the conditions imposed in Pogodin et al. (2025). Specifically, they derive the null distribution under requirements that are similar to (or even stronger than) $n_{1}B_{n_{1}}^{X}(P_{n})=o_{p}(1)$ and $n_{1}B_{n_{1}}^{Y}(Q_{n})=o_{p}(1)$. Meanwhile, by utilizing the CI property, our framework significantly relaxes these requirements to Assumption 4.3 for deriving the asymptotic distribution. Assumption 4.4 rules out degenerate asymptotic distributions.
Remark 4.1.
A refined theoretical analysis of the bias of the CME estimators for the test statistic of Pogodin et al. (2025) is provided in the Supplementary Material (Appendix E).
4.1 Asymptotic Behavior
We characterize the limiting behavior of our statistic under the null and alternative hypotheses. Assumption 2.2 is only required for the test’s consistency, but is not needed to establish the asymptotic distribution under the null.
Theorem 4.1.
If Assumptions 2.1, 4.14.4 hold, then under $H_{0}$, we have
$n_{1}\hat{T}_{n_{1}}\overset{d}{\longrightarrow}\sum_{\ell=1}^{\infty}\xi_{\ell}\,G_{\ell}^{2},$
where $\{G_{\ell}\}_{\ell=1}^{\infty}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$ and $\{\xi_{\ell}\}_{\ell=1}^{\infty}$ are the positive eigenvalues of the covariance operator $\operatorname{\mathbb{E}}[\Psi\otimes\Psi]\neq 0$. Moreover, if Assumptions 2.1, 2.2, 4.1, and 4.2 hold, then under $H_{1}$, we have for every fixed $M\in\mathbb{R}$,
$\mathbb{P}\left(n_{1}\hat{T}_{n_{1}}>M\right)\longrightarrow 1.$
4.2 Size control
We then establish the validity of the wild-bootstrap and resulting size control. Consider the bootstrap version of our statistic,
$n_{1}\,\hat{T}_{n_{1}}^{\ast}=\frac{1}{n_{1}}\sum_{i,j\in I_{1}}\sum_{p=1}^{P}\sum_{q=1}^{Q}\hat{\zeta}^{(i)}_{X\mid Z,p}\,\hat{\zeta}^{(i)}_{Y\mid Z,q}\,\hat{\zeta}^{(j)}_{X\mid Z,p}\,\hat{\zeta}^{(j)}_{Y\mid Z,q}\,k_{\mathcal{Z}}(Z_{i},Z_{j})W_{i}^{\ast}\,W_{j}^{\ast},$
where $\{W_{i}^{\ast}\}_{i\in I_{1}}$ are i.i.d. bootstrap weights characterized in Section 3.2.
Theorem 4.2.
If Assumptions 2.1, 4.14.4 hold, then under $H_{0}$, we have
$n_{1}\hat{T}_{n_{1}}^{\ast}\overset{d^{\ast}}{\longrightarrow}\sum_{\ell=1}^{\infty}\xi_{\ell}G_{\ell}^{2}~~~\text{in probability}.$
This theorem shows the bootstrap validity of our procedure. This result immediately yields pointwise asymptotic size control.
Corollary 4.3.
If Assumptions 2.1, 4.14.4 hold, then, under $H_{0}$, we have
$\mathbb{P}\left(n_{1}\hat{T}_{n_{1}}>Q_{n,1-\alpha}^{\ast(B)}\right)\longrightarrow\alpha.$
Moreover, if Assumptions 2.1, 2.2, 4.1, and 4.2 hold, then, under $H_{1}$, we have
$\mathbb{P}\left(n_{1}\hat{T}_{n_{1}}>Q_{n,1-\alpha}^{\ast(B)}\right)\longrightarrow 1.$
We next aim to uniformly control the asymptotic size. Denote by $\mathcal{P}_{0}$ the family of Borel probability measures on $\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}$ satisfying $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$. As discussed in Section 1.1, Shah and Peters (2020) show that every sequence of CI tests with uniform asymptotic level $\alpha$ suffers from trivial power against alternatives unless further assumptions are imposed. As discussed in Section 1.1.1, regression-based procedures often rely on surrogate dependence measures for CI. In contrast, Corollary 4.3 shows that our test based on the SGCM attains asymptotic power one against a broad class of fixed alternatives under suitable conditions such as double robustness. The question is then under which additional conditions this procedure can be made uniformly valid in size.
In our framework, we assume that all conditional expectation components entering the statistic admit estimators whose bias converges to zero uniformly over $P\in\mathcal{P}_{0}$ under a few extra assumptions, as in Shah and Peters (2020); Lundborg, Shah and Peters (2022). Under these assumptions, the test based on the SGCM enjoys uniform asymptotic size control while still retaining nontrivial power against a broad class of general alternatives.
We define a subfamily $\tilde{\mathcal{P}}_{0}\subset\mathcal{P}_{0}$ that satisfies Assumptions 4.54.8 below.
Assumption 4.5.
The eigenvalues $\{\lambda_{j}\}_{j\geq 1}$ of $C_{XX}$ and $\{\mu_{j}\}_{j\geq 1}$ of $C_{YY}$ satisfy the following: (i) They are strictly decreasing, $\lambda_{1}>\lambda_{2}>\cdots\downarrow 0$ and $\mu_{1}>\mu_{2}>\cdots\downarrow 0$; (ii) The minimum spectral gaps $c_{P_{n}}^{X}\coloneqq\inf_{P\in\tilde{\mathcal{P}}_{0}}(\lambda_{P_{n}}-\lambda_{P_{n}+1})$ and $c_{Q_{n}}^{Y}\coloneqq\inf_{P\in\tilde{\mathcal{P}}_{0}}(\mu_{Q_{n}}-\mu_{Q_{n}+1})$ satisfy the growing-gap conditions, $c_{P_{n}}^{X}\sqrt{n_{2}}\to\infty$ and $c_{Q_{n}}^{Y}\sqrt{n_{2}}\to\infty$.
Assumption 4.6.
Assume that $B^{X}_{n_{1}}(P_{n})=o_{\tilde{\mathcal{P}}_{0}}(1)$ and $B^{Y}_{n_{1}}(Q_{n})=o_{\tilde{\mathcal{P}}_{0}}(1)$. Moreover, assume $n_{1}B^{X}_{n_{1}}(P_{n})B^{Y}_{n_{1}}(Q_{n})=o_{\tilde{\mathcal{P}}_{0}}(1)$.
Assumption 4.7.
There exist orthonormal bases $\{\tilde{e}_{p}\}_{p=1}^{\infty}\subset\mathcal{H}_{\mathcal{X}}$, $\{\tilde{f}_{p}\}_{p=1}^{\infty}\subset\mathcal{H}_{\mathcal{Y}}$, and $\{\tilde{g}_{p}\}_{p=1}^{\infty}\subset\mathcal{H}_{\mathcal{Z}}$ such that
$\lim_{K\to\infty}\sup_{P\in\tilde{\mathcal{P}}_{0}}\sum_{p>K}\left\{\operatorname{\mathbb{E}}_{P}[\tilde{e}_{p}(X)^{2}]+\operatorname{\mathbb{E}}_{P}[\tilde{f}_{p}(Y)^{2}]+\operatorname{\mathbb{E}}_{P}[\tilde{g}_{p}(Z)^{2}]\right\}=0.$
Assumption 4.8.
Assume $\inf_{P\in\tilde{\mathcal{P}}_{0}}\left\|\operatorname{\mathbb{E}}_{P}[\Psi\otimes\Psi]\right\|_{\textrm{op}}>0$.
Assumptions 4.5 and 4.6 are uniform versions of Assumptions 4.1, 4.2 and 4.3. Regarding Assumption 4.5, if we consider parametric distributions for $X$ and $Y$ where the parameter set is compact and the density functions are continuous with respect to the parameter $\theta$, it can be shown that minimum spectral gaps and the associated $P_{n}$ and $Q_{n}$ exist, satisfying this assumption; e.g., for the Gaussian family, this holds if $P_{n}$ and $Q_{n}$ grow more slowly than $\log n$; see the Supplementary Material (Appendix D) for details. Assumption 4.6 can be achieved using methods like kernel ridge regression or a series estimator, provided that the divergence rate of the truncation level is sufficiently slow and some additional assumptions are met. Sufficient conditions for these approaches are provided in Shah and Peters (2020) and Lundborg et al. (2024), respectively. Assumption 4.7 is sufficient for uniform tightness and a uniform central limit theorem with respect to $\tilde{\mathcal{P}}_{0}$, akin to Lundborg, Shah and Peters (2022). Assumption 4.8 means that the covariance operator of $\Psi$ is uniformly nondegenerate over $\tilde{\mathcal{P}}_{0}$.
Theorem 4.4.
Suppose Assumption 2.1 holds. For each bootstrap replication $b=1,\dots,B$, let the multipliers $\{W_{i}^{(b)}\}_{i\in I_{1}}$ be i.i.d. $\mathcal{N}(0,1)$. Then, with respect to $\tilde{\mathcal{P}}_{0}$ satisfying Assumptions 4.54.8, the test $\phi_{n,P,B}=\mathds{1}\{n_{1}\hat{T}_{n_{1}}>Q_{n,1-\alpha}^{\ast(B)}\}$ achieves a uniform asymptotic level $\alpha$:
$\lim_{n\to\infty}\sup_{P\in\tilde{\mathcal{P}}_{0}}\left|\mathbb{P}_{P}(\phi_{n,P,B}=1)-\alpha\right|=0.$
5 Kernel Design for Random Objects
Kernel-based procedures have the practical advantage of a modular design. In our framework, it suffices to select a bounded, characteristic kernel on the space of interest. In Euclidean spaces, Gaussian and Laplacian kernels are characteristic (Sriperumbudur, Fukumizu and Lanckriet, 2011); on separable Hilbert spaces, the exponential (Gaussian) kernel of the squared Hilbert norm remains characteristic (Wynne and Duncan, 2022). Beyond Hilbert spaces, if the underlying metric is of strongly negative type (Lyons2013), one can thereby construct characteristic kernels (Ziegeletal2024). However, verifying this assumption is often difficult in practice. Moreover, on spheres, strong negative type typically requires a stringent support restriction (Lyons2020), but spaces such as $L^{1}$ fail to satisfy it altogether (Lyons2013). This limitation is substantial because a broad family of metrics, including trees and graph shortest-path metrics, admits isometric embeddings into $L^{1}$ (DezaandLaurent1997). Hence, a meaningful gap remains in constructing broadly applicable characteristic kernels on general metric spaces. To address this, we do not require metrics of strongly negative type; instead, we work with semimetrics of negative type to construct characteristic kernels on Polish spaces.
5.1 Theoretical Foundations for Characteristic Kernels
We develop a general construction of characteristic kernels on Polish spaces. First, we show that for any continuous semimetric $\rho$ of negative type on a Polish space $\mathcal{M}$, an exponential kernel is characteristic. We then extend this to all non-constant completely monotone transforms, and finally show that finite tensor products preserve characteristicity.
Definition 5.1 (Semimetric).
Let $\mathcal{S}\neq\varnothing$ and $\rho\colon\mathcal{S}\times\mathcal{S}\to[0,\infty)$. If, for all $s,s^{\prime}\in\mathcal{S}$, (i) $\rho(s,s^{\prime})=0$ if and only if $s=s^{\prime}$, and (ii) $\rho(s,s^{\prime})=\rho(s^{\prime},s)$, then $\rho$ is called a semimetric on $\mathcal{S}$. We do not necessarily require the triangle inequality.
Definition 5.2 (Negative type).
A semimetric $\rho$ on $\mathcal{S}$ is of negative type if, for every $n\geq 2$, all $s_{1},\dots,s_{n}\in\mathcal{S}$, and all $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$ with $\sum_{i=1}^{n}\alpha_{i}=0$, we have $\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{i}\alpha_{j}\rho(s_{i},s_{j})\leq 0$.
Remark 5.1.
In the RKHS literature, this property is also referred to as (conditionally) negative definite.
Theorem 5.1.
Let $\mathcal{M}$ be Polish and let $\rho\colon\mathcal{M}\times\mathcal{M}\to[0,\infty)$ be a continuous semimetric of negative type. Then, for every $\gamma>0$, the exponential kernel
$k_{\mathrm{exp},\rho}(u,v)\coloneqq\exp\left(-\gamma\,\rho(u,v)\right),\qquad u,v\in\mathcal{M},$
is characteristic on $\mathcal{M}$.
The proof uses that the pull-back kernel obtained from a characteristic kernel via an injective map remains characteristic; see the Supplementary Material (Lemma C.1). If a semimetric $\rho$ is of negative type, then any fractional power $\rho^{\,q}$ with $q\in(0,1)$ is also a semimetric of negative type (Berg, Christensen and Ressel, 1984, Chapter 3, Corollary 2.10). Moreover, for a Polish space $(\mathcal{M},d_{\mathcal{M}})$, if $d_{\mathcal{M}}^{\,p}$ is of negative type, then the distance-induced exponential kernel $k_{\mathrm{exp},d_{\mathcal{M}}^{p}}(u,v)\coloneqq\exp(-\gamma\,d_{\mathcal{M}}(u,v)^{p})$ is characteristic on $\mathcal{M}$.
The following corollary provides a further generalization, showing that a characteristic kernel can be constructed from any non-constant completely monotone function, not just the exponential.
Definition 5.3 (Completely monotone function).
A function $f\colon(0,\infty)\to\mathbb{R}$ is completely monotone if it satisfies $f\in C^{\infty}(0,\infty)$ and $(-1)^{n}f^{(n)}(t)\geq 0$ for every $t>0$ and $n\in\{0\}\cup\mathbb{N}$.
Corollary 5.2.
Let $\mathcal{M}$ be Polish and let $\rho\colon\mathcal{M}\times\mathcal{M}\to[0,\infty)$ be a continuous semimetric of negative type. If $f\colon[0,\infty)\to\mathbb{R}$ is non-constant, completely monotone on $(0,\infty)$, and satisfies $f(0)=\lim_{t\downarrow 0}f(t)$, then
$k_{f}(u,v)\coloneqq f(\rho(u,v))$
is bounded, continuous, positive definite, and characteristic on $\mathcal{M}$.
The condition on $f$ in Corollary 5.2 holds if and only if there exists a unique finite positive Borel measure $\mu$ on $0,\infty)$ with $\mu((0,\infty))>0$ such that
$f(t)=\int_{0}^{\infty}e^{-st}\,\mu(\mathop{}\!\mathrm{d}s),\qquad t\in[0,\infty)$
by Bernstein’s theorem (Schilling, Song and Vondracek, 2012, Theorem 1.4). In other words, every such $f$ can be written as a finite mixture of Laplace kernels $t\mapsto e^{-st}$. Theorem 5.1 is recovered as the special case where $\mu$ is the Dirac measure at $\gamma$, so that $f(t)=e^{-\gamma t}$. Therefore, Corollary 5.2 is a generalization of Theorem 5.1. While this class matches Ziegeletal2024, we prefer to formulate the assumptions directly in terms of $f$ because this is often more convenient for constructing examples. For instance, it contains the rational-quadratic $f(t)=\left(1+c\,t\right)^{-\alpha}$ with $\alpha,c>0$; see MillerandSamko2001 for further examples.
For RKHS-based methods, positive definiteness is the basic requirement: it guarantees the existence of an RKHS with the exponential kernel as its reproducing kernel and a well-defined feature map. By Schoenberg’s theorem (Berg, Christensen and Ressel, 1984, Chapter 3, Theorem 2.2), the exponential kernel is positive definite for any $\gamma>0$ if and only if $\rho$ is negative type. Moreover, when $\mathcal{M}$ is an infinite-dimensional separable Hilbert space and $\rho$ is the squared Hilbert distance, it is known that a kernel of the form $f\circ\rho$ is positive definite if and only if $f$ is completely monotone (Berg, Christensen and Ressel, 1984, Chapter 5, Section 2). Motivated by these results, Theorem 5.1 and Corollary 5.2 show that, for the exponential kernel and for kernels with the squared Hilbert distance built from non-constant completely monotone functions, positive definiteness upgrades to characteristicness on Polish spaces without additional geometric assumptions.
Theorem 5.1 and Corollary 5.2 therefore reduce the search for characteristic kernels to verifying the negative-type property, aligning our analysis with established results. For Riemannian manifolds, criteria ensuring that geodesic distances are of negative type are surveyed by Feragenetal2015; conditions under which metric spaces are of negative type (e.g., hyperbolic spaces and the sphere) are compiled by Istas2012; and for graphs and trees, DezaandLaurent1997 and ObataandZakiyyah2018 provide criteria. We further develop these themes with additional examples and new results in Section 5.2.
Remark 5.2.
Bhattacharjee, Li and Xue (2025) introduce the weak conditional Fréchet mean and show that it coincides with the usual conditional Fréchet mean under cc-universality of the kernel on the response space; see also Zhang, Xue and Li (2024). However, cc-universality may fail to ensure this identification on non-compact spaces, whereas a bounded characteristic kernel does provide a sufficient condition. Hence, the bounded characteristic kernels on metric spaces obtained in Theorem 5.1 and Corollary 5.2 are well suited for Fréchet regression. See the Supplementary Material (Appendix C.3).
Before turning to concrete examples, we record a tensor-product result that will underpin our CI test, which requires the product kernel $k_{\mathcal{X}}\otimes k_{\mathcal{Z}}$ or $k_{\mathcal{Y}}\otimes k_{\mathcal{Z}}$ to be characteristic (Assumption 2.2). As highlighted in SzaboandSriperumbudur2018, such tensor-product results are foundational for kernel two-sample and independence tests. The following tensorization result ensures that product kernels inherit characteristicness.
Corollary 5.3.
Let $L\in\mathbb{N}$ and $\mathcal{N}\coloneqq\prod_{\ell=1}^{L}\mathcal{N}_{\ell}$, where each $\mathcal{N}_{\ell}$ is Polish. For each $\ell\in\{1,\dots,L\}$, let $\rho_{\ell}\colon\mathcal{N}_{\ell}\times\mathcal{N}_{\ell}\to[0,\infty)$ be a continuous semimetric of negative type and Let $f_{\ell}\colon[0,\infty)\to[0,\infty)$ be non-constant, completely monotone on $(0,\infty)$, satisfy $f_{\ell}(0)=\lim_{t\downarrow 0}f_{\ell}(t)$. Define, for $u_{\ell},v_{\ell}\in\mathcal{N}_{\ell}$,
$k^{(\ell)}(u_{\ell},v_{\ell})\coloneqq f_{\ell}(\rho_{\ell}(u_{\ell},v_{\ell})),$
and, for $u=(u_{1},\dots,u_{L}),v=(v_{1},\dots,v_{L})\in\mathcal{N}$,
$k_{\otimes,L}(u,v)\coloneqq\prod_{\ell=1}^{L}k^{(\ell)}(u_{\ell},v_{\ell}).$
Then $k_{\otimes,L}$ is bounded, continuous, positive definite, and characteristic on $\mathcal{N}$.
Taken together, these results imply that whenever $\rho$ is of negative type, composing $\rho$ with any non-constant completely monotone function including the exponential yields bounded, continuous characteristic kernels; finite products of such kernels retain these properties and fulfill Assumptions 2.1 and 2.2.
5.2 Illustrative Spaces
We now illustrate the toolkit on five emblematic non-Euclidean domains including time-varying random objects. Our guiding principle is simple: if a metric space $(\mathcal{M},d_{\mathcal{M}})$ is Polish and its powered distance $d_{\mathcal{M}}^{\,q}$ is of negative type for some $q>0$, then by Theorem 5.1 the distance-induced exponential kernel $k(x,y)=\exp\left(-\gamma\,d_{\mathcal{M}}(x,y)^{q}\right)$ is characteristic on $\mathcal{M}$. Therefore, in what follows we only verify Polishness and negative type for each example; characteristicity then follows immediately and will not be repeated.
Example 5.1 (The $L^{p}$ Spaces).
The Lebesgue spaces $L^{p}(\Omega)$, where $\Omega$ is a separable measurable space, are standard domains for functional data (e.g., growth curves and temperature records; see RamsayandSilverman2005). For $1\leq p\leq 2$, Theorem 2 in Bretagnolleetal1965 implies that the powered norm $\|f-g\|_{p}^{q}$ is of negative type for every $0<q\leq p$. In sharp contrast, for $p>2$, no nontrivial exponent $q>0$ renders $\|f-g\|_{p}^{q}$ of negative type on $L^{p}$ (Lennardetal1997, Theorem 1.3 and 2.4). Since $L^{p}(\Omega)$ is a separable Banach space under its norm metric, it is Polish (Brezis2011).
Example 5.2 (The Sphere).
The unit sphere $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:\left\|x\right\|_{2}=1\}$ is the standard domain for directional data. The great-circle (geodesic) distance is $d_{\mathbb{S}}(x,y)\coloneqq\arccos(\langle x,y\rangle)$. For this distance, Istas2012 proves that $d_{\mathbb{S}}^{\,q}$ is of negative type for all $0<q\leq 1$. The same conclusion holds when $\mathbb{S}^{d-1}$ is replaced by the unit sphere in a Hilbert space $\mathcal{H}$, as shown in the Supplementary Material (Proposition C.2). If $\mathcal{H}$ is a separable Hilbert space, then its unit sphere $\mathbb{S}(\mathcal{H})$ is Polish.
Example 5.3 (The Simplex).
The simplex $\Delta^{d-1}\coloneqq\{x\in\mathbb{R}^{d}:x_{i}\geq 0,\ \sum_{i=1}^{d}x_{i}=1\}$ serves as the canonical space for compositional data. Define the element-wise square-root map $\sqrt{x}\coloneqq(\sqrt{x_{1}},\dots,\sqrt{x_{d}})$, which embeds $\Delta^{d-1}$ isometrically into the positive orthant of the unit sphere $\mathbb{S}^{d-1}$. The Fisher–Rao distance is $d_{\mathrm{FR}}(x,y)\coloneqq\arccos(\langle\sqrt{x},\sqrt{y}\rangle_{\mathbb{R}^{d}}).$ Since $d_{\mathbb{S}}^{\,q}$ is of negative type for all $0<q\leq 1$ on $\mathbb{S}^{d-1}$, the same holds for $d_{\mathrm{FR}}^{\,q}$ on $\Delta^{d-1}$. This construction extends to the space $\mathfrak{D}\coloneqq\{f\geq 0:\int_{\mathcal{X}}f\mathop{}\!\mathrm{d}\mu=1\}$ of probability densities on $(\mathcal{X},\mu)$, with
$d_{\mathrm{FR}}(f,g)\coloneqq\arccos\left(\langle\sqrt{f},\sqrt{g}\rangle_{L^{2}}\right).$
If $L^{2}(\mu)$ is separable, $(\mathfrak{D},d_{\mathrm{FR}})$ is Polish, and $d_{\mathrm{FR}}^{\,q}$ is of negative type for all $0<q\leq 1$.
Example 5.4 (The $p$-Wasserstein Space).
The $p$-Wasserstein space is the canonical space for distributional data (BritoandDias2022). Let $I\subseteq\mathbb{R}$ be an interval and $1\leq p\leq 2$. Denote by $\mathcal{W}_{p}(I)$ the set of Borel probability measures on $I$ with finite $p$-th moment. In one dimension,
$d_{\mathcal{W}_{p}}(\mu,\nu)\coloneqq\left(\int_{0}^{1}\left|F_{\mu}^{-1}(q)-F_{\nu}^{-1}(q)\right|^{p}\mathop{}\!\mathrm{d}q\right)^{1/p}=\left\|F_{\mu}^{-1}-F_{\nu}^{-1}\right\|_{L^{p}(0,1)}.$
Here $F_{\mu}^{-1}$ and $F_{\nu}^{-1}$ are quantile functions. Since on $L^{p}(0,1)$ the powered metric $\|u-v\|_{p}^{q}$ is of negative type for every $0<q\leq p$, it follows that $d_{\mathcal{W}_{p}}^{\,q}$ is of negative type on $\mathcal{W}_{p}(I)$ for all $0<q\leq p$. For $\mathbb{R}^{d}$ with $d\geq 2$, the native geometry of $\mathcal{W}_{2}(\mathbb{R}^{d})$ is non-Hilbertian. Nevertheless, the Monge (Brenier) map provides an injective map into an $L^{2}$-space (Brenier1991; Merigotetal2020) and yields characteristic kernels induced by this injective embedding by Lemma C.1 in the Supplementary Material, in the same spirit as Ziegeletal2024. While the tree-Wasserstein distance is of negative type (Leetal2019, Proposition 2), whether this holds for other ground metrics remains an open question. Regarding topology, if the value space $\mathcal{X}$ is Polish, then the Wasserstein space $(\mathcal{W}_{p}(\mathcal{X}),d_{\mathcal{W}_{p}})$ is also Polish; see Villani2009.
Example 5.5 (The $L^{p}$ space for time-varying random objects).
Here, we define an $L^{p}$ space of metric-valued functions to handle time-varying random objects. As LinandYao2019 emphasize, spherical functional data—for example, hurricane trajectories or spatio-temporal intensity/wind fields on $\mathbb{S}^{2}$—fit naturally within this framework. Further applications include evolving networks and covariance matrices; see DubeyandMuller2020.
Let $(\mathcal{T},\rho_{\mathcal{T}})$ be a separable metric space with a finite Radon measure $\nu$, and let $(\mathcal{M},d_{\mathcal{M}})$ be a metric space. Denote the class of $\nu$-measurable maps by $\mathsf{Meas}_{\nu}(\mathcal{T},\mathcal{M})$. For $1\leq p<\infty$,
$L^{p}(\mathcal{T},\mathcal{M})\coloneqq\Big\{f\in\mathsf{Meas}_{\nu}(\mathcal{T},\mathcal{M}):\int_{\mathcal{T}}d_{\mathcal{M}}(f(t),o)^{p}\nu(\mathop{}\!\mathrm{d}t)<\infty\text{ for some }o\in\mathcal{M}\Big\}\Big/\sim,$
with metric
$D_{p}(f,g)\coloneqq\left(\int_{\mathcal{T}}d_{\mathcal{M}}(f(t),g(t))^{p}\nu(\mathop{}\!\mathrm{d}t)\right)^{1/p},$
where $f\sim g$ if and only if $d_{\mathcal{M}}(f(t),g(t))=0$ for $\nu$-a.e. $t\in\mathcal{T}$. Corollary C.9 in the Supplementary Material proves that $D_{p}^{\,q}$ is of negative type on $L^{p}(\mathcal{T},\mathcal{M})$ whenever $0<q\leq p$ and $d_{\mathcal{M}}^{\,q}$ is of negative type on $\mathcal{M}$; moreover, the Polishness of $(L^{p}(\mathcal{T},\mathcal{M}),D_{p})$ is also established there, provided that $(\mathcal{M},d_{\mathcal{M}})$ is Polish.
Remark 5.3.
To our knowledge, this is the first work to construct characteristic kernels on the Hilbert sphere (geodesic distance), the space of probability densities (FR distance), and metric-valued $L^{p}$ spaces (for the admissible exponents $q>0$ specified above). Proofs are provided in the Supplementary Material (Appendix C).
6 Simulation Study
We assess the finite-sample behavior of our testing procedure based on the SGCM across various data-generating processes (DGPs). Empirical sizes and powers are reported at the $\alpha=0.05$ level, based on 1,0001{,}000 Monte Carlo replications. We use B=2, [^0] B=2{,}000 bootstrap replications with i.i.d. standard normal multipliers. Conditional expectations are estimated via three-fold cross-fitting with XGBoost as the regression learner. To separate eigenfunction estimation, we allocate 20% of observations to $I_{2}$ and use the remaining 80% ($I_{1}$) for subsequent analysis and test statistic evaluation.
We use the distance-induced exponential kernel $k_{\mathcal{S}}(s,s^{\prime})=\exp(-\gamma_{\mathcal{S}}d_{\mathcal{S}}(s,s^{\prime}))$ for $\mathcal{S}\in\{\mathcal{X},\mathcal{Y},\mathcal{Z}\}$, where the bandwidth is set by the median heuristic, $\gamma_{\mathcal{S}}=(\operatorname{med}_{i<j}d_{\mathcal{S}}(s_{i},s_{j}))^{-1}$. Specifically, we consider Euclidean data with the Euclidean distance in Section 6.1, and distributional data with the 1- and 2-Wasserstein, Fisher–Rao (FR), and Hellinger distances in Section 6.2.
6.1 Euclidean Data
We benchmark our CI test against representative methods drawn from regression-, kernel-, metric-, and classification-based families:
  • GCM
    Regression-based: Generalized Covariance Measure (Shah and Peters, 2020), via the R function GeneralisedCovarianceMeasure::gcm.test.
  • WGCM
    Regression-based: Weighted Generalized Covariance Measure (Scheidegger, Hörrmann and Bühlmann, 2022), via the R function weightedGCM::wgcm.est.
  • CDCOV
    Metric-based: Conditional Distance Covariance test (Wangetal2015), via the R function cdcsis::CDCOV.test.
  • KCI
    Kernel-based: Kernel Conditional Independence test (Zhang2011), via the R function CondIndTests::KCI with fixed default parameters.
  • MMDCI
    Kernel-based: Doubly robust MMD-based conditional independence test with conditional generators (Zhangetal2025); we employ the authors’ reference implementation with the recommended settings.
  • CCIT
    Others: Classifier Conditional Independence Test (Senetal2017); we employ the authors’ reference implementation with the recommended settings.
  • SGCM
    Proposed: our Spectral Generalized Covariance Measure.
Here, all the regression-based methods (GCM, WGCM) employ XGBoost as in our implementation.
6.1.1 Low-Dimensional DGPs
Following the low-dimensional settings by Shah and Peters (2020), we introduce the Gaussian-damped sine wave $f_{a}(z)=\exp(-z^{2}/2)\sin(az)$. For $Z,\varepsilon_{X},\varepsilon_{Y}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$ and each $a\in\{2,4\}$, we generate data under the null
$X=f_{a}(Z)+0.3\,\varepsilon_{X},\qquad Y=f_{a}(Z)+0.3\,\varepsilon_{Y}.$
We refer to the cases $a=2$ and $a=4$ as scenarios (a2) and (a4), respectively. As $a$ increases, the oscillations of $f_{a}(Z)$ become faster within the same Gaussian envelope, making departures from the null hypothesis progressively harder to detect. Figure 1 displays $f_{a}(\cdot)$ for $a=2$ and $a=4$.
Figure 1: Gaussian-damped sine wave $f_{a}(z)=\exp(-z^{2}/2)\sin(az)$ for $a=2$ (blue) and $a=4$ (red).
For each scenario (a2) and (a4), we consider three designs for the alternative hypotheses. The first design (DGP1-1) is a simple linear perturbation. The following two (DGP1-2 and DGP1-3) are designed to probe regression-based tests in more detail, as they involve an unobserved latent variable $U$, inducing a conditional dependence between $X$ and $Y$ given $Z$.
  1. DGP 1-1.
    $Y\leftarrow Y+0.2\,X$.
  2. DGP 1-2.
    $U\sim\mathcal{N}(0,1)$, with transformations $X\leftarrow X+0.3(U^{2}-1)$ and $Y\leftarrow Y+0.3\,U$.
  3. DGP 1-3.
    Same transformations as in DGP1-2, but with $U\leftarrow f_{1}(Z)+\varepsilon_{U}$ where $\varepsilon_{U}\sim\mathcal{N}(0,1)$.
We consider the cases with sample sizes $n\in\{100,200,300,400\}$. The theoretical expectation is that while both GCM and WGCM have difficulty achieving high power in DGP1-2, this difficulty is limited to only GCM in DGP1-3. This is because GCM only guarantees power for DGPs where $\operatorname{\mathbb{E}}[\operatorname{Cov}(X,Y\mid Z)]\neq 0$, whereas WGCM guarantees power for DGPs where $\operatorname{Cov}(X,Y\mid Z)\neq 0$ with a positive probability; see the Supplementary Material (Appendix F) for details.
6.1.2 Results for Low-Dimensional Settings
All the results are summarized in Table 2. As for the size, our SGCM is close to or slightly above the nominal level ($0.05$) in all the cases, even with small samples. The GCM and WGCM do not depart much from $0.05$, but are no better than the SGCM overall. The KCI shows acceptable control only with the milder scenario (a2) and larger $n$, while it exhibits a systematic size inflation with (a4). The CDCOV, MMDCI, and CCIT fail to control the size across every design and sample size, indicating that they are not valid in these settings.
Regarding the power, under the simple additive alternative (DGP1-1), the GCM and WGCM achieve high power and remain competitive. Our SGCM exhibits nontrivial power that increases with $n$, but is lower than the regression-based baselines. Under DGP1-2, where no power guarantee is available for the regression-based procedures, both GCM and WGCM yield only trivial power, whereas SGCM delivers consistently nontrivial power that strengthens with $n$. In DGP1-3, GCM remains essentially powerless, WGCM improves with $n$ yet stays below SGCM, and SGCM reaches high power. KCI often exhibits very high apparent power; however, its size distortions in (a4) preclude a fair power comparison. Since CCIT, CDCOV, and MMDCI fail to maintain the specified size, they are not included in the power comparison.
Overall, our SGCM achieves reliable finite-sample size control and power guarantees across the more challenging cases, outperforming GCM and WGCM in DGP1-2 and DGP1-3 while remaining competitive in DGP1-1.
Table 2: Empirical size ($\alpha=0.05$) and power summary.
Table 2.1: Results for DGP (a2) Sample Size DGP Method 100 200 300 400 Size gcm 0.088 0.070 0.064 0.065 wgcm 0.131 0.086 0.078 0.078 cdcov 1.000 1.000 1.000 1.000 kci 0.176 0.297 0.070 0.066 mmdci 1.000 1.000 1.000 1.000 ccit 0.447 0.423 0.439 0.450 sgcm 0.053 0.044 0.070 0.067 Power (DGP1-1) gcm 0.574 0.841 0.953 0.976 wgcm 0.538 0.678 0.774 0.850 cdcov kci 0.785 0.972 0.810 0.907 mmdci ccit sgcm 0.212 0.480 0.700 0.820 Power (DGP1-2) gcm 0.076 0.064 0.065 0.072 wgcm 0.110 0.081 0.081 0.086 cdcov kci 0.190 0.480 0.709 0.859 mmdci ccit sgcm 0.143 0.337 0.490 0.675 Power (DGP1-3) gcm 0.072 0.057 0.058 0.064 wgcm 0.159 0.279 0.434 0.553 cdcov kci 0.596 0.944 0.980 0.999 mmdci ccit sgcm 0.295 0.691 0.903 0.974
Table 2.2: Results for DGP (a4) Sample Size DGP Method 100 200 300 400 Size gcm 0.138 0.090 0.096 0.086 wgcm 0.208 0.118 0.102 0.091 cdcov 1.000 1.000 1.000 1.000 kci 0.985 0.999 0.981 0.997 mmdci 1.000 1.000 1.000 1.000 ccit 0.465 0.464 0.412 0.441 sgcm 0.060 0.071 0.088 0.073 Power (DGP1-1) gcm 0.669 0.863 0.968 0.984 wgcm 0.649 0.743 0.822 0.884 cdcov kci mmdci ccit sgcm 0.280 0.569 0.762 0.868 Power (DGP1-2) gcm 0.088 0.075 0.075 0.079 wgcm 0.154 0.100 0.097 0.094 cdcov kci mmdci ccit sgcm 0.145 0.310 0.470 0.626 Power (DGP1-3) gcm 0.100 0.079 0.075 0.066 wgcm 0.212 0.285 0.456 0.566 cdcov kci mmdci ccit sgcm 0.233 0.527 0.780 0.910
6.1.3 High-Dimensional DGPs
We generate high-dimensional data based on the post-nonlinear noise model of Shi et al. (2021). For each dimension $d\in\{10,20,30,40,50\}$, draw $Z\sim\mathcal{N}(0,I_{d})$ and $\varepsilon_{X},\varepsilon_{Y}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$, and define
$X=\sin\left(\beta_{X}^{\top}Z+0.5\,\varepsilon_{X}\right),\qquad Y=\cos\left(\beta_{Y}^{\top}Z+b\,X+0.5\,\varepsilon_{Y}\right),$
where $\beta_{X},\beta_{Y}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$ are some parameters. Here, we draw $\beta_{X},\beta_{Y}$ independently from $\mathrm{Unif}\left([-1,1]\right)$ and then normalize $\beta_{X}\leftarrow\beta_{X}/\left\|\beta_{X}\right\|_{1}$, $\beta_{Y}\leftarrow\beta_{Y}/\left\|\beta_{Y}\right\|_{1}$, and set $b=0.5\cdot\mathds{1}\{H_{1}\}$, which controls the direct effect from $X$ to $Y$.
6.1.4 Results for High-Dimensional Settings
Figure 2 reports empirical rejection rates under the null (left panels) and the alternative (right panels) as functions of the ambient dimension $d\in\{10,20,30,40,50\}$ for sample sizes $n\in\{200,400,600\}$. We do not include MMDCI from this comparison due to its high computational cost.
As for the size, SGCM, GCM, and WGCM keep Type I error close to the nominal level uniformly over $(d,n)$; their null curves are essentially flat around the target level. CDCOV is markedly conservative throughout. In contrast, KCI and CCIT exhibit systematic size inflation under the null, with the distortion attenuating as $d$ increases and, to a lesser extent, as $n$ grows.
Regarding the power, restricting attention to size-valid procedures, SGCM achieves the highest power uniformly over $(d,n)$. Its power decreases monotonically with $d$ and increases with $n$, remaining nontrivial across all settings considered. WGCM is the strongest of the remaining size-valid baselines but trails SGCM at every $(d,n)$ and degrades notably as $d$ increases. GCM yields the lowest power among size-valid methods, approaching the nominal level as $d$ grows. KCI and CCIT often show moderate to high apparent power; however, because they do not control size under the null, these power curves are not directly comparable as hypothesis tests. CDCOV displays negligible power across all settings.
Figure 2: Empirical rejection rates by dimension for the high-dimensional post-nonlinear noise model. Left column: null; right column: alternative. Rows correspond to sample sizes $n\in\{200,400,600\}$ (top to bottom).
6.2 Distributional Data
We study the behavior of our test when $X$ and $Y$ are probability distributions represented by finite samples. We examine how the choice of metric on the space of probability measures affects the behavior of our SGCM. In particular, we compare the 1–, 2–Wasserstein, Fisher–Rao (FR) distances and Hellinger distance. For one-dimensional probability distributions, $d_{\mathcal{W}_{1}}$ and $d_{\mathcal{W}_{2}}$ reduce to $L^{1}/L^{2}$ distances between quantile functions and are computed directly from empirical quantiles. For FR and Hellinger, we estimate densities via kernel density estimators and numerically evaluate the corresponding inner product on square-root densities; the distances are then obtained from this quantity.
6.2.1 DGPs
Each observation consists of a pair $(X_{i},Y_{i})$ of univariate empirical distributions together with a Euclidean covariate $Z_{i}\in\mathbb{R}^{10}$. For every $i$, both $X_{i}$ and $Y_{i}$ are observed as point clouds of common size $m=150$. We report results for sample sizes $n\in\left\{200,\;400,\;600\right\}$. Generate covariates and noise via
$Z_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,I_{10}),\qquad\varepsilon_{X,i},\varepsilon_{Y,i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1),$
mutually independent. Draw $\beta_{X},\beta_{Y}\in\mathbb{R}^{10}$ with coordinates i.i.d. from $\mathrm{Unif}([-1,1])$, then set $\beta_{X}\leftarrow\beta_{X}/\left\|\beta_{X}\right\|_{1}$ and $\beta_{Y}\leftarrow\beta_{Y}/\left\|\beta_{Y}\right\|_{1}$. A signal parameter
$c\in\left\{0,\;0.05,\;0.10,\;0.15,\;0.20\right\}$
controls conditional dependence by sharing $\varepsilon_{X,i}$ across $X$ and $Y$ in the alternatives below: $c=0$ corresponds to $H_{0}:X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$, whereas $c>0$ induces dependence of controlled strength. All remaining SGCM settings follow Section 6.1.
  1. DGP 2-1 (mean-varying).
    Define the latent means
    $\mu_{X,i}=\beta_{X}^{\top}Z_{i}+0.5\,\varepsilon_{X,i},\qquad\mu_{Y,i}(c)=\beta_{Y}^{\top}Z_{i}+c\,\varepsilon_{X,i}+0.5\,\varepsilon_{Y,i}.$
    Conditional on $(Z_{i},\varepsilon_{X,i},\varepsilon_{Y,i})$, sample
    $X_{i}=\{X_{i\ell}\}_{\ell=1}^{m},\quad X_{i\ell}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(\mu_{X,i},1),\qquad Y_{i}=\{Y_{i\ell}\}_{\ell=1}^{m},\quad Y_{i\ell}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(\mu_{Y,i}(c),1).$
  2. DGP 2-2 (variance-varying).
    Let
    $s_{X,i}=\beta_{X}^{\top}Z_{i}+0.5\,\varepsilon_{X,i},\qquad s_{Y,i}(c)=\beta_{Y}^{\top}Z_{i}+c\,\varepsilon_{X,i}+0.5\,\varepsilon_{Y,i},$
    and define variances $\sigma_{X,i}^{2}=\exp(s_{X,i})$ and $\sigma_{Y,i}^{2}(c)=\exp(s_{Y,i}(c))$. Conditional on these,
    $X_{i\ell}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma_{X,i}^{2}),\qquad Y_{i\ell}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma_{Y,i}^{2}(c)),\qquad\ell=1,\dots,m.$
6.2.2 Results for Distributional Data
Across metrics and sample sizes, the empirical size at $c=0$ aligns with the nominal $\alpha=0.05$. In the mean-varying design (DGP 2-1), power increases with $c$. Across all $n$ and $c$ considered, rejection rates under $\mathcal{W}_{1}$, $\mathcal{W}_{2}$, FR, and Hellinger are nearly indistinguishable, with pairwise differences typically $\leq 0.02$ (and at most about $0.04$). See Table 3. In the variance‐varying design (DGP 2-2), 1– and 2–Wasserstein remain closely matched, whereas FR and Hellinger show lower power for $c>0$ with gaps widening as $c$ and $n$ increase. See Table 4. Overall, 1– and 2–Wasserstein perform similarly across designs. The FR and Hellinger distances generate positive-definite distance-induced exponential kernels for distributions in any dimension, but they appear to effectively capture shifts in the mean and are less sensitive to changes in variance.
Table 3: Power of the CI test for the mean-varying setting (DGP 1). The column for $c=0$ shows the empirical Type I error rate (nominal level $\alpha=0.05$), while columns for $c>0$ show statistical power.
Signal Strength $(c)$
$n$ Metric 0.00 0.05 0.10 0.15 0.20
200 1–Wasserstein 0.049 0.050 0.094 0.170 0.299
2–Wasserstein 0.054 0.060 0.093 0.173 0.309
Fisher–Rao 0.045 0.058 0.088 0.165 0.287
Hellinger 0.050 0.062 0.088 0.170 0.300
400 1–Wasserstein 0.052 0.078 0.182 0.425 0.672
2–Wasserstein 0.051 0.090 0.184 0.423 0.682
Fisher–Rao 0.044 0.077 0.178 0.394 0.674
Hellinger 0.053 0.078 0.193 0.404 0.643
600 1–Wasserstein 0.044 0.091 0.282 0.620 0.890
2–Wasserstein 0.046 0.084 0.298 0.635 0.908
Fisher–Rao 0.039 0.080 0.270 0.592 0.895
Hellinger 0.045 0.082 0.267 0.593 0.885
Table 4: Power of the CI test for the variance-varying setting (DGP 2). The column for $c=0$ shows the empirical Type I error rate (nominal level $\alpha=0.05$), while columns for $c>0$ show statistical power.
Signal Strength $(c)$
$n$ Metric 0.00 0.05 0.10 0.15 0.20
200 1–Wasserstein 0.041 0.067 0.121 0.243 0.424
2–Wasserstein 0.044 0.075 0.128 0.246 0.440
Fisher–Rao 0.054 0.061 0.090 0.147 0.272
Hellinger 0.046 0.061 0.090 0.148 0.256
400 1–Wasserstein 0.042 0.088 0.238 0.527 0.774
2–Wasserstein 0.038 0.097 0.279 0.548 0.784
Fisher–Rao 0.045 0.066 0.153 0.353 0.615
Hellinger 0.047 0.066 0.148 0.356 0.598
600 1–Wasserstein 0.043 0.126 0.418 0.761 0.954
2–Wasserstein 0.046 0.133 0.427 0.788 0.960
Fisher–Rao 0.044 0.086 0.246 0.571 0.869
Hellinger 0.038 0.078 0.238 0.567 0.863
7 Conclusion
We have introduced the CI test based on SGCM, which unifies the strengths of kernel and regression-based methods. Through a spectral decomposition based basis expansion, our approach bypasses the notoriously difficult estimation of CMEs, instead reframing the problem as a series of simple, Euclidean-valued regressions. This design not only allows for the flexible integration of modern machine learning regressors but also confers a crucial double robustness property, ensuring reliable performance even in high-dimensional settings.
Theoretically, we established the consistency of our test and its wild bootstrap counterpart, guaranteeing both pointwise and uniform asymptotic size control under doubly robust conditions. A key enabler for our method’s applicability to complex data is our second main contribution: a general theory for constructing characteristic kernels on any Polish space equipped with a conditionally negative definite metric. This result significantly broadens the scope of kernel-based inference for non-Euclidean data, including time-varying random objects and distributions.
Empirical results confirmed the superiority of our method, which demonstrated accurate size control and dominant power over state-of-the-art competitors across a wide array of challenging Euclidean and distribution-valued data scenarios. While computationally intensive, the SGCM framework provides a powerful, versatile, and theoretically sound tool for modern statistical inference.
{acks}
[Acknowledgments] The authors would like to thank …
{funding}
This work was supported by JSPS KAKENHI Grant Numbers JP25K00625 and JP24K02904, and by Nomura Foundation Grant Number N25-4-F30-021.
Supplementary Material for
Testing Conditional Independence via the Spectral Generalized Covariance Measure: Beyond Euclidean Data
Ryunosuke Miyazaki and Yoshimasa Uematsu
Department of Social Data Science, Hitotsubashi University
Throughout this Appendix we work under Assumption 2.1. In particular, we fix the finite constants
$K_{X}\coloneqq\sup_{x\in\mathcal{X}}k_{\mathcal{X}}(x,x),\qquad K_{Y}\coloneqq\sup_{y\in\mathcal{Y}}k_{\mathcal{Y}}(y,y),\qquad K_{Z}\coloneqq\sup_{z\in\mathcal{Z}}k_{\mathcal{Z}}(z,z),$
which serve as uniform bounds for the corresponding feature-map norms. All statements in this Appendix are to be understood under the standing convention that $K_{X},K_{Y},K_{Z}<\infty$. Unless stated otherwise, $\|\cdot\|$ and $\langle\cdot,\cdot\rangle$ denote, respectively, the norm and the inner product of the Hilbert space to which the arguments belong.
Henceforth, write $P_{X},P_{Y},P_{Z}$ for the marginals of $X,Y,Z$ and $P_{XZ},P_{YZ}$ for the joint laws of $(X,Z)$ and $(Y,Z)$. For $1\leq p\leq\infty$, let $L_{X}^{p}$ denote $L^{p}(\mathcal{X},P_{X})$; the spaces $L_{Y}^{p}$, $L_{Z}^{p}$, $L_{XZ}^{p}$, and $L_{YZ}^{p}$ are defined analogously.
We define the canonical feature maps associated with the kernels as $\phi_{\mathcal{X}}(x)\coloneqq k_{\mathcal{X}}(x,\cdot)$, $\phi_{\mathcal{Y}}(y)\coloneqq k_{\mathcal{Y}}(y,\cdot)$, and $\phi_{\mathcal{Z}}(z)\coloneqq k_{\mathcal{Z}}(z,\cdot)$.
We denote by $\mathbb{P}^{*}$, $\operatorname{\mathbb{E}}^{*}$, and $\operatorname{Var}^{*}$ the probability, expectation, and variance conditional on the data $\mathcal{D}_{n}=\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{n}$. The symbol $a_{n}\asymp b_{n}$ means that there exist constants $0<c\leq C<\infty$ such that $c\,b_{n}\leq a_{n}\leq C\,b_{n}$ for sufficiently large $n$.
Appendix A Proofs of Main Theorems of the conditional cross-covariance operator
We write $\ddot{X}\coloneqq(X,Z)$ and $\ddot{\mathcal{X}}=\mathcal{X}\times\mathcal{Z}$. By classical results on product RKHSs, the RKHS $\mathcal{H}_{\ddot{\mathcal{X}}}$ of the product kernel $k_{\ddot{\mathcal{X}}}((x,z),(x^{\prime},z^{\prime}))=k_{\mathcal{X}}(x,x^{\prime})k_{\mathcal{Z}}(z,z^{\prime})$ is canonically isometric to the Hilbert tensor product $\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}$; see, e.g., Steinwart and Christmann (2008, Section 4.2), Paulsen and Raghupathi (2016, Section 5.5).
Proof of Proposition 2.1.
Since the kernels are bounded by Assumption 2.1, all feature maps and RKHS elements are uniformly bounded (Steinwart and Christmann, 2008, Lemma 4.28). In particular, for each kernel $k$ used below we have $\operatorname{\mathbb{E}}\|k(U,\cdot)\|_{\mathcal{H}_{k}}<\infty$ for the corresponding random input $U$, so all Bochner expectations and conditional Bochner expectations appearing in the proof are well defined.
By Lemma A.1 (identifying $\mathcal{H}_{\ddot{\mathcal{X}}}\cong\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}$),
$\operatorname{\mathbb{E}}\left[k_{\ddot{\mathcal{X}}}((X,Z),\cdot)\mid Z\right]=\operatorname{\mathbb{E}}\left[k_{\mathcal{X}}(X,\cdot)\mid Z\right]\otimes k_{\mathcal{Z}}(Z,\cdot)\quad\text{in }\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}\ \text{a.s.}$
Consequently,
$k_{\ddot{\mathcal{X}}}((X,Z),\cdot)-\mu_{\ddot{X}\mid Z}(Z)=\left(k_{\mathcal{X}}(X,\cdot)-\mu_{X\mid Z}(Z)\right)\otimes k_{\mathcal{Z}}(Z,\cdot)\quad\text{a.s.}$ (A.1)
Define
$\Sigma_{\ddot{X}Y\mid Z}\coloneqq\operatorname{\mathbb{E}}\left[\bigl(k_{\ddot{\mathcal{X}}}(\ddot{X},\cdot)-\mu_{\ddot{X}\mid Z}(Z)\bigr)\otimes\bigl(k_{\mathcal{Y}}(Y,\cdot)-\mu_{Y\mid Z}(Z)\bigr)\right]\in\mathcal{H}_{\ddot{\mathcal{X}}}\otimes\mathcal{H}_{\mathcal{Y}}.$
Let us denote the tensor product of the Hilbert spaces by $\mathcal{H}_{XZY}\coloneqq\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}\otimes\mathcal{H}_{\mathcal{Y}}$. By (A.1), $\Sigma_{\ddot{X}Y\mid Z}$ identifies with the tensor
$\widetilde{\Sigma}_{XZY\mid Z}=\operatorname{\mathbb{E}}\left[\left(k_{\mathcal{X}}(X,\cdot)-\mu_{X\mid Z}(Z)\right)\otimes k_{\mathcal{Z}}(Z,\cdot)\otimes\left(k_{\mathcal{Y}}(Y,\cdot)-\mu_{Y\mid Z}(Z)\right)\right]\in\mathcal{H}_{XZY}.$
Let $U_{(23)}$ be the unitary that permutes tensor factors as $U_{(23)}(u\otimes v\otimes w)=u\otimes w\otimes v$. Then
$\Sigma_{XYZ\mid Z}=0\quad\Longleftrightarrow\quad\widetilde{\Sigma}_{XZY\mid Z}=0\quad\Longleftrightarrow\quad\Sigma_{\ddot{X}Y\mid Z}=0,$ (A.2)
since $U_{(23)}$ is unitary and hence injective.
Assuming $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$, fix arbitrary $h_{\mathcal{X}}\in\mathcal{H}_{\mathcal{X}}$, $h_{\mathcal{Y}}\in\mathcal{H}_{\mathcal{Y}}$, $h_{\mathcal{Z}}\in\mathcal{H}_{\mathcal{Z}}$. Since $c(Z)\coloneqq\left\langle\phi_{\mathcal{Z}}(Z),h_{\mathcal{Z}}\right\rangle$ is $\sigma(Z)$-measurable and bounded, Proposition 2.6.31 in Hytönen et al. (2016) and $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$ give
$\operatorname{\mathbb{E}}\left[\left\langle\phi_{\mathcal{X}}(X)-\mu_{X\mid Z}(Z),h_{\mathcal{X}}\right\rangle\left\langle\phi_{\mathcal{Y}}(Y)-\mu_{Y\mid Z}(Z),h_{\mathcal{Y}}\right\rangle\,c(Z)\right]$
$\quad=\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}[\left\langle\phi_{\mathcal{X}}(X)-\mu_{X\mid Z}(Z),h_{\mathcal{X}}\right\rangle\mid Z]\operatorname{\mathbb{E}}[\left\langle\phi_{\mathcal{Y}}(Y)-\mu_{Y\mid Z}(Z),h_{\mathcal{Y}}\right\rangle\mid Z]\,c(Z)\right]=0$
Hence $\left\langle\Sigma_{XYZ\mid Z},\,h_{\mathcal{X}}\otimes h_{\mathcal{Y}}\otimes h_{\mathcal{Z}}\right\rangle=0$ for all elementary tensors. By density of the algebraic tensor product in $\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Y}}\otimes\mathcal{H}_{\mathcal{Z}}$, it follows that $\Sigma_{XYZ\mid Z}=0$.
By (A.2), we have $\Sigma_{\ddot{X}Y\mid Z}=0$. Under Assumption 2.2 (i), both $k_{\ddot{\mathcal{X}}}=k_{\mathcal{X}}\otimes k_{\mathcal{Z}}$ and $k_{\mathcal{Y}}$ are characteristic. Because each kernel is bounded and characteristic by Assumptions 2.1 and 2.2 (i), the direct sum $\mathcal{H}_{\ddot{\mathcal{X}}}+\mathbb{R}$ is dense in $L^{2}(P_{XZ})$ for any probability measure $P_{XZ}$ on $(\ddot{\mathcal{X}},\mathcal{B}(\ddot{\mathcal{X}}))$, and similarly $\mathcal{H}_{\mathcal{Y}}+\mathbb{R}$ is dense in $L^{2}(P_{Y})$ for any probability measure $P_{Y}$ on $(\mathcal{Y},\mathcal{B}(\mathcal{Y}))$ (Fukumizu, Bach and Jordan, 2009, Proposition 5).
Define a bilinear form $B\colon L^{2}_{XZ}\times L^{2}_{Y}\to\mathbb{R}$ by
$B(f,g)\coloneqq\operatorname{\mathbb{E}}\left[(f-\operatorname{\mathbb{E}}[f\mid Z])(g-\operatorname{\mathbb{E}}\left[g\mid Z\right])\right].$
Using the Cauchy–Schwarz inequality and that conditional expectation is an $L^{2}$-contraction, we obtain
$|B(f,g)|\leq\|f-\operatorname{\mathbb{E}}[f\mid Z]\|_{L^{2}}\,\|g-\operatorname{\mathbb{E}}[g\mid Z]\|_{L^{2}}\leq 4\|f\|_{L_{XZ}^{2}}\|g\|_{L_{Y}^{2}},$
confirming the continuity of $B$.
Next, for any $f_{H}\in\mathcal{H}_{\ddot{\mathcal{X}}}$ and $g_{H}\in\mathcal{H}_{\mathcal{Y}}$, expanding the definition of $\Sigma_{\ddot{X}Y\mid Z}$ yields
$\left\langle\Sigma_{\ddot{X}Y\mid Z},f_{H}\otimes g_{H}\right\rangle=\operatorname{\mathbb{E}}\left[(f_{H}(X,Z)-\operatorname{\mathbb{E}}[f_{H}(X,Z)\mid Z])(g_{H}(Y)-\operatorname{\mathbb{E}}[g_{H}(Y)\mid Z])\right]=B(f_{H},g_{H}).$
Thus $\Sigma_{\ddot{X}Y\mid Z}=0$ implies
$B(f_{H},g_{H})=0\quad\forall f_{H}\in\mathcal{H}_{\ddot{\mathcal{X}}},\ g_{H}\in\mathcal{H}_{\mathcal{Y}}.$
Moreover, constants vanish under centering, so $B$ is identically zero on $(\mathcal{H}_{\ddot{\mathcal{X}}}+\mathbb{R})\times(\mathcal{H}_{\mathcal{Y}}+\mathbb{R})$.
By the denseness and continuity of $B$, we conclude that $B(f,g)=0$ for all $f\in L^{2}_{XZ}$ and $g\in L^{2}_{Y}$; that is,
$\operatorname{\mathbb{E}}\left[\left(f(X,Z)-\operatorname{\mathbb{E}}[f(X,Z)\mid Z]\right)\left(g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z]\right)\right]=0\quad\forall f\in L^{2}_{XZ},g\in L^{2}_{Y}.$
Thus, Corollary A.4 implies $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$.
The case of Assumption 2.2 (ii) is identical by symmetry, exchanging the roles of $X$ and $Y$.
Finally, under Assumption 2.2 (iii), we establish the converse. Fix arbitrary $f\in L^{2}(P_{X})$ and $g\in L^{2}(P_{Y})$, and define the centered versions
$\tilde{f}\coloneqq f(X)-\operatorname{\mathbb{E}}[f(X)\mid Z],\qquad\tilde{g}\coloneqq g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z],$
together with the conditional moment
$M_{f,g}(Z)\coloneqq\operatorname{\mathbb{E}}[\tilde{f}\,\tilde{g}\mid Z]\in L^{1}(P_{Z}).$
From $\Sigma_{XYZ\mid Z}=0$ we know that
$\operatorname{\mathbb{E}}\left[\tilde{f}_{H}\,\tilde{g}_{H}\,h_{H}(Z)\right]=0\quad\text{for all }f_{H}\in\mathcal{H}_{\mathcal{X}}+\mathbb{R},\ g_{H}\in\mathcal{H}_{\mathcal{Y}}+\mathbb{R},\ h_{H}\in\mathcal{H}_{\mathcal{Z}}.$
For such $f_{H}\in\mathcal{H}_{\mathcal{X}}+\mathbb{R}$ and $g_{H}\in\mathcal{H}_{\mathcal{Y}}+\mathbb{R}$, bounded kernels imply that $f_{H}(X)$ and $g_{H}(Y)$ are essentially bounded; hence $\tilde{f}_{H}$ and $\tilde{g}_{H}$ are essentially bounded as well, so $\tilde{f}_{H}\,\tilde{g}_{H}\in L^{2}(\mathbb{P})$ and therefore
$M_{f_{H},g_{H}}(Z)\coloneqq\operatorname{\mathbb{E}}\left[\tilde{f}_{H}\,\tilde{g}_{H}\mid Z\right]\in L^{2}(P_{Z}).$
Equivalently,
$\left\langle M_{f_{H},g_{H}},\,h_{H}\right\rangle_{L^{2}(P_{Z})}=0\quad\text{for all }h_{H}\in\mathcal{H}_{\mathcal{Z}}.$
Since $\mathcal{H}_{\mathcal{Z}}$ is dense in $L^{2}(P_{Z})$ by Assumption 2.2 (iii), the identity
$\langle M_{f_{H},g_{H}},h_{H}\rangle_{L^{2}(P_{Z})}=0\quad\forall\,h_{H}\in\mathcal{H}_{\mathcal{Z}}$
implies $M_{f_{H},g_{H}}=0$ in $L^{2}(P_{Z})$ by the self-duality of $L^{2}(P_{Z})$.
By Assumptions 2.1-2.2 (iii) and Fukumizu, Bach and Jordan (2009, Proposition 5),
$\overline{\mathcal{H}_{\mathcal{X}}+\mathbb{R}}^{L^{2}(P_{X})}=L^{2}(P_{X}),\qquad\overline{\mathcal{H}_{\mathcal{Y}}+\mathbb{R}}^{L^{2}(P_{Y})}=L^{2}(P_{Y}).$
Hence there exist sequences $f_{n}\in\mathcal{H}_{\mathcal{X}}+\mathbb{R}$ and $g_{n}\in\mathcal{H}_{\mathcal{Y}}+\mathbb{R}$ with $\|f_{n}-f\|_{L_{X}^{2}}\to 0$ and $\|g_{n}-g\|_{L_{Y}^{2}}\to 0$. Let $\tilde{f}_{n}=f_{n}(X)-\operatorname{\mathbb{E}}[f_{n}(X)\mid Z]$ and $\tilde{g}_{n}=g_{n}(Y)-\operatorname{\mathbb{E}}[g_{n}(Y)\mid Z]$. Since conditional expectation is an $L^{2}$-contraction,
$\|\tilde{f}_{n}-\tilde{f}\|_{L^{2}}\leq 2\|f_{n}-f\|_{L_{X}^{2}}\to 0,\quad\|\tilde{g}_{n}-\tilde{g}\|_{L^{2}}\leq 2\|g_{n}-g\|_{L_{Y}^{2}}\to 0.$
Therefore
$\left\|M_{f_{n},g_{n}}-M_{f,g}\right\|_{L_{Z}^{1}}=\Bigl\|\operatorname{\mathbb{E}}\bigl[(\tilde{f}_{n}\tilde{g}_{n}-\tilde{f}\tilde{g})\mid Z\bigr]\Bigr\|_{L_{Z}^{1}}\leq\operatorname{\mathbb{E}}\bigl|\tilde{f}_{n}\tilde{g}_{n}-\tilde{f}\tilde{g}\bigr|\to 0,$
using Cauchy–Schwarz and the fact that the approximating sequences are bounded in $L^{2}$: indeed, $(f_{n})$ and $(g_{n})$ converge in $L^{2}$, hence are $L^{2}$-bounded, and conditional expectation is a contraction on $L^{2}$, so $(\tilde{f}_{n})$ and $(\tilde{g}_{n})$ are $L^{2}$-bounded as well.
Since $M_{f_{n},g_{n}}=0$ in $L^{2}(P_{Z})$, it follows that $M_{f,g}=0$ in $L^{1}(P_{Z})$ hence also almost surely.
Consequently,
$\operatorname{\mathbb{E}}\left[\bigl(f(X)-\operatorname{\mathbb{E}}[f(X)\mid Z]\bigr)\bigl(g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z]\bigr)\,\bigm|\,Z\right]=0\quad\text{a.s.-}P_{Z}\quad\forall f\in L_{X}^{2},\,g\in L_{Y}^{2}.$ (A.3)
Therefore, by Proposition A.3, (A.3) implies $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$. ∎
Lemma A.1.
Assume Assumptions 2.1, 2.2. Define the product kernel on $\mathcal{X}\times\mathcal{Z}$ by
$k_{\ddot{\mathcal{X}}}((x,z),(x^{\prime},z^{\prime}))\coloneqq k_{\mathcal{X}}(x,x^{\prime})\,k_{\mathcal{Z}}(z,z^{\prime}),$
and denote its RKHS by $\mathcal{H}_{\ddot{\mathcal{X}}}$. Then, as $\mathcal{H}_{\ddot{\mathcal{X}}}$-valued random variables,
$\operatorname{\mathbb{E}}\left[k_{\ddot{\mathcal{X}}}\bigl((X,Z),\cdot\bigr)\mid Z\right]=\operatorname{\mathbb{E}}\left[k_{\mathcal{X}}(X,\cdot)\mid Z\right]\otimes k_{\mathcal{Z}}(Z,\cdot),$
where we identify $\mathcal{H}_{\ddot{\mathcal{X}}}$ isometrically with the Hilbert tensor product $\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}$ and use the elementary tensor on the right-hand side.
Proof.
Let $U\colon\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}\to\mathcal{H}_{\ddot{\mathcal{X}}}$ be the isometric isomorphism specified on elementary tensors by
$U\left(\phi_{\mathcal{X}}(x)\otimes\phi_{\mathcal{Z}}(z)\right)=k_{\ddot{\mathcal{X}}}((x,z),\cdot),$
and extended by linearity and completion. For all $(x,z),(x^{\prime},z^{\prime})$ one has
$\left\langle\phi_{\mathcal{X}}(x)\otimes\phi_{\mathcal{Z}}(z),\phi_{\mathcal{X}}(x^{\prime})\otimes\phi_{\mathcal{Z}}(z^{\prime})\right\rangle=k_{\mathcal{X}}(x,x^{\prime})k_{\mathcal{Z}}(z,z^{\prime})=\left\langle k_{\ddot{\mathcal{X}}}((x,z),\cdot),k_{\ddot{\mathcal{X}}}((x^{\prime},z^{\prime}),\cdot)\right\rangle,$
so $U$ is indeed an isometry. Below we identify $\mathcal{H}_{\ddot{\mathcal{X}}}\cong\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}$ via $U$.
Since the kernels are bounded, $\phi_{\mathcal{X}}(X)\otimes\phi_{\mathcal{Z}}(Z)$ is Bochner integrable and admits a Bochner conditional expectation given $\sigma(Z)$.
Let $\mathcal{D}$ be the algebraic span of simple tensors $h_{\mathcal{X}}\otimes h_{\mathcal{Z}}$ with $h_{\mathcal{X}}\in\mathcal{H}_{\mathcal{X}}$ and $h_{\mathcal{Z}}\in\mathcal{H}_{\mathcal{Z}}$. Then $\mathcal{D}$ is dense in $\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}$. For any $h_{\mathcal{X}}\otimes h_{\mathcal{Z}}\in\mathcal{D}$, by Proposition 2.6.31 in Hytönen et al. (2016) we have
$\left\langle\operatorname{\mathbb{E}}\left[\phi_{\mathcal{X}}(X)\otimes\phi_{\mathcal{Z}}(Z)\mid Z\right],\,h_{\mathcal{X}}\otimes h_{\mathcal{Z}}\right\rangle=\operatorname{\mathbb{E}}\left[\langle\phi_{\mathcal{X}}(X),h_{\mathcal{X}}\rangle\left\langle\phi_{\mathcal{Z}}(Z),h_{\mathcal{Z}}\right\rangle\mid Z\right].$
Since $\left\langle\phi_{\mathcal{Z}}(Z),h_{\mathcal{Z}}\right\rangle$ is $\sigma(Z)$-measurable and bounded, we have
$\operatorname{\mathbb{E}}\left[\langle\phi_{\mathcal{X}}(X),h_{\mathcal{X}}\rangle\left\langle\phi_{\mathcal{Z}}(Z),h_{\mathcal{Z}}\right\rangle\mid Z\right]=\left\langle\phi_{\mathcal{Z}}(Z),h_{\mathcal{Z}}\right\rangle\,\operatorname{\mathbb{E}}\left[\langle\phi_{\mathcal{X}}(X),h_{\mathcal{X}}\rangle\mid Z\right].$
On the other hand,
$\left\langle\operatorname{\mathbb{E}}\left[\phi_{\mathcal{X}}(X)\mid Z\right]\otimes\phi_{\mathcal{Z}}(Z),\,h_{\mathcal{X}}\otimes h_{\mathcal{Z}}\right\rangle$ $=\left\langle\operatorname{\mathbb{E}}\left[\phi_{\mathcal{X}}(X)\mid Z\right],h_{\mathcal{X}}\right\rangle\,\left\langle\phi_{\mathcal{Z}}(Z),h_{\mathcal{Z}}\right\rangle$
$=\operatorname{\mathbb{E}}\left[\langle\phi_{\mathcal{X}}(X),h_{\mathcal{X}}\rangle\mid Z\right]\,\left\langle\phi_{\mathcal{Z}}(Z),h_{\mathcal{Z}}\right\rangle.$
Therefore, for all $h_{\mathcal{X}}\otimes h_{\mathcal{Z}}\in\mathcal{D}$,
$\left\langle\operatorname{\mathbb{E}}\left[\phi_{\mathcal{X}}(X)\otimes\phi_{\mathcal{Z}}(Z)\mid Z\right],\,h_{\mathcal{X}}\otimes h_{\mathcal{Z}}\right\rangle=\left\langle\operatorname{\mathbb{E}}\left[\phi_{\mathcal{X}}(X)\mid Z\right]\otimes\phi_{\mathcal{Z}}(Z),\,h_{\mathcal{X}}\otimes h_{\mathcal{Z}}\right\rangle\quad\text{a.s.}$
By density of $\mathcal{D}$ and continuity of the inner product, it follows that
$\operatorname{\mathbb{E}}\left[\phi_{\mathcal{X}}(X)\otimes\phi_{\mathcal{Z}}(Z)\mid Z\right]=\operatorname{\mathbb{E}}\left[\phi_{\mathcal{X}}(X)\mid Z\right]\otimes\phi_{\mathcal{Z}}(Z)\quad\text{almost surely in }\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Z}}.$
Applying the isometry $U$ and using $\phi_{\mathcal{X}}(x)=k_{\mathcal{X}}(x,\cdot)$ and $\phi_{\mathcal{Z}}(z)=k_{\mathcal{Z}}(z,\cdot)$ yields the claimed identity in $\mathcal{H}_{\ddot{\mathcal{X}}}$. ∎
We show that the CCCO defined via CMEs coincides, under a suitable identification, with the classical CCCO introduced by Fukumizu, Bach and Jordan (2004). The classical definition was formulated without CMEs as an RKHS analog of the conditional covariance for Gaussian random variables.
Proposition A.2.
Assume Assumptions 2.1 and 2.2. Let the product kernel on $\ddot{\mathcal{Y}}=\mathcal{Y}\times\mathcal{Z}$ be $k_{\ddot{\mathcal{Y}}}((y,z),(y^{\prime},z^{\prime}))\coloneqq k_{\mathcal{Y}}(y,y^{\prime})\,k_{\mathcal{Z}}(z,z^{\prime})$ with RKHS $\mathcal{H}_{\ddot{\mathcal{Y}}}$, and let $U_{YZ}\colon\mathcal{H}_{\mathcal{Y}}\otimes\mathcal{H}_{\mathcal{Z}}\xrightarrow{\ \cong\ }\mathcal{H}_{\ddot{\mathcal{Y}}}$ be the canonical isometry.
Define the conditional cross-covariance operator $\Sigma_{X\ddot{Y}\mid Z}\in\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\ddot{\mathcal{Y}}}$ by
$\Sigma_{X\ddot{Y}\mid Z}=\Sigma_{X\ddot{Y}}-\Sigma_{XX}^{1/2}V_{XZ}V_{Z\ddot{Y}}\Sigma_{\ddot{Y}\ddot{Y}}^{1/2}.$
Then the canonical product-kernel identification yields
$\Sigma_{X\ddot{Y}\mid Z}=\left(I\otimes U_{YZ}\right)\,\Sigma_{XYZ\mid Z}\in\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\ddot{\mathcal{Y}}}.$ (A.4)
In particular,
$\|\Sigma_{X\ddot{Y}\mid Z}\|_{\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\ddot{\mathcal{Y}}}}=\|\Sigma_{XYZ\mid Z}\|_{\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Y}}\otimes\mathcal{H}_{\mathcal{Z}}}$
Proof.
Fix $f_{\mathcal{X}}\in\mathcal{H}_{\mathcal{X}}$ and write $h_{\ddot{Y}}=U_{YZ}\left(h_{\mathcal{Y}}\otimes h_{\mathcal{Z}}\right)$ with $h_{\mathcal{Y}}\in\mathcal{H}_{\mathcal{Y}}$, $h_{\mathcal{Z}}\in\mathcal{H}_{\mathcal{Z}}$. By Fukumizu, Bach and Jordan (2004, Proposition 5) applied to $\ddot{Y}=(Y,Z)$,
$\left\langle f_{\mathcal{X}}\otimes h_{\ddot{Y}},\,\Sigma_{X\ddot{Y}\mid Z}\right\rangle=\operatorname{\mathbb{E}}\left[\operatorname{Cov}\left(\left\langle f_{\mathcal{X}},k_{\mathcal{X}}(X,\cdot)\right\rangle,\ \left\langle h_{\ddot{Y}},k_{\ddot{\mathcal{Y}}}((Y,Z),\cdot)\right\rangle\ \middle|\ Z\right)\right].$
Using $U_{YZ}$ and the product-kernel evaluation,
$\left\langle h_{\ddot{Y}},k_{\ddot{\mathcal{Y}}}((Y,Z),\cdot)\right\rangle=\left\langle h_{\mathcal{Y}},k_{\mathcal{Y}}(Y,\cdot)\right\rangle\,\left\langle h_{\mathcal{Z}},k_{\mathcal{Z}}(Z,\cdot)\right\rangle.$
By Park and Muandet (2020, Lemma 2),
$\left\langle f_{\mathcal{X}}\otimes h_{\mathcal{Y}}\otimes h_{\mathcal{Z}},\Sigma_{XYZ\mid Z}\right\rangle=\operatorname{\mathbb{E}}\left[\operatorname{Cov}\left(\left\langle f_{\mathcal{X}},k_{\mathcal{X}}(X,\cdot)\right\rangle,\left\langle h_{\mathcal{Y}},k_{\mathcal{Y}}(Y,\cdot)\right\rangle\mid Z\right)\left\langle h_{\mathcal{Z}},k_{\mathcal{Z}}(Z,\cdot)\right\rangle\right].$
Combining the last two displays gives
$\left\langle f_{\mathcal{X}}\otimes h_{\ddot{Y}},\,\Sigma_{X\ddot{Y}\mid Z}\right\rangle=\left\langle f_{\mathcal{X}}\otimes h_{\ddot{Y}},\ \left(I\otimes U_{YZ}\right)\Sigma_{XYZ\mid Z}\right\rangle.$
Since elementary tensors are total, (A.4) follows. ∎
A.1 Daudin’s Characterization of CI on Polish Spaces
Here, $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$ is defined via factorization of regular conditional probabilities: there exists a $P_{Z}$-null set $N\subset\mathcal{Z}$ such that, for all $z\in\mathcal{Z}\setminus N$ and all $A\in\mathcal{B}(\mathcal{X})$, $B\in\mathcal{B}(\mathcal{Y})$,
$\mathbb{P}(X\in A,\,Y\in B\mid Z=z)=\mathbb{P}(X\in A\mid Z=z)\,\mathbb{P}(Y\in B\mid Z=z).$
In addition to this measure-theoretic definition, we provide below a characterization of CI by Daudin (1980) in terms of vanishing conditional covariances of centered square-integrable functions. While Daudin (1980) established such a characterization for real-valued random variables, the equivalence with the above conditional-measure factorization was not stated explicitly there; for completeness, we state and prove the equivalence here in the present Polish setting.
Proposition A.3.
Let $(\mathcal{X},\mathcal{Y},\mathcal{Z})$ be Polish spaces, and let $X\colon\Omega\to\mathcal{X}$, $Y\colon\Omega\to\mathcal{Y}$, $Z\colon\Omega\to\mathcal{Z}$ be random elements on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. Assume that for all $f\in L^{2}_{X}$ and $g\in L^{2}_{Y}$,
$\operatorname{\mathbb{E}}\left[\left(f(X)-\operatorname{\mathbb{E}}[f(X)\mid Z]\right)\left(g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z]\right)\mid Z\right]=0\quad\text{a.s.-}P_{Z}.$
Then $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$.
Proof of Proposition A.3.
Taking indicators $f=\mathds{1}_{A}$, $g=\mathds{1}_{B}$ with $A\in\mathcal{B}(\mathcal{X})$, $B\in\mathcal{B}(\mathcal{Y})$ (these belong to $L^{2}$ on a probability space), we obtain almost surely
$\mathbb{P}(X\in A,\,Y\in B\mid Z)=\mathbb{P}(X\in A\mid Z)\,\mathbb{P}(Y\in B\mid Z).$
Taking indicators establishes the factorization for any given pair of Borel sets $(A,B)$ almost surely. However, the definition of CI requires this identity to hold simultaneously for all $A\in\mathcal{B}(\mathcal{X})$ and $B\in\mathcal{B}(\mathcal{Y})$ outside a single null set. To bridge this gap from a pointwise to a simultaneous almost-sure statement, we employ a standard measure-theoretic argument. Let $\mathcal{P}_{\mathcal{X}}\subset\mathcal{B}(\mathcal{X})$ and $\mathcal{P}_{\mathcal{Y}}\subset\mathcal{B}(\mathcal{Y})$ be countable $\pi$-systems that generate the respective Borel $\sigma$-algebras (e.g. finite intersections of balls with rational centers and radii). For each $A\in\mathcal{P}_{\mathcal{X}},B\in\mathcal{P}_{\mathcal{Y}}$, the above factorization holds almost surely; denote by $N_{A,B}$ the corresponding $P_{Z}$-null set where it may fail. Let $N\coloneqq\bigcup_{A,B}N_{A,B}$, still $P_{Z}$-null. Fix $z\in\mathcal{Z}\setminus N$. The proof proceeds in two steps by applying the $\pi$-$\lambda$ theorem twice.
First, fix an arbitrary set $A\in\mathcal{P}_{\mathcal{X}}$. We define a class of sets $\Lambda_{A}\subset\mathcal{B}(\mathcal{Y})$ as follows:
$\Lambda_{A}\coloneqq\{B\in\mathcal{B}(\mathcal{Y}):P_{XY\mid Z=z}(A\times B)=P_{X\mid Z=z}(A)P_{Y\mid Z=z}(B)\}.$
It is a standard exercise to verify that $\Lambda_{A}$ is a $\lambda$-system. By hypothesis, the product identity holds for all sets in the $\pi$-system $\mathcal{P}_{\mathcal{Y}}$, which implies that $\mathcal{P}_{\mathcal{Y}}\subset\Lambda_{A}$. Since $\Lambda_{A}$ is a $\lambda$-system containing the $\pi$-system $\mathcal{P}_{\mathcal{Y}}$ that generates $\mathcal{B}(\mathcal{Y})$, the $\pi$-$\lambda$ theorem (Klenke, 2020, Theorem 1.19) yields that $\sigma(\mathcal{P}_{\mathcal{Y}})\subset\Lambda_{A}$, and thus $\Lambda_{A}=\mathcal{B}(\mathcal{Y})$.
Second, having shown that the equality holds for any $A\in\mathcal{P}_{\mathcal{X}}$ and all $B\in\mathcal{B}(\mathcal{Y})$, we now extend it to all $A\in\mathcal{B}(\mathcal{X})$. Define the class $\Lambda\subset\mathcal{B}(\mathcal{X})$ by:
$\Lambda\coloneqq\{A\in\mathcal{B}(\mathcal{X}):P_{XY\mid Z=z}(A\times B)=P_{X\mid Z=z}(A)P_{Y\mid Z=z}(B)\text{ for all }B\in\mathcal{B}(\mathcal{Y})\}.$
Again, $\Lambda$ is a $\lambda$-system. From the conclusion of our first step, we know that $\mathcal{P}_{\mathcal{X}}\subset\Lambda$. Since $\Lambda$ is a $\lambda$-system containing the $\pi$-system $\mathcal{P}_{\mathcal{X}}$ that generates $\mathcal{B}(\mathcal{X})$, we may once more apply the $\pi$-$\lambda$ theorem to conclude that $\Lambda=\mathcal{B}(\mathcal{X})$.
Therefore, the identity $P_{XY\mid Z=z}(A\times B)=P_{X\mid Z=z}(A)P_{Y\mid Z=z}(B)$ holds for all $A\in\mathcal{B}(\mathcal{X})$ and $B\in\mathcal{B}(\mathcal{Y})$. This is precisely the statement that the conditional measure $P_{XY\mid Z=z}$ is equal to the product of its marginals $P_{X\mid Z=z}\otimes P_{Y\mid Z=z}$. As this holds for all $z\in\mathcal{Z}\setminus N$ where $P_{Z}(N)=0$, we have shown $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$. ∎
Corollary A.4.
Under the setting of Proposition A.3, suppose that for every $f\in L^{2}_{XZ}$ and $g\in L^{2}_{Y}$,
$\operatorname{\mathbb{E}}\left[\left(f(X,Z)-\operatorname{\mathbb{E}}[f(X,Z)\mid Z]\right)\left(g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z]\right)\mid Z\right]=0\quad\text{a.s.-}P_{Z}.$
Then $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$.
Proof of Corollary A.4.
Work under the setting of the preceding proposition. Fix $\tilde{f}\in L^{2}_{X}$, $g\in L^{2}_{Y}$, and $h\in L^{\infty}_{Z}$. First, $f(x,z)\coloneqq\tilde{f}(x)\,h(z)$ belongs to $L^{2}_{XZ}$, since
$\operatorname{\mathbb{E}}\left[\bigl|\tilde{f}(X)\,h(Z)\bigr|^{2}\right]\leq\left\|h\right\|_{\infty}^{2}\,\operatorname{\mathbb{E}}\left[\tilde{f}(X)^{2}\right]<\infty.$
Moreover, by the tower property and the fact that $h(Z)$ is $\sigma(Z)$-measurable and bounded,
$f(X,Z)-\operatorname{\mathbb{E}}[f(X,Z)\mid Z]=h(Z)\,\Bigl(\tilde{f}(X)-\operatorname{\mathbb{E}}\bigl[\tilde{f}(X)\,\mid Z\bigr]\Bigr).$
Applying the hypothesis with this $f$ and the given $g$ yields, almost surely,
$0=\operatorname{\mathbb{E}}\left[\left(f(X,Z)-\operatorname{\mathbb{E}}[f(X,Z)\mid Z]\right)\left(g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z]\right)\mid Z\right]=h(Z)\,\Phi(Z),$
where
$\Phi(Z)\coloneqq\operatorname{\mathbb{E}}\left[\bigl(\tilde{f}(X)-\operatorname{\mathbb{E}}\bigl[\tilde{f}(X)\mid Z\bigr]\bigr)\bigl(g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z]\bigr)\mid Z\right]$
is $\sigma(Z)$-measurable and integrable.
Since the equality $h(Z)\,\Phi(Z)=0$ holds for all $h\in L^{\infty}_{Z}$, it holds in particular for all indicator functions $h=\mathds{1}_{C}$ with $C\in\mathcal{B}(\mathcal{Z})$. Hence $\mathds{1}_{C}(Z)\,\Phi(Z)=0$ a.s. for every Borel set $C$, which implies $\Phi(Z)=0$ a.s.-$P_{Z}$ (take $C=\{\Phi>0\}$ and $C=\{\Phi<0\}$, both $Z$-measurable). Consequently, for every $\tilde{f}\in L^{2}_{X}$ and $g\in L^{2}_{Y}$,
$\operatorname{\mathbb{E}}\left[\bigl(\tilde{f}(X)-\operatorname{\mathbb{E}}\bigl[\tilde{f}(X)\mid Z\bigr]\bigr)\bigl(g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z]\bigr)\mid Z\right]=0\quad\text{a.s.-}P_{Z}.$
Thus the assumption of Proposition A.3 is satisfied with test functions depending only on $X$ (for $\tilde{f}$) and on $Y$ (for $g$), and the proposition yields $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$. ∎
Appendix B Asymptotic Properties of the Test Statistic
In the sequel, we define the random element $\Psi_{i}$ in $\mathcal{H}_{XYZ}=\mathcal{H}_{\mathcal{X}}\otimes\mathcal{H}_{\mathcal{Y}}\otimes\mathcal{H}_{\mathcal{Z}}$ as
$\Psi_{i}\coloneqq\left(k_{\mathcal{X}}(X_{i},\cdot)-\mu_{X\mid Z}(Z_{i})\right)\otimes\left(k_{\mathcal{Y}}(Y_{i},\cdot)-\mu_{Y\mid Z}(Z_{i})\right)\otimes\phi_{\mathcal{Z}}(Z_{i}).$
Moreover, $\hat{\Psi}_{i}^{(P_{n},Q_{n})}$, is then given by
$\hat{\Psi}_{i}^{(P_{n},Q_{n})}\coloneqq\hat{u}^{(i)}_{X}\otimes\hat{u}^{(i)}_{Y}\otimes\phi_{\mathcal{Z}}(Z_{i}),$
where
$\hat{u}^{(i)}_{X}\coloneqq\sum_{p=1}^{P_{n}}\hat{\zeta}^{(i)}_{X\mid Z,p}\,\hat{e}_{p}\in\mathcal{H}_{\mathcal{X}},\qquad\hat{u}^{(i)}_{Y}\coloneqq\sum_{q=1}^{Q_{n}}\hat{\zeta}^{(i)}_{Y\mid Z,q}\,\hat{f}_{q}\in\mathcal{H}_{\mathcal{Y}}.$
For $p\in\{1,\dots,P_{n}\}$ and $q\in\{1,\dots,Q_{n}\}$ define
$\zeta^{(i)}_{X\mid Z,p}$ $\coloneqq\hat{e}_{p}(X_{i})-\operatorname{\mathbb{E}}[\hat{e}_{p}(X_{i})\mid\hat{e}_{p},Z_{i}],$ $\zeta^{(i)}_{Y\mid Z,q}$ $\coloneqq\hat{f}_{q}(Y_{i})-\operatorname{\mathbb{E}}[\hat{f}_{q}(Y_{i})\mid\hat{f}_{q},Z_{i}].$
Let the uncentered covariance operators $C_{XX}$ and $C_{YY}$ admit spectral decompositions
$C_{XX}e_{p}=\lambda_{p}e_{p}\quad\text{in}\ \mathcal{H}_{\mathcal{X}},\qquad C_{YY}f_{q}=\mu_{q}f_{q}\quad\text{in}\ \mathcal{H}_{\mathcal{Y}},$
where $\{e_{p}\}_{p\geq 1}\subset\mathcal{H}_{\mathcal{X}}$ and $\{f_{q}\}_{q\geq 1}\subset\mathcal{H}_{\mathcal{Y}}$ form orthonormal systems of eigenfunctions, and the corresponding eigenvalues are arranged in nonincreasing order. We define the population spectral projectors as
$P_{X,P_{n}}\coloneqq\sum_{p=1}^{P_{n}}e_{p}\otimes e_{p},\qquad P_{Y,Q_{n}}\coloneqq\sum_{q=1}^{Q_{n}}f_{q}\otimes f_{q}.$
Similarly, let $\hat{P}_{X,P_{n}}$ and $\hat{P}_{Y,Q_{n}}$ be the empirical orthogonal projectors on $\mathcal{H}_{\mathcal{X}}$ and $\mathcal{H}_{\mathcal{Y}}$ constructed from the empirical eigenfunctions $\hat{e}_{p}$ and $\hat{f}_{q}$:
$\hat{P}_{X,P_{n}}\coloneqq\sum_{p=1}^{P_{n}}\hat{e}_{p}\otimes\hat{e}_{p},\qquad\hat{P}_{Y,Q_{n}}\coloneqq\sum_{q=1}^{Q_{n}}\hat{f}_{q}\otimes\hat{f}_{q}.$
B.1 Pointwise Asymptotic Properties
Proof of Theorem 4.1.
Define
$T_{n_{1}}\coloneqq\|\Sigma_{n_{1}}\|^{2},\qquad\Sigma_{n_{1}}\coloneqq\frac{1}{n_{1}}\sum_{i\in I_{1}}\Psi_{i},$
and write $V_{i}\coloneqq(X_{i},Y_{i},Z_{i})$. Under $H_{0}$, Lemmas B.4, B.5, and B.6 yield
$n_{1}\left|\hat{T}_{n_{1}}-T_{n_{1}}\right|=o_{p}(1).$
Hence it suffices to analyze the scaled quadratic functional $n_{1}T_{n_{1}}$.
Define the symmetric kernel
$h(V_{i},V_{j})\coloneqq\langle\Psi_{i},\Psi_{j}\rangle=k_{\mathcal{Z}}(Z_{i},Z_{j})\,\left\langle\epsilon_{X\mid Z}^{(i)},\epsilon_{X\mid Z}^{(j)}\right\rangle\,\left\langle\epsilon_{Y\mid Z}^{(i)},\epsilon_{Y\mid Z}^{(j)}\right\rangle,$
where $\epsilon_{X\mid Z}^{(i)}\coloneqq k_{\mathcal{X}}(X_{i},\cdot)-\mu_{X\mid Z}(Z_{i})$ and $\epsilon_{Y\mid Z}^{(i)}\coloneqq k_{\mathcal{Y}}(Y_{i},\cdot)-\mu_{Y\mid Z}(Z_{i})$. Then
$n_{1}T_{n_{1}}=\frac{1}{n_{1}}\sum_{i,j\in I_{1}}h(V_{i},V_{j}),$
so $n_{1}T_{n_{1}}$ is a degree-2 $V$-statistic with kernel $h$. By the Cauchy–Schwarz inequality, $\left|k_{\mathcal{Z}}(Z_{i},Z_{j})\right|\leq K_{Z}$. Moreover,
$\left\|\epsilon_{X\mid Z}^{(i)}\right\|\leq\left\|k_{\mathcal{X}}(X_{i},\cdot)\right\|+\left\|\mu_{X\mid Z}(Z_{i})\right\|\leq\sqrt{K_{X}}+\sqrt{K_{X}}=2\sqrt{K_{X}},$
and similarly $\left\|\epsilon_{Y\mid Z}^{(i)}\right\|\leq 2\sqrt{K_{Y}}$. Hence, we have
$|h(V_{i},V_{j})|\leq K_{Z}\cdot\left\|\epsilon_{X\mid Z}^{(i)}\right\|\left\|\epsilon_{X\mid Z}^{(j)}\right\|\left\|\epsilon_{Y\mid Z}^{(i)}\right\|\left\|\epsilon_{Y\mid Z}^{(j)}\right\|\leq 16\,K_{X}K_{Y}K_{Z},$
so $h$ is uniformly bounded.
Thus, under $H_{0}$ and by first-order degeneracy and boundedness (e.g., Lee 2019, Section 3.2.2, Theorem 1),
$n_{1}\hat{T}_{n_{1}}\overset{d}{\longrightarrow}\sum_{\ell=1}^{\infty}\xi_{\ell}G_{\ell}^{2},\qquad\{G_{\ell}\}_{\ell\geq 1}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,1).$
By Lemma B.9, the weights $\{\xi_{\ell}\}$ are exactly the nonzero eigenvalues of
$C\coloneqq\operatorname{\mathbb{E}}[\Psi\otimes\Psi]\colon\mathcal{H}_{XYZ}\to\mathcal{H}_{XYZ},$
listed in nonincreasing order and repeated according to multiplicity.
Divergence under $H_{1}$ using $\left|\hat{T}_{n_{1}}-T_{n_{1}}\right|=o_{p}(1)$. Under $H_{1}$, by the Hilbert-space LLN, $\Sigma_{n_{1}}\xrightarrow{p}\Sigma_{XYZ\mid Z}$ in $\mathcal{H}_{XYZ}$. By continuity of the Hilbert–Schmidt norm and Assumption 2.2,
$T_{n_{1}}=\|\Sigma_{n_{1}}\|^{2}\overset{p}{\longrightarrow}\left\|\Sigma_{XYZ\mid Z}\right\|^{2}\eqqcolon\theta>0.$
Assumptions 2.1 and 4.1, 4.2 imply $\left|\hat{T}_{n_{1}}-T_{n_{1}}\right|=o_{p}(1)$. Fix $\varepsilon\in(0,\theta/2)$. Then for all sufficiently large $n_{1}$,
$\mathbb{P}\left(T_{n_{1}}>\theta-\varepsilon\right)\to 1,\qquad\mathbb{P}\left(\left|\hat{T}_{n_{1}}-T_{n_{1}}\right|<\varepsilon\right)\to 1,$
so on an event whose probability tends to one, $\hat{T}_{n_{1}}\geq T_{n_{1}}-\varepsilon>\theta-2\varepsilon\geq\theta/2$. Therefore $n_{1}\hat{T}_{n_{1}}\geq n_{1}\,\theta/2\to\infty$ in probability. ∎
Proof of Theorem 4.2.
Let $\zeta_{X\mid Z,p}^{(i)}$ and $\zeta_{Y\mid Z,q}^{(i)}$ be the population residual coordinates and write $\hat{\zeta}=\zeta-b$ with estimation errors $b^{(i)}_{X,p},b^{(i)}_{Y,q}$.
Introduce the population-residual bootstrap counterpart
$n_{1}T_{n_{1}}^{\ast}\coloneqq\frac{1}{n_{1}}\sum_{i,j\in I_{1}}\sum_{p=1}^{P_{n}}\sum_{q=1}^{Q_{n}}k_{\mathcal{Z}}(Z_{i},Z_{j})\,W_{i}^{\ast}W_{j}^{\ast}\,\zeta^{(i)}_{X\mid Z,p}\,\zeta^{(i)}_{Y\mid Z,q}\,\zeta^{(j)}_{X\mid Z,p}\,\zeta^{(j)}_{Y\mid Z,q}$.
Conditioning on the data and expanding $\hat{\zeta}=\zeta-b$, the same estimation-error bounds used in Proposition B.4 give
$\operatorname{\mathbb{E}}^{\ast}\left[\left|n_{1}\hat{T}_{n_{1}}^{\ast}-n_{1}T_{n_{1}}^{\ast}\right|\right]\leq C\left\{\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{p=1}^{P_{n}}\left(b^{(i)}_{X,p}\right)^{2}+\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{q=1}^{Q_{n}}\left(b^{(i)}_{Y,q}\right)^{2}\right\}.$
By Assumptions 4.3, the right-hand side is $o_{p}(1)$, hence
$\left|n_{1}\hat{T}_{n_{1}}^{\ast}-n_{1}T_{n_{1}}^{\ast}\right|=o_{p}^{\ast}(1).$ (A.5)
Define
$S_{n_{1}}^{\ast}\coloneqq\frac{1}{\sqrt{n_{1}}}\sum_{i\in I_{1}}W_{i}^{\ast}\Psi_{i},\qquad A\coloneqq P_{X,P_{n}}\otimes P_{Y,Q_{n}}\otimes I,\quad\hat{A}\coloneqq\hat{P}_{X,P_{n}}\otimes\hat{P}_{Y,Q_{n}}\otimes I,$
where $I$ denotes the identity operator on $\mathcal{H}_{\mathcal{Z}}$. A direct calculation shows
$n_{1}T_{n_{1}}^{\ast}=\left\|A\,S_{n_{1}}^{\ast}\right\|^{2},\qquad n_{1}\bar{T}_{n_{1}}^{\ast}\coloneqq\left\|\hat{A}\,S_{n_{1}}^{\ast}\right\|^{2}.$
By the projector-consistency argument in Lemma B.5 (see (A.20)) and noting that $\|P_{X,P_{n}}\|_{\textrm{op}}=\|\hat{P}_{X,P_{n}}\|_{\textrm{op}}=1$ (and similarly for $Y$), we have
$\|\hat{A}-A\|_{\textrm{op}}$ $\leq\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\|_{\textrm{op}}+\|\hat{P}_{Y,Q_{n}}-P_{Y,Q_{n}}\|_{\textrm{op}}$ (A.6)
$\quad+\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\|_{\textrm{op}}\,\|\hat{P}_{Y,Q_{n}}-P_{Y,Q_{n}}\|_{\textrm{op}}=o_{p}(1).$
Moreover, since $\operatorname{\mathbb{E}}^{\ast}[W_{i}^{\ast}]=0$, $\operatorname{\mathbb{E}}^{\ast}[(W_{i}^{\ast})^{2}]=1$, bounded kernels imply $\sup_{i}\|\Psi_{i}\|^{2}\leq C<\infty$ a.s. (Lemma B.11 and B.12), and $\operatorname{\mathbb{E}}^{\ast}[(W_{i}^{\ast})^{4}]<\infty$, a standard multiplier-moment bound yields
$\operatorname{\mathbb{E}}^{\ast}\left[\left\|S_{n_{1}}^{\ast}\right\|^{4}\right]\leq\frac{C}{n_{1}^{2}}\sum_{i\in I_{1}}\|\Psi_{i}\|^{4}\ +\ \frac{C}{n_{1}^{2}}\left(\sum_{i\in I_{1}}\|\Psi_{i}\|^{2}\right)^{2}=O_{p}(1),$ (A.7)
whence $\|S_{n_{1}}^{\ast}\|^{2}=O_{p}^{\ast}(1)$ by Chebyshev. Combining (A.6) and (A.7),
$\left|\,n_{1}\bar{T}_{n_{1}}^{\ast}-n_{1}T_{n_{1}}^{\ast}\,\right|=\left|\left\|\hat{A}S_{n_{1}}^{\ast}\right\|^{2}-\left\|AS_{n_{1}}^{\ast}\right\|^{2}\right|\leq\|\hat{A}-A\|_{\textrm{op}}\,\|S_{n_{1}}^{\ast}\|^{2}=o_{p}^{\ast}(1).$ (A.8)
Let $T_{n_{1}}^{\ast,\infty}$ denote the version of $T_{n_{1}}^{\ast}$ with the infinite population projectors $\Pi_{X},\Pi_{Y}$ in place of $P_{X,P_{n}},P_{Y,Q_{n}}$. Thus, by Lemma B.6, $T_{n_{1}}^{\ast,\infty}=\left\|\sum_{i\in I_{1}}\Psi_{i}W_{i}^{\ast}\right\|^{2}$. By the same contraction and tail arguments (now conditional on the data and using $\sup_{i}\|\Psi_{i}\|^{2}\leq C$ a.s.), one obtains
$\left|n_{1}T_{n_{1}}^{\ast}-n_{1}T_{n_{1}}^{\ast,\infty}\right|=o_{p}^{\ast}(1).$ (A.9)
Here, the standard wild-bootstrap theory for first-order degenerate $V$-statistics applies; see, e.g., Lee, Zhang and Shao (2020, Theorem 4). It follows that, under $H_{0}$,
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}^{\ast}\left(n_{1}T_{n_{1}}^{\ast,\infty}\leq t\right)-\mathbb{P}(W\leq t)\right|\overset{p}{\longrightarrow}0,$ (A.10)
where $W=\sum_{\ell=1}^{\infty}\xi_{\ell}G_{\ell}^{2}$ is the limit of $n_{1}\hat{T}_{n_{1}}$.
By (A.5), (A.8), (A.9), we have
$\left|\,n_{1}\hat{T}_{n_{1}}^{\ast}-n_{1}T_{n_{1}}^{\ast,\infty}\,\right|=o_{p}^{\ast}(1).$
Therefore, for every continuity point $t$ of $F_{W}$,
$\mathbb{P}^{\ast}\left(n_{1}\hat{T}_{n_{1}}^{\ast}\leq t\right)=\mathbb{P}^{\ast}\left(n_{1}T_{n_{1}}^{\ast,\infty}\leq t\right)+o_{p}(1).$
Proof of Corollary 4.3.
Let
$F_{n}^{\ast}(t)\coloneqq\mathbb{P}^{\ast}\left(n_{1}\hat{T}_{n_{1}}^{\ast}\leq t\right),\qquad\hat{F}_{n,B}^{\ast}(t)\coloneqq\frac{1}{B}\sum_{b=1}^{B}\mathds{1}\left\{n_{1}\hat{T}_{n_{1}}^{\ast(b)}\leq t\right\}$
denote, respectively, the ideal bootstrap distribution function and its empirical version based on $B$ bootstrap replicates. Let $W$ be the limiting distribution obtained in Theorem 4.1.
By the conditional Dvoretzky–Kiefer–Wolfowitz inequality (Massart, 1990), if $B\to\infty$ then
$\sup_{t\in\mathbb{R}}\bigl|\hat{F}_{n,B}^{\ast}(t)-F_{n}^{\ast}(t)\bigr|=o_{p}(1).$
Therefore, replacing the empirical bootstrap cdf $\hat{F}_{n,B}^{\ast}$ by the ideal bootstrap cdf $F_{n}^{\ast}$ does not affect the asymptotic level result. In what follows, we work directly with $F_{n}^{\ast}$ for simplicity. The same reduction applies to part (ii). As $W$ is an infinite weighted sum of chi-squared variables with strictly positive weights, its cdf $F_{W}$ is continuous on $\mathbb{R}$.
By Theorem 4.2, $F_{n}^{\ast}(t)\xrightarrow{p}F_{W}(t)$ for every $t\in\mathbb{R}$. The continuity of $F_{W}$ allows us to apply Pólya’s theorem (Lehmann and Romano, 2022, Theorem 11.2.9), which strengthens the pointwise convergence to uniform convergence in probability:
$\sup_{t\in\mathbb{R}}\,\bigl\lvert F_{n}^{\ast}(t)-F_{W}(t)\bigr\rvert\overset{p}{\longrightarrow}0.$
Let $Q_{1-\alpha}=F_{W}^{-1}(1-\alpha)$ be the $(1-\alpha)$-quantile of $F_{W}$, and let $Q_{n,1-\alpha}^{\ast}=(F_{n}^{*})^{-1}(1-\alpha)$ be the bootstrap quantile. By Lemma 11.2.1(ii) of Lehmann and Romano (2022), the uniform convergence of distribution functions implies the convergence of the corresponding quantiles in probability:
$Q_{n,1-\alpha}^{\ast}\overset{p}{\longrightarrow}Q_{1-\alpha}.$ (A.11)
We have established two key convergences: $n_{1}\hat{T}_{n_{1}}\xrightarrow{d}W$ from Theorem 4.3, and $Q_{n,1-\alpha}^{\ast}\overset{p}{\longrightarrow}Q_{1-\alpha}$ from (A.11). Since the limiting distribution of $n\hat{T}_{n}$ has a continuous distribution function $F_{W}$, an application of Slutsky’s Theorem directly yields:
$\lim_{n\to\infty}\mathbb{P}\left(n_{1}\hat{T}_{n_{1}}>Q_{n,1-\alpha}^{\ast}\right)=\mathbb{P}\left(W>Q_{1-\alpha}\right).$
By the definition of the quantile $Q_{1-\alpha}$ and the continuity of $F_{W}$,
$\mathbb{P}(W>Q_{1-\alpha})=1-\mathbb{P}(W\leq Q_{1-\alpha})=1-F_{W}(Q_{1-\alpha})=1-(1-\alpha)=\alpha.$
Next, we show that the latter statement. Since the kernel is bounded by Assumption 2.1., the bootstrap variance satisfies
$\operatorname{Var}^{\ast}\left(n_{1}\hat{T}_{n_{1}}^{\ast}\right)\leq C^{2},$
and Chebyshev’s inequality yields, for any $K>0$,
$\mathbb{P}^{\ast}\left(\lvert n_{1}\hat{T}_{n_{1}}^{\ast}\rvert>K\right)\leq\frac{C^{2}}{K^{2}}.$
Choosing $K>C/\sqrt{\alpha}$ gives
$\mathbb{P}^{\ast}\left(n_{1}\hat{T}_{n_{1}}^{\ast}\leq K\right)\geq 1-\alpha$
so that $Q_{n,1-\alpha}^{\ast}\leq K$ with unconditional probability at least $1-\alpha$; in particular, $Q_{n,1-\alpha}^{\ast}=O_{p}(1)$.
For any $\varepsilon\in(0,\alpha)$ pick the same $K$ as above; then, for all sufficiently large $n$,
$\mathbb{P}\left(Q_{n,1-\alpha}^{\ast}\leq K\right)\geq 1-\varepsilon.$
Combining this bound with Theorem 4.4 gives
$\mathbb{P}\left(n_{1}\hat{T}_{n_{1}}>Q_{n,1-\alpha}^{\ast}\right)\geq\mathbb{P}\left(n_{1}\hat{T}_{n_{1}}>K\right)-\varepsilon\longrightarrow 1-\varepsilon.$
Because $\varepsilon>0$ is arbitrary, we conclude
$\mathbb{P}\left(n_{1}\hat{T}_{n_{1}}>Q_{n,1-\alpha}^{\ast}\right)\longrightarrow 1.$
B.2 Uniformly Asymptotic Properties
Following Lundborg, Shah and Peters (2022), we say that $W_{n}$ converges uniformly in distribution to $W$ over a subclass $\mathcal{A}\subseteq\mathcal{P}$ and write $W_{n}\xrightarrow{\mathcal{D}}W$ uniformly over $\mathcal{A}$ if
$\lim_{n\to\infty}\ \sup_{P\in\mathcal{A}}d_{\mathrm{BL}}\left(\mathcal{L}_{P}(W_{n}),\,\mathcal{L}_{P}(W)\right)=0,$
where $d_{\mathrm{BL}}$ denotes the bounded–Lipschitz metric and $\mathcal{L}_{P}(\cdot)$ the law under $\mathbb{P}_{P}$. Furthermore, for the Gaussian measures on Hilbert spaces used in the following proof, we also refer to Lundborg, Shah and Peters (2022).
Proof of Theorem 4.4.
First, by Proposition B.4, and Lemmas B.5 and B.7, the biases of conditional expectations and the truncations are asymptotically negligible uniformly:
$n_{1}\left|\hat{T}_{n_{1}}-\|\Sigma_{n_{1}}\|^{2}\right|=o_{\tilde{\mathcal{P}}_{0}}(1).$
Thus, by Lemma S5 in Lundborg et al. (2024), we only consider
$n_{1}\|\Sigma_{n_{1}}\|^{2}=\left\|\frac{1}{\sqrt{n_{1}}}\sum_{i\in I_{1}}\Psi_{i}\right\|$
By Theorem 2 in Lundborg, Shah and Peters (2022),
$\frac{1}{\sqrt{n_{1}}}\sum_{i\in I_{1}}\Psi_{i}\overset{\mathcal{D}}{\to}G_{P},$
uniformly in $\tilde{\mathcal{P}}_{0}$, where $G_{P}$ is a zero-mean Gaussian measure on the Hilbert space $\mathcal{H}_{XYZ}$ with covariance operator $C_{P}$ given by
$C_{P}\coloneqq\operatorname{\mathbb{E}}_{P}\left[\Psi\otimes\Psi\right]\in\mathcal{H}_{XYZ}\otimes\mathcal{H}_{XYZ}.$
Here, we can show that the conditions in Theorem 2 of Lundborg, Shah and Peters (2022) are satisfied under our Assumptions. First, under Assumptions 2.1 and 4.7, condition (iii) of Theorem 2 in Lundborg, Shah and Peters (2022) holds in our setting. Next, we show that the uniform tail condition required in Theorem 2 (iv) also holds. For the triple tensor case, let us define
$U_{p}\coloneqq\tilde{e}_{p}(X)-\operatorname{\mathbb{E}}\bigl[\tilde{e}_{p}(X)\mid Z\bigr],\quad V_{q}\coloneqq\tilde{f}_{q}(Y)-\operatorname{\mathbb{E}}\bigl[\tilde{f}_{q}(Y)\mid Z\bigr],\quad W_{r}\coloneqq\tilde{g}_{r}(Z).$
Then, we have
$\sup_{P}\sum_{p+q+r\geq K}\operatorname{\mathbb{E}}_{P}[U_{p}^{2}V_{q}^{2}W_{r}^{2}]$
$\leq 4K_{Y}K_{Z}\sup_{P}\sum_{p\geq K/3}\operatorname{\mathbb{E}}_{P}[U_{p}^{2}]+4K_{X}K_{Z}\sup_{P}\sum_{q\geq K/3}\operatorname{\mathbb{E}}_{P}[V_{q}^{2}]+16K_{X}K_{Y}\sup_{P}\sum_{r\geq K/3}\operatorname{\mathbb{E}}_{P}[W_{r}^{2}],$
which vanishes as $K\to\infty$ under Assumption 4.7. Hence, the condition holds.
Now, we want to show that
$\mathbb{P}_{P}(\phi_{n,P,B}=1)=\mathbb{P}_{P}\left(n_{1}T_{n_{1}}>Q_{n,1-\alpha}^{\ast(B)}\right)$
converges to $\alpha$ uniformly, so it suffices to show that for any $\varepsilon>0$ there exists $N\in\mathbb{N}$, such that for all $n\geq N$,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(n_{1}T_{n_{1}}>Q_{n,1-\alpha}^{\ast(B)}\right)<\alpha+\varepsilon$
and
$\inf_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(n_{1}T_{n_{1}}>Q_{n,1-\alpha}^{\ast(B)}\right)>\alpha-\varepsilon$ (A.12)
Fix an arbitrary $\delta>0$ (its value will be specified later). Whenever $\left|Q_{n,1-\alpha}^{\ast(B)}-Q_{1-\alpha}(C_{P})\right|<\delta$ and $T_{n}>Q_{n,1-\alpha}^{\ast(B)}$, it follows that $T_{n}>Q_{1-\alpha}(C_{P})-\delta$. Hence
$\mathbb{P}_{P}\left(T_{n}>Q_{n,1-\alpha}^{\ast(B)}\right)\leq\mathbb{P}_{P}\left(T_{n}>Q_{1-\alpha}(C_{P})-\delta\right)+\mathbb{P}_{P}\left(\left|Q_{n,1-\alpha}^{\ast(B)}-Q_{1-\alpha}(C_{P})\right|\geq\delta\right).$
Taking the supremum over $P\in\tilde{\mathcal{P}}_{0}$ yields
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(T_{n}>Q_{n,1-\alpha}^{\ast(B)}\right)\leq\mathrm{I}_{n}+\mathrm{I}\hskip-1.2pt\mathrm{I}_{n},$
where
$\mathrm{I}_{n}\coloneqq\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(T_{n}>Q_{1-\alpha}(C_{P})-\delta\right),\quad\mathrm{I}\hskip-1.2pt\mathrm{I}_{n}\coloneqq\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left|Q_{n,1-\alpha}^{\ast(B)}-Q_{1-\alpha}(C_{P})\right|\geq\delta\right).$
Add and subtract the Gaussian tail to obtain
$\mathrm{I}_{n}$ $=\sup_{P\in\tilde{\mathcal{P}}_{0}}\Bigl[\mathbb{P}_{P}\left(T_{n}>Q_{1-\alpha}(C_{P})-\delta\right)-\mathbb{P}\left(W>Q_{1-\alpha}(C_{P})-\delta\right)\Bigr]$
$\quad+\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}\left(W>Q_{1-\alpha}(C_{P})-\delta\right)-\alpha$
$\quad+\alpha$
$\eqqcolon\mathrm{I}\hskip-1.2pt\mathrm{I}\hskip-1.2pt\mathrm{I}_{n}+\mathrm{I}\hskip-1.2pt\mathrm{V}_{n}+\alpha.$
We will prove that, for all sufficiently large $n$,
$\mathrm{I}\hskip-1.2pt\mathrm{I}_{n}<\tfrac{\varepsilon}{3},\quad\mathrm{I}\hskip-1.2pt\mathrm{I}\hskip-1.2pt\mathrm{I}_{n}<\tfrac{\varepsilon}{3},\quad\mathrm{I}\hskip-1.2pt\mathrm{V}_{n}<\tfrac{\varepsilon}{3}.$
Combining these bounds, we obtain
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(nT_{n}>Q_{1-\alpha,n}^{\ast(B)}\right)\leq\alpha+\varepsilon,$
which is precisely the desired uniform upper bound on the rejection probability. The corresponding lower bound (A.12) follows by an analogous argument.
For $\mathrm{I}\hskip-1.2pt\mathrm{I}_{n}$, by Proposition B.1 and Lemma B.13, we have
$\mathrm{I}\hskip-1.2pt\mathrm{I}_{n}<\tfrac{\varepsilon}{3}.$
For $\mathrm{I}\hskip-1.2pt\mathrm{I}\hskip-1.2pt\mathrm{I}_{n}$ and $\mathrm{I}\hskip-1.2pt\mathrm{V}_{n}$, it can be proved by the same argument in the proof of Theorem 3 in Lundborg, Shah and Peters (2022). An analogous argument yields (A.12). ∎
Proposition B.1.
Fix $\alpha\in(0,1)$ and assume Assumptions 2.1 and that there exists $c_{0}>0$ such that
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\hat{C}_{n}-C_{P}\right\|_{\operatorname{tr}}>\varepsilon\right)\xrightarrow[n\to\infty]{}0\qquad\text{for all }\varepsilon>0.$ (A.13)
For a positive trace-class operator $C$, define the $(1-\alpha)$-quantile of $\left\|\mathcal{N}(0,C)\right\|_{\mathrm{HS}}^{2}$ by
$Q_{1-\alpha}(C)\coloneqq\inf\left\{t\in\mathbb{R}\,\colon\,\mathbb{P}\left(\|\mathcal{N}(0,C)\|_{\mathrm{HS}}^{2}\leq t\right)\geq 1-\alpha\right\}.$
Then, as $n,B\to\infty$,
$\left|Q_{n,1-\alpha}^{\ast(B)}-Q_{1-\alpha}(C_{P})\right|=o_{\tilde{\mathcal{P}}_{0}}(1).$
Proof.
For any cumulative distribution function $F$ on $\mathbb{R}$ and $\tau\in(0,1)$, define the (left-continuous) quantile map
$Q_{\tau}[F]\coloneqq\inf\left\{t\in\mathbb{R}\colon\ F(t)\geq\tau\right\}.$
For a positive trace-class operator $C$, let $F_{C}$ denote the law of $\left\|\mathcal{N}(0,C)\right\|_{\textrm{HS}}^{2}$ and define
$Q_{\tau}(C)\coloneqq Q_{\tau}[F_{C}].$
Given the data $\mathcal{D}_{n}\coloneqq\{X_{i},Y_{i},Z_{i}\}_{i=1}^{n}$, write $F_{n}\coloneqq F_{\hat{C}_{n}}$ for the conditional law of $\left\|N\left(0,\hat{C}_{n}\right)\right\|_{\textrm{HS}}^{2}$. Let $Z_{1}^{\ast},\dots,Z_{B}^{\ast}$ be the bootstrap draws produced by the multiplier scheme with i.i.d. standard normal multipliers; conditionally on $\mathcal{D}_{n}$, they are i.i.d. with cdf $F_{n}$. Define the empirical bootstrap cdf by
$F_{n,B}^{\ast}(t)\coloneqq\frac{1}{B}\sum_{b=1}^{B}\mathds{1}\left\{\,Z_{b}^{\ast}\leq t\,\right\},$
and define the empirical bootstrap quantile at level $\tau$ by
$Q_{\tau}^{\ast,B}(\hat{C}_{n})\coloneqq Q_{\tau}\left[F_{n,B}^{\ast}\right].$
In particular, at level $1-\alpha$, since the multipliers are standard normal and hence $Z_{b}^{\ast}$ are conditionally i.i.d. from $F_{n}$,
$Q_{1-\alpha}^{\ast(B)}=Q_{1-\alpha}^{\ast,B}(\hat{C}_{n})=Q_{1-\alpha}\left[F_{n,B}^{\ast}\right].$
By Assumption 2.1, there exists $c_{1}<\infty$ such that
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\operatorname{tr}(C_{P})\leq c_{1}.$ (A.14)
Moreover, by Assumption 4.8, there exists $\lambda_{\star}>0$ with
$\inf_{P\in\tilde{\mathcal{P}}_{0}}\ \lambda_{\max}(C_{P})\geq\lambda_{\star}.$ (A.15)
By (A.14) and (A.13), for large $n$,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\operatorname{tr}(\hat{C}_{n})\leq 2c_{1}\right)\to 1,$
since $\left|\operatorname{tr}(\hat{C}_{n})-\operatorname{tr}(C_{P})\right|\leq\left\|\hat{C}_{n}-C_{P}\right\|_{\operatorname{tr}}$. Further, by Weyl’s inequality and (A.15),
$\lambda_{\max}(\hat{C}_{n})\geq\lambda_{\max}(C_{P})-\left\|\hat{C}_{n}-C_{P}\right\|_{\textrm{op}}\geq\lambda_{\star}-\left\|\hat{C}_{n}-C_{P}\right\|_{\operatorname{tr}},$
hence, again by (A.13),
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\lambda_{\max}(\hat{C}_{n})\geq\lambda_{\star}/2\ \right)\to 1.$
Define the compact quantile band $\mathcal{T}=[\tau_{-},\tau_{+}]\subset(0,1)$ and the deterministic constants
$a_{\mathcal{T}}\coloneqq\frac{\lambda_{\star}}{2}\,Q_{\tau_{-}}\left(\chi^{2}_{1}\right),\qquad b_{\mathcal{T}}\coloneqq\frac{2c_{1}}{1-\tau_{+}},\qquad\eta_{\mathcal{T}}\coloneqq\min\left\{\frac{\lambda_{\star}}{4},\ \frac{c_{1}}{2},\ \frac{a_{\mathcal{T}}}{16}\right\}.$
Introduce the event, for each $P\in\tilde{\mathcal{P}}$,
$\mathcal{E}_{n}(P)\coloneqq\left\{\left\|\hat{C}_{n}-C_{P}\right\|_{\operatorname{tr}}\leq\eta_{\mathcal{T}}\ \text{ and }\ \sum_{k>K(\eta_{\mathcal{T}})}\lambda_{k}(C_{P})\leq\eta_{\mathcal{T}}\right\}.$
By Assumption 4.7 together with (A.13),
$\rho_{n}\coloneqq\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\mathcal{E}_{n}(P)^{\mathrm{c}}\right)\ \longrightarrow\ 0.$
Fix $n$ and work on $\mathcal{E}_{n}(P)$. We have
$\sum_{k>K(\eta_{\mathcal{T}})}\hat{\lambda}_{k}\leq\sum_{k>K(\eta_{\mathcal{T}})}\lambda_{k}(C_{P})\ +\ \sum_{k\geq 1}\left|\hat{\lambda}_{k}-\lambda_{k}(C_{P})\right|\leq 2\eta_{\mathcal{T}}\leq\frac{a_{\mathcal{T}}}{8}.$
By Lemma B.2, on the event $\mathcal{E}_{n}(P)$ one has $Q_{\tau}(\hat{C}_{n})\in\left[a_{\mathcal{T}},\,b_{\mathcal{T}}\right]$ for all $\tau\in\mathcal{T}$, and there exist $r_{0}=a_{\mathcal{T}}/4$ and $m_{0}>0$ depending only on $\lambda_{\star},c_{1},\mathcal{T}$ such that the conditional density $f_{n}$ of $Z_{n}$ satisfies $f_{n}(x)\geq m_{0}$ for all $x\in\left[Q_{\tau}(\hat{C}_{n})-r_{0},\,Q_{\tau}(\hat{C}_{n})+r_{0}\right]$ uniformly in $\tau\in\mathcal{T}$. Given $\mathcal{D}_{n}$, let $F_{n,B}^{\ast}$ be the empirical cdf of i.i.d. draws $Z_{1}^{\ast},\dots,Z_{B}^{\ast}\sim F_{n}$, and set $Q_{\tau}^{\ast,B}(\hat{C}_{n})\coloneqq Q_{\tau}\left(F_{n,B}^{\ast}\right)$. By the Dvoretzky–Kiefer–Wolfowitz inequality (Massart, 1990),
$\mathbb{P}^{\ast}(\ \left\|F_{n,B}^{\ast}-F_{n}\right\|_{\infty}>\delta\ )\leq 2\,\exp(-2B\delta^{2})\qquad\text{for all }\delta>0.$
By Lemma B.3, on $\mathcal{E}_{n}(P)$ the quantile map is locally $1/m_{0}$-Lipschitz in the $\ell_{\infty}$ distance, hence
$\mathbb{P}^{\ast}\left(\sup_{\tau\in\mathcal{T}}\left|Q_{\tau}^{\ast,B}(\hat{C}_{n})-Q_{\tau}(\hat{C}_{n})\right|>\varepsilon\right)\leq\mathbb{P}^{\ast}\left(\left\|F_{n,B}^{\ast}-F_{n}\right\|_{\infty}>m_{0}\varepsilon\right)\leq 2\exp\left(-2Bm_{0}^{2}\varepsilon^{2}\right)$
for any $\varepsilon\in\left(0,r_{0}\right)$. Taking expectation with respect to $\mathbb{P}_{P}$ and using $\mathds{1}_{\mathcal{E}_{n}(P)}\leq 1$ gives
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\ \sup_{\tau\in\mathcal{T}}\left|\,Q_{\tau}^{\ast,B}(\hat{C}_{n})-Q_{\tau}(\hat{C}_{n})\,\right|>\varepsilon\ \right)\leq\rho_{n}\ +\ 2\,\exp\left(-2B\,m_{0}^{2}\varepsilon^{2}\right),$
which yields the desired uniform convergence in probability first as $B\to\infty$ for fixed $n$, and then as $n\to\infty$ because $\rho_{n}\to 0$.
Define
$\mathcal{K}_{\operatorname{tr}}\coloneqq\left\{\,C\in\mathfrak{S}_{1}^{+}(\mathcal{H})\ :\ \operatorname{tr}(C)\leq 2c_{1}\,\right\},$
where $\mathfrak{S}_{1}^{+}(\mathcal{H})$ denotes the cone of positive trace-class operators on $\mathcal{H}$, i.e.,
$\mathfrak{S}_{1}^{+}(\mathcal{H})=\left\{C\in\mathcal{B}(\mathcal{H})\ :\ C=C^{\ast},\ \left\langle Cf,f\right\rangle_{\mathcal{H}}\geq 0\ \text{for all }f\in\mathcal{H},\ \operatorname{tr}(C)<\infty\,\right\},$
where $\mathcal{B}(\mathcal{H})$ denotes the bounded linear operators on $\mathcal{H}$. In particular, $\mathcal{K}_{\operatorname{tr}}$ is trace-bounded because $\operatorname{tr}(C)\leq 2c_{1}$ for all $C\in\mathcal{K}_{\operatorname{tr}}$.
Lemma 12 in Lundborg, Shah and Peters (2022) implies that $C\mapsto Q_{1-\alpha}(C)$ is uniformly continuous on $\mathcal{K}_{\operatorname{tr}}$ with respect to the trace norm. Combining this with (A.13),
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\left|Q_{1-\alpha}(\hat{C}_{n})-Q_{1-\alpha}(C_{P})\right|>\varepsilon\right)\ \xrightarrow[n\to\infty]{}0.$ (A.16)
For any $\varepsilon>0$,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\left|Q_{n,1-\alpha}^{\ast(B)}-Q_{1-\alpha}(C_{P})\right|>\varepsilon\right)$ $\leq\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\left|Q_{n,1-\alpha}^{\ast(B)}-Q_{1-\alpha}(\hat{C}_{n})\right|>\varepsilon/2\right)$
$\quad+\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left|Q_{1-\alpha}(\hat{C}_{n})-Q_{1-\alpha}(C_{P})\right|>\varepsilon/2\right).$
The first term is bounded by $\rho_{n}+o_{B}(1)$ by the preceding display; the second term vanishes by (A.16). Therefore
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\left|Q_{n,1-\alpha}^{\ast(B)}-Q_{1-\alpha}(C_{P})\right|>\varepsilon\right)\longrightarrow 0,$
that is, $\left|Q_{n,1-\alpha}^{\ast(B)}-Q_{1-\alpha}(C_{P})\right|=o_{\tilde{\mathcal{P}}_{0}}(1)$. ∎
Lemma B.2.
On $\mathcal{E}_{n}(P)$, for all $\tau\in\mathcal{T}$,
$Q_{\tau}(\hat{C}_{n})\in\left[a_{\mathcal{T}},\,b_{\mathcal{T}}\right],\quad f_{n}(x)\geq m_{0}~~\text{for all}~x\in\left[Q_{\tau}(\hat{C}_{n})-r_{0},\,Q_{\tau}(\hat{C}_{n})+r_{0}\right],$
with $r_{0}=a_{\mathcal{T}}/4$ and some $m_{0}>0$ depending only on $\lambda_{\star},c_{1},\mathcal{T}$.
Proof.
Fix $P\in\tilde{\mathcal{P}}$ and work on the event $\mathcal{E}_{n}(P)$. Since $\|\cdot\|_{\mathrm{op}}\leq\|\cdot\|_{\mathrm{tr}}$ and $\eta_{\mathcal{T}}\leq\lambda_{\star}/4$, we have
$\|\hat{C}_{n}-C_{P}\|_{\mathrm{op}}\leq\eta_{\mathcal{T}}\leq\frac{\lambda_{\star}}{4}.$
By Weyl’s inequality for self-adjoint operators,
$\lambda_{\max}(\hat{C}_{n})\geq\lambda_{\max}(C_{P})-\|\hat{C}_{n}-C_{P}\|_{\mathrm{op}}\geq\lambda_{\star}-\frac{\lambda_{\star}}{4}\geq\frac{\lambda_{\star}}{2}.$
Moreover, using $|\operatorname{tr}(A)-\operatorname{tr}(B)|\leq\|A-B\|_{\mathrm{tr}}$ and $\eta_{\mathcal{T}}\leq c_{1}/2$,
$\operatorname{tr}(\hat{C}_{n})\leq\operatorname{tr}(C_{P})+\|\hat{C}_{n}-C_{P}\|_{\mathrm{tr}}\leq c_{1}+\frac{c_{1}}{2}\leq 2c_{1}.$
Let $(\hat{\lambda}_{k})_{k\geq 1}$ denote the eigenvalues of $\hat{C}_{n}$ in non-increasing order, and let $(\chi^{2}_{1,k})_{k\geq 1}$ be i.i.d. $\chi^{2}_{1}$ random variables, independent of $\hat{C}_{n}$. Conditionally on $\hat{C}_{n}$, the random variable
$Z_{n}\coloneqq\sum_{k\geq 1}\hat{\lambda}_{k}\,\chi^{2}_{1,k}$
satisfies $Z_{n}\geq\hat{\lambda}_{1}\chi^{2}_{1,1}=\lambda_{\max}(\hat{C}_{n})\chi^{2}_{1,1}$ almost surely. Hence, for any $\tau\in\mathcal{T}$, by monotonicity of quantiles and the scaling property $Q_{\tau}(aX)=aQ_{\tau}(X)$ for $a>0$,
$Q_{\tau}(\hat{C}_{n})=Q_{\tau}(Z_{n}\mid\hat{C}_{n})\geq\frac{\lambda_{\star}}{2}\,Q_{\tau}(\chi^{2}_{1})\geq\frac{\lambda_{\star}}{2}\,Q_{\tau_{-}}(\chi^{2}_{1})=a_{\mathcal{T}}.$
On the other hand, $Z_{n}\geq 0$ and $\operatorname{\mathbb{E}}[Z_{n}\mid\hat{C}_{n}]=\operatorname{tr}(\hat{C}_{n})\leq 2c_{1}$, so Markov’s inequality implies
$Q_{\tau}(\hat{C}_{n})\leq\frac{\operatorname{\mathbb{E}}[Z_{n}\mid\hat{C}_{n}]}{1-\tau}\leq\frac{2c_{1}}{1-\tau}\leq\frac{2c_{1}}{1-\tau_{+}}=b_{\mathcal{T}}.$
Therefore $Q_{\tau}(\hat{C}_{n})\in[a_{\mathcal{T}},b_{\mathcal{T}}]$ for all $\tau\in\mathcal{T}$.
Let $K\coloneqq K(\eta_{\mathcal{T}})$ and $S\coloneqq\{1,\dots,K\}$. Decompose
$Z_{n}=U_{S}+R,\qquad U_{S}\coloneqq\sum_{k\in S}\hat{\lambda}_{k}\,\chi^{2}_{1,k},\qquad R\coloneqq\sum_{k\notin S}\hat{\lambda}_{k}\,\chi^{2}_{1,k}.$
Then $U_{S}$ and $R$ are independent. On $\mathcal{E}_{n}(P)$ we have $\sum_{k>K}\hat{\lambda}_{k}\leq a_{\mathcal{T}}/8$ (as established in the main text from the definition of $\mathcal{E}_{n}(P)$ and the choice $\eta_{\mathcal{T}}\leq a_{\mathcal{T}}/16$), hence
$\operatorname{\mathbb{E}}[R\mid\hat{C}_{n}]=\sum_{k\notin S}\hat{\lambda}_{k}\leq\frac{a_{\mathcal{T}}}{8}.$
With $r_{0}\coloneqq a_{\mathcal{T}}/4$, Markov’s inequality yields
$\mathbb{P}(R\leq r_{0}\mid\hat{C}_{n})\geq 1-\frac{\operatorname{\mathbb{E}}[R\mid\hat{C}_{n}]}{r_{0}}\geq\frac{1}{2}.$
We next obtain a uniform lower bound for the density of $U_{S}$ on a fixed compact interval. For $\hat{\lambda}>0$ define
$h_{\hat{\lambda}}(u)\coloneqq\frac{1}{\hat{\lambda}}\,f_{\chi^{2}_{1}}\left(\frac{u}{\hat{\lambda}}\right)=\frac{1}{\sqrt{2\pi\hat{\lambda}u}}\exp\left(-\frac{u}{2\hat{\lambda}}\right),\qquad u>0,$ (A.17)
and set $h_{\hat{\lambda}}(u)\coloneqq 0$ for $u\leq 0$. Let $I_{x}\coloneqq[a_{\mathcal{T}}/4,\,2b_{\mathcal{T}}]$. Write $U_{S}=X+V$ with
$X\coloneqq\hat{\lambda}_{1}\chi^{2}_{1,1},\qquad V\coloneqq\sum_{k=2}^{K}\hat{\lambda}_{k}\,\chi^{2}_{1,k}.$
Conditionally on $\hat{C}_{n}$, $X$ and $V$ are independent, and $X$ has density $h_{\hat{\lambda}_{1}}$. Hence $U_{S}$ admits a density $g_{S}(\cdot;\hat{\lambda}_{1:K})$ given by the convolution formula
$g_{S}(x;\hat{\lambda}_{1:K})=\operatorname{\mathbb{E}}\left[h_{\hat{\lambda}_{1}}(x-V)\,\middle|\,\hat{C}_{n}\right],\qquad x\in\mathbb{R},$ (A.18)
where $h_{\hat{\lambda}_{1}}(x-V)=0$ whenever $x\leq V$.
On $\mathcal{E}_{n}(P)$, we have $\hat{\lambda}_{1}\in[\lambda_{\star}/2,\,2c_{1}]$ and $\hat{\lambda}_{k}\leq\operatorname{tr}(\hat{C}_{n})\leq 2c_{1}$ for all $k\leq K$. Define the constant
$\underline{h}\coloneqq\inf_{\lambda\in[\lambda_{\star}/2,\,2c_{1}]}\ \inf_{u\in[a_{\mathcal{T}}/8,\,2b_{\mathcal{T}}]}\ h_{\lambda}(u).$
Since $(\lambda,u)\mapsto h_{\lambda}(u)$ is continuous and strictly positive on the compact set $\lambda_{\star}/2,2c_{1}]\times[a_{\mathcal{T}}/8,2b_{\mathcal{T}}]$, we have $\underline{h}>0$.
For $x\in I_{x}$, consider the event $A\coloneqq\{V\leq a_{\mathcal{T}}/8\}$. On $A$ we have $x-V\in[a_{\mathcal{T}}/8,\,2b_{\mathcal{T}}]$, and therefore $h_{\hat{\lambda}_{1}}(x-V)\geq\underline{h}$. Using (A.18),
$g_{S}(x;\hat{\lambda}_{1:K})\geq\underline{h}\,\mathbb{P}(A\mid\hat{C}_{n}).$
To bound $\mathbb{P}(A\mid\hat{C}_{n})$ from below uniformly, note that
$\bigcap_{k=2}^{K}\left\{\hat{\lambda}_{k}\chi^{2}_{1,k}\leq\frac{a_{\mathcal{T}}}{8(K-1)}\right\}\subseteq\left\{V\leq\frac{a_{\mathcal{T}}}{8}\right\}=A.$
By independence of $(\chi^{2}_{1,k})_{k=2}^{K}$ and the bound $\hat{\lambda}_{k}\leq 2c_{1}$,
$\mathbb{P}(A\mid\hat{C}_{n})\geq\prod_{k=2}^{K}\mathbb{P}\left(\chi^{2}_{1}\leq\frac{a_{\mathcal{T}}}{16c_{1}(K-1)}\right)=p_{0}^{K-1},\quad p_{0}\coloneqq\mathbb{P}\left(\chi^{2}_{1}\leq\frac{a_{\mathcal{T}}}{16c_{1}(K-1)}\right)>0.$
Consequently, for all $x\in I_{x}$,
$g_{S}(x;\hat{\lambda}_{1:K})\geq\underline{h}\,p_{0}^{K-1}\eqqcolon m_{S},$
where $m_{S}>0$ depends only on $\lambda_{\star},c_{1},\mathcal{T}$ (through $a_{\mathcal{T}},b_{\mathcal{T}},\eta_{\mathcal{T}}$ and the deterministic index $K(\eta_{\mathcal{T}})$).
Let $f_{n}$ denote the conditional density of $Z_{n}$ given $\hat{C}_{n}$. Since $Z_{n}=U_{S}+R$ with $U_{S}$ independent of $R$ and $U_{S}$ having density $g_{S}(\cdot;\hat{\lambda}_{1:K})$, we have, for all $x\in\mathbb{R}$,
$f_{n}(x)=\int_{[0,\infty)}g_{S}(x-v;\hat{\lambda}_{1:K})\,F_{R}(\mathop{}\!\mathrm{d}v),$
where $F_{R}$ is the conditional distribution function of $R$ given $\hat{C}_{n}$. For any $x\in[a_{\mathcal{T}}/2,\,2b_{\mathcal{T}}]$ and any $v\in[0,r_{0}]$,
$x-v\in\left[a_{\mathcal{T}}/2-r_{0},\,2b_{\mathcal{T}}\right]=\left[a_{\mathcal{T}}/4,\,2b_{\mathcal{T}}\right]=I_{x},$
so $g_{S}(x-v;\hat{\lambda}_{1:K})\geq m_{S}$. Therefore,
$f_{n}(x)\geq\int_{[0,r_{0}]}g_{S}(x-v;\hat{\lambda}_{1:K})\,F_{R}(\mathop{}\!\mathrm{d}v)\geq m_{S}\,\mathbb{P}(R\leq r_{0}\mid\hat{C}_{n})\geq\frac{m_{S}}{2}\eqqcolon m_{0}.$
Finally, for any $\tau\in\mathcal{T}$, the already established bounds $Q_{\tau}(\hat{C}_{n})\in[a_{\mathcal{T}},b_{\mathcal{T}}]$ imply
$Q_{\tau}(\hat{C}_{n})-r_{0}\geq a_{\mathcal{T}}-\frac{a_{\mathcal{T}}}{4}=\frac{a_{\mathcal{T}}}{2},\qquad Q_{\tau}(\hat{C}_{n})+r_{0}\leq b_{\mathcal{T}}+\frac{a_{\mathcal{T}}}{4}\leq 2b_{\mathcal{T}},$
and hence
$\left[Q_{\tau}(\hat{C}_{n})-r_{0},\ Q_{\tau}(\hat{C}_{n})+r_{0}\right]\subseteq\left[\frac{a_{\mathcal{T}}}{2},\ 2b_{\mathcal{T}}\right].$
Thus $f_{n}(x)\geq m_{0}$ holds for all $x$ in the above interval, uniformly over $\tau\in\mathcal{T}$, as claimed. ∎
Lemma B.3.
Let $F,G$ be cdfs on $\mathbb{R}$ and $\tau\in(0,1)$. Suppose $F$ admits a density $f$ with $f(x)\geq m_{0}>0$ for all $x\in[q-r_{0},q+r_{0}]$, where $q\coloneqq Q_{\tau}\left(F\right)$. If $\left\|G-F\right\|_{\infty}\leq\Delta<m_{0}r_{0}$, then
$\left|Q_{\tau}(G)-Q_{\tau}\left(F\right)\right|\leq\frac{\Delta}{m_{0}}.$
Proof.
Set $q\coloneqq Q_{\tau}(F)$. Since $F$ admits a density $f$ on $q-r_{0},q+r_{0}]$ with $f\geq m_{0}>0$ there, $F$ is continuous and strictly increasing on $q-r_{0},q+r_{0}]$. In particular, $F(q)=\tau$ (otherwise, if $F(q)>\tau$, continuity would yield some $t<q$ with $F(t)\geq\tau$, contradicting the definition of $q$).
Fix $\Delta$ with $\|G-F\|_{\infty}\leq\Delta<m_{0}r_{0}$ and define $h\coloneqq\Delta/m_{0}\in(0,r_{0})$. The lower bound on $f$ yields
$F(q+h)-F(q)=\int_{q}^{q+h}f(u)\,\mathop{}\!\mathrm{d}u\geq m_{0}h=\Delta,\qquad F(q)-F(q-h)=\int_{q-h}^{q}f(u)\,\mathop{}\!\mathrm{d}u\geq m_{0}h=\Delta.$
Since $F(q)=\tau$, we obtain
$F(q+h)\geq\tau+\Delta,\qquad F(q-h)\leq\tau-\Delta.$
Using $\|G-F\|_{\infty}\leq\Delta$, it follows that
$G(q+h)\geq F(q+h)-\Delta\geq\tau,$
and therefore $Q_{\tau}(G)\leq q+h$ by the definition of the $\tau$-quantile.
For the lower bound, let t<qht<q-h. Then $t\leq q-r_{0}$ need not hold, but we may still compare $F(t)$ to F(qh)F(q-h) by splitting the integral:
$F(q-h)-F(t)=\int_{t}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u\geq\int_{q-h}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u+\int_{t}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u.$
Since $f\geq m_{0}$ holds on $q-r_{0},q+r_{0}]$ and $q-h\in[q-r_{0},q+r_{0}]$, we have
$F(q-h)-F(t)\geq\int_{q-h}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u+\int_{t}^{q-h}0\,\mathop{}\!\mathrm{d}u\geq 0,$
and more importantly, for any t<qht<q-h,
$F(q-h)-F(t)\geq\int_{q-h}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u+\int_{t}^{q-h}0\,\mathop{}\!\mathrm{d}u\geq m_{0}\,(q-h-t),$
where the last inequality uses that on the portion of the interval lying inside $q-r_{0},q+r_{0}]$ the integrand is at least $m_{0}$, while outside it the integrand is nonnegative. Hence,
$F(t)\leq F(q-h)-m_{0}(q-h-t)\leq(\tau-\Delta)-m_{0}(q-h-t)<\tau-\Delta.$
Consequently,
$G(t)\leq F(t)+\Delta<\tau.$
Since $G(t)<\tau$ for all t<qht<q-h, we conclude that $Q_{\tau}(G)\geq q-h$.
Combining the two bounds gives
$q-\frac{\Delta}{m_{0}}\leq Q_{\tau}(G)\leq q+\frac{\Delta}{m_{0}},$
and hence $\left|Q_{\tau}(G)-Q_{\tau}(F)\right|\leq\Delta/m_{0}$. ∎
B.3 Lemmas on Asymptotically Negligible Biases
B.3.1 Error from conditional expectations
Proposition B.4.
Define
$n_{1}\,\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}\coloneqq\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{j\in I_{1}}\sum_{p=1}^{P_{n}}\sum_{q=1}^{Q_{n}}k_{\mathcal{Z}}(Z_{i},Z_{j})\,\zeta^{(i)}_{X\mid Z,p}\,\zeta^{(i)}_{Y\mid Z,q}\,\zeta^{(j)}_{X\mid Z,p}\,\zeta^{(j)}_{Y\mid Z,q}$.
Assume $H_{0}\colon\,X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$, Assumption 2.1, 4.2, 4.3. Then
$n_{1}\left(\hat{T}_{n_{1}}-\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}\right)=o_{p}(1).$
Hence, under $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$, the regression-induced bias in $n_{1}\hat{T}_{n_{1}}$ is asymptotically negligible. Furthermore, under Assumption 4.6, this convergence in probability holds uniformly in $\tilde{\mathcal{P}}_{0}$.
Proof.
Define
$\hat{S}_{n_{1}}\coloneqq\frac{1}{\sqrt{n_{1}}}\sum_{i\in I_{1}}\hat{\Psi}_{i}^{(P_{n},Q_{n})}$
Using $\langle\hat{e}_{p},\hat{e}_{p^{\prime}}\rangle=\delta_{pp^{\prime}}$, $\langle\hat{f}_{q},\hat{f}_{q^{\prime}}\rangle=\delta_{qq^{\prime}}$ and $\langle\phi_{\mathcal{Z}}(Z_{i}),\phi_{\mathcal{Z}}(Z_{j})\rangle=k_{\mathcal{Z}}(Z_{i},Z_{j})$, where $\delta_{rs}$ denotes the Kronecker delta (i.e., $\delta_{rs}=1$ if $r=s$ and 0 otherwise), we have
$\left\|\hat{S}_{n_{1}}\right\|^{2}=\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{j\in I_{1}}\left(\sum_{p=1}^{P_{n}}\hat{\zeta}^{(i)}_{X\mid Z,p}\hat{\zeta}^{(j)}_{X\mid Z,p}\right)\left(\sum_{q=1}^{Q_{n}}\hat{\zeta}^{(i)}_{Y\mid Z,q}\hat{\zeta}^{(j)}_{Y\mid Z,q}\right)k_{\mathcal{Z}}(Z_{i},Z_{j})\,=\,n_{1}\hat{T}_{n_{1}}.$
Similarly, with $u^{(i)}_{X}\coloneqq\sum_{p=1}^{P_{n}}\zeta^{(i)}_{X\mid Z,p}\hat{e}_{p}$ and $u^{(i)}_{Y}\coloneqq\sum_{q=1}^{Q_{n}}\zeta^{(i)}_{Y\mid Z,q}\hat{f}_{q}$, define
$S_{n_{1}}\coloneqq\frac{1}{\sqrt{n_{1}}}\sum_{i\in I_{1}}u^{(i)}_{X}\otimes u^{(i)}_{Y}\otimes\phi_{\mathcal{Z}}(Z_{i}),\qquad\left\|S_{n_{1}}\right\|^{2}=n_{1}\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}.$
Hence
$n_{1}\left(\hat{T}_{n_{1}}-\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}\right)=\left\|\hat{S}_{n_{1}}\right\|^{2}-\left\|S_{n_{1}}\right\|^{2}=\left\|\hat{S}_{n_{1}}-S_{n_{1}}\right\|^{2}+2\left\langle S_{n_{1}},\,\hat{S}_{n_{1}}-S_{n_{1}}\right\rangle.$
Therefore it suffices to show
$\left\|\hat{S}_{n_{1}}-S_{n_{1}}\right\|^{2}=o_{p}(1)~~~\text{and}~~~\left|\left\langle S_{n_{1}},\hat{S}_{n_{1}}-S_{n_{1}}\right\rangle\right|=o_{p}(1).$
Write $\hat{\zeta}^{(i)}_{X\mid Z,p}=\zeta^{(i)}_{X\mid Z,p}-b^{(i)}_{X,p}$ and $\hat{\zeta}^{(i)}_{Y\mid Z,q}=\zeta^{(i)}_{Y\mid Z,q}-b^{(i)}_{Y,q}$. Then
$\hat{S}_{n_{1}}-S_{n_{1}}=B_{n_{1}}^{(1)}+B_{n_{1}}^{(2)}+B_{n_{1}}^{(3)},$
where
$B_{n_{1}}^{(1)}\coloneqq\frac{1}{\sqrt{n_{1}}}\sum_{i\in I_{1}}\left(\sum_{p=1}^{P_{n}}b^{(i)}_{X,p}\,\hat{e}_{p}\right)\otimes\left(\sum_{q=1}^{Q_{n}}\zeta^{(i)}_{Y\mid Z,q}\,\hat{f}_{q}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}),$
$B_{n_{1}}^{(2)}\coloneqq\frac{1}{\sqrt{n_{1}}}\sum_{i\in I_{1}}\left(\sum_{p=1}^{P_{n}}\zeta^{(i)}_{X\mid Z,p}\,\hat{e}_{p}\right)\otimes\left(\sum_{q=1}^{Q_{n}}b^{(i)}_{Y,q}\,\hat{f}_{q}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}),$
$B_{n_{1}}^{(3)}\coloneqq\frac{1}{\sqrt{n_{1}}}\sum_{i\in I_{1}}\left(\sum_{p=1}^{P_{n}}b^{(i)}_{X,p}\,\hat{e}_{p}\right)\otimes\left(\sum_{q=1}^{Q_{n}}b^{(i)}_{Y,q}\,\hat{f}_{q}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}).$
Then
$\left\|\hat{S}_{n_{1}}-S_{n_{1}}\right\|^{2}=\left\|B_{n_{1}}^{(1)}\right\|^{2}+\left\|B_{n_{1}}^{(2)}\right\|^{2}+\left\|B_{n_{1}}^{(3)}\right\|^{2}+2\langle B_{n_{1}}^{(1)},B_{n_{1}}^{(2)}\rangle+2\langle B_{n_{1}}^{(1)},B_{n_{1}}^{(3)}\rangle+2\langle B_{n_{1}}^{(2)},B_{n_{1}}^{(3)}\rangle.$
By $2\langle u,v\rangle\leq\|u\|^{2}+\|v\|^{2}$,
$\left\|\hat{S}_{n_{1}}-S_{n_{1}}\right\|^{2}\leq 3\sum_{i=1}^{3}\left\|B_{n_{1}}^{(i)}\right\|^{2}.$
Hence it suffices to show $\left\|B_{n_{1}}^{(i)}\right\|^{2}\xrightarrow{p}0$ for $i=1,2,3$.
First, consider $B_{n_{1}}^{(1)}$. By orthonormality and $\langle\phi_{\mathcal{Z}}(Z_{i}),\phi_{\mathcal{Z}}(Z_{j})\rangle=k_{\mathcal{Z}}(Z_{i},Z_{j})$,
$\left\|B_{n_{1}}^{(1)}\right\|^{2}=\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{j\in I_{1}}\sum_{p=1}^{P_{n}}\sum_{q=1}^{Q_{n}}k_{\mathcal{Z}}(Z_{i},Z_{j})\,b^{(i)}_{X,p}b^{(j)}_{X,p}\,\zeta^{(i)}_{Y\mid Z,q}\zeta^{(j)}_{Y\mid Z,q}$.
Split into diagonal and off-diagonal parts.
Diagonal $i=j$. Using $\left|k_{\mathcal{Z}}(Z_{i},Z_{j})\right|\leq K_{Z}$ by Lemma B.12, Lemma B.11 and Cauchy–Schwarz,
$\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{p=1}^{P_{n}}\sum_{q=1}^{Q_{n}}k_{\mathcal{Z}}(Z_{i},Z_{j})\left(b^{(i)}_{X,p}\right)^{2}\,\left(\zeta^{(i)}_{Y\mid Z,q}\right)^{2}$ $\leq\frac{K_{Z}}{n_{1}}\sum_{i\in I_{1}}\left(\sum_{p=1}^{P_{n}}\left(b^{(i)}_{X,p}\right)^{2}\right)\left(\sum_{q=1}^{Q_{n}}\left(\zeta^{(i)}_{Y\mid Z,q}\right)^{2}\right)$
$\leq\frac{4K_{Y}K_{Z}}{n_{1}}\sum_{i\in I_{1}}\left(\sum_{p=1}^{P_{n}}\left(b^{(i)}_{X,p}\right)^{2}\right)$
Assumption 4.2 yield
$\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{p=1}^{P_{n}}\left(b^{(i)}_{X,p}\right)^{2}=o_{p}(1),$
hence the diagonal contribution is $o_{p}(1)$.
Off-diagonal $i\neq j$. Let
$\mathcal{F}_{n}\coloneqq\sigma(Z_{i}\colon i\in I_{1})\vee\sigma((X_{\ell},Y_{\ell},Z_{\ell})\colon\ell\in I_{2}),$
where $\sigma(\cdot)$ denotes the $\sigma$-algebra generated by the random variables. Under $H_{0}\colon X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$, we have
$\operatorname{\mathbb{E}}\left[\zeta^{(i)}_{Y\mid Z,q}\mid Z_{i}\right]=0\quad\text{and}\quad\operatorname{\mathbb{E}}\left[\zeta^{(i)}_{Y\mid Z,q}\,\zeta^{(j)}_{Y\mid Z,q}\mid Z_{i},Z_{j}\right]=0\quad(\forall i\neq j),$
so
$\operatorname{\mathbb{E}}\left[\frac{1}{n_{1}}\sum_{\begin{subarray}{c}i,j\in I_{1}\\ j\neq i\end{subarray}}\sum_{p=1}^{P_{n}}\sum_{q=1}^{Q_{n}}k_{\mathcal{Z}}(Z_{i},Z_{j})b^{(i)}_{X,p}b^{(j)}_{X,p}\zeta^{(i)}_{Y\mid Z,q}\zeta^{(j)}_{Y\mid Z,q}\mid\mathcal{F}_{n}\right]=0.$
Therefore, by Lemma B.15, we have
$\left\|B_{n_{1}}^{(1)}\right\|^{2}=o_{p}(1).$
The argument is symmetric, exchanging the roles of $X$ and $Y$. Therefore, we can obtain
$\left\|B_{n_{1}}^{(2)}\right\|^{2}=o_{p}(1).$
By orthonormality,
$\left\|B_{n_{1}}^{(3)}\right\|^{2}$ $=\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{j\in I_{1}}k_{\mathcal{Z}}(Z_{i},Z_{j})\left(\sum_{p=1}^{P_{n}}b^{(i)}_{X,p}b^{(j)}_{X,p}\right)\left(\sum_{q=1}^{Q_{n}}b^{(i)}_{Y,q}b^{(j)}_{Y,q}\right)$
$\leq n_{1}K_{Z}\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{p=1}^{P_{n}}\left(b_{X\mid Z,p}^{(i)}\right)^{2}\right)\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\sum_{q=1}^{Q_{n}}\left(b_{Y\mid Z,q}^{(i)}\right)^{2}\right),$
so Assumption 4.3 yields $\left\|B_{n_{1}}^{(3)}\right\|^{2}=o_{p}(1)$.
By Cauchy–Schwarz,
$\left|\langle B_{n_{1}}^{(1)},B_{n_{1}}^{(2)}\rangle\right|\leq\left\|B_{n_{1}}^{(1)}\right\|\left\|B_{n_{1}}^{(2)}\right\|=o_{p}(1),$
and similarly for $\langle B_{n_{1}}^{(1)},B_{n_{1}}^{(3)}\rangle$ and $\langle B_{n_{1}}^{(2)},B_{n_{1}}^{(3)}\rangle$. Hence
$\left\|\hat{S}_{n_{1}}-S_{n_{1}}\right\|^{2}=o_{p}(1).$
Moreover, by Cauchy–Schwarz again and the assumption $\left\|S_{n_{1}}\right\|^{2}=O_{p}(1)$,
$\left|\left\langle S_{n_{1}},\,\hat{S}_{n_{1}}-S_{n_{1}}\right\rangle\right|\leq\left\|S_{n_{1}}\right\|\left\|\hat{S}_{n_{1}}-S_{n_{1}}\right\|=O_{p}(1)\cdot o_{p}(1)^{1/2}=o_{p}(1).$
Combining the displays proves
$n_{1}\left|\hat{T}_{n_{1}}-\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}\right|=o_{p}(1).$
According to Lemmas S5 and S6 in Lundborg et al. (2024), this convergence in probability can be replaced by uniform convergence under Assumption 4.6. ∎
B.3.2 Error from Eigenfunction Truncation
Lemma B.5.
Define
$T_{n_{1}}^{(P_{n},Q_{n})}\coloneqq\left\|\left(P_{X,P_{n}}\otimes P_{Y,Q_{n}}\otimes I\right)\Sigma_{n_{1}}\right\|^{2},$
where $I$ denotes the identity operator on $\mathcal{H}_{\mathcal{Z}}$. Then, under $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$ and Assumption 4.1, we have
$n_{1}\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|=o_{p}(1).$
Furthermore, under Assumption 4.5, this convergence in probability holds uniformly in $\tilde{\mathcal{P}}_{0}$.
Proof.
Define
$U\coloneqq\left(\hat{P}_{X,P_{n}}\otimes\hat{P}_{Y,Q_{n}}\otimes I\right)\Sigma_{n_{1}},\qquad V\coloneqq\left(P_{X,P_{n}}\otimes P_{Y,Q_{n}}\otimes I\right)\Sigma_{n_{1}}.$
By the polarization identity,
$n_{1}\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|$ $=n_{1}\left|\left\|U\right\|^{2}-\left\|V\right\|^{2}\right|$
$=n_{1}\left|\left\langle U+V,\ U-V\right\rangle\right|$
$\leq n_{1}\left(\left\|U\right\|+\left\|V\right\|\right)\left\|U-V\right\|.$
Since orthogonal projectors have operator norm at most $1$,
$\left\|U\right\|\leq\left\|\Sigma_{n_{1}}\right\|,\qquad\left\|V\right\|\leq\left\|\Sigma_{n_{1}}\right\|,$
hence
$n_{1}\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|\leq 2\,n_{1}\,\left\|\Sigma_{n_{1}}\right\|\,\left\|U-V\right\|.$ (A.19)
Define
$\Delta_{X,P_{n}}\coloneqq\hat{P}_{X,P_{n}}-P_{X,P_{n}},\qquad\Delta_{Y,Q_{n}}\coloneqq\hat{P}_{Y,Q_{n}}-P_{Y,Q_{n}}.$
By bilinearity of the tensor product,
$\hat{P}_{X,P_{n}}\otimes\hat{P}_{Y,Q_{n}}$ $=\left(P_{X,P_{n}}+\Delta_{X,P_{n}}\right)\otimes\left(P_{Y,Q_{n}}+\Delta_{Y,Q_{n}}\right)$
$=P_{X,P_{n}}\otimes P_{Y,Q_{n}}+\Delta_{X,P_{n}}\otimes P_{Y,Q_{n}}+P_{X,P_{n}}\otimes\Delta_{Y,Q_{n}}+\Delta_{X,P_{n}}\otimes\Delta_{Y,Q_{n}}.$
Therefore
$U-V=\left[\left(\Delta_{X,P_{n}}\otimes P_{Y,Q_{n}}\otimes I\right)+\left(P_{X,P_{n}}\otimes\Delta_{Y,Q_{n}}\otimes I\right)+\left(\Delta_{X,P_{n}}\otimes\Delta_{Y,Q_{n}}\otimes I\right)\right]\Sigma_{n_{1}}.$
Using submultiplicativity and $\left\|A\otimes B\otimes C\right\|_{\mathrm{op}}=\left\|A\right\|_{\mathrm{op}}\left\|B\right\|_{\mathrm{op}}\left\|C\right\|_{\mathrm{op}}$ together with $\left\|P_{X,P_{n}}\right\|_{\mathrm{op}}=\left\|P_{Y,Q_{n}}\right\|_{\mathrm{op}}=1$, we obtain
$\left\|U-V\right\|\leq\left(\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}+\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}+\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}\right)\left\|\Sigma_{n_{1}}\right\|.$
Insert this bound into (A.19) to conclude
$\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|\leq 2n_{1}\left\|\Sigma_{n_{1}}\right\|^{2}\left\{\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}+\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}+\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}\right\}.$ (A.20)
We detail the $X$-part; the $Y$-part is identical with $(C_{YY},\mu_{j},\gamma^{Y}_{Q_{n}})$ in place of $(C_{XX},\lambda_{j},\gamma^{X}_{P_{n}})$.
Let
$A\coloneqq C_{XX},\qquad\hat{A}\coloneqq\hat{C}_{XX}=A+H,\qquad H\coloneqq\hat{C}_{XX}-C_{XX}.$
Let $\lambda_{1}>\lambda_{2}>\cdots$ be the (simple) positive eigenvalues of $A$ and fix $P_{n}$ such that
$\gamma^{X}_{P_{n}}\coloneqq\lambda_{P_{n}}-\lambda_{P_{n}+1}>0.$
Define the midpoint threshold
$r_{P_{n}}\coloneqq\frac{\lambda_{P_{n}}+\lambda_{P_{n}+1}}{2},\qquad\sigma_{X}\coloneqq(r_{P_{n}},\infty),\qquad\Sigma_{X}\coloneqq(-\infty,r_{P_{n}}].$
Note that
$\mathrm{dist}\left(r_{P_{n}},\mathrm{spec}(A)\right)=\inf_{\lambda\in\mathrm{spec}(A)}\left|r_{P_{n}}-\lambda\right|=\min\left\{\lambda_{P_{n}}-r_{P_{n}},\,r_{P_{n}}-\lambda_{P_{n}+1}\right\}=\frac{\gamma^{X}_{P_{n}}}{2}.$
Let $E_{A}(\cdot)$ denote the spectral measure of the bounded self-adjoint operator $A$. Then the population projector onto the top-$P_{n}$ eigenspace can be written as the spectral projection
$P_{X,P_{n}}=E_{A}(\sigma_{X}).$
Define the “good” event
$\Omega^{X}_{n}\coloneqq\left\{\left\|H\right\|_{\mathrm{op}}<\frac{\gamma^{X}_{P_{n}}}{4}\right\}.$
On $\Omega^{X}_{n}$, Weyl’s eigenvalue inequality gives, for the ordered eigenvalues $\hat{\lambda}_{j}$ of $\hat{A}$,
$\hat{\lambda}_{P_{n}}\geq\lambda_{P_{n}}-\left\|H\right\|_{\mathrm{op}}>\lambda_{P_{n}}-\frac{\gamma^{X}_{P_{n}}}{4}=r_{P_{n}}+\frac{\gamma^{X}_{P_{n}}}{4}>r_{P_{n}},$
and
$\hat{\lambda}_{P_{n}+1}\leq\lambda_{P_{n}+1}+\left\|H\right\|_{\mathrm{op}}<\lambda_{P_{n}+1}+\frac{\gamma^{X}_{P_{n}}}{4}=r_{P_{n}}-\frac{\gamma^{X}_{P_{n}}}{4}<r_{P_{n}}.$
Hence, on $\Omega^{X}_{n}$, $\hat{A}$ has exactly $P_{n}$ eigenvalues in $(r_{P_{n}},\infty)$, so the empirical projector onto the top-$P_{n}$ eigenspace coincides with the spectral projection:
$\hat{P}_{X,P_{n}}=E_{\hat{A}}(\sigma_{X})\qquad\text{on }\Omega^{X}_{n}.$
Let $\Theta_{X}$ be the (operator) angle between the subspaces $\mathrm{Ran}(P_{X,P_{n}})$ and $\mathrm{Ran}(\hat{P}_{X,P_{n}})$. By the symmetric $\sin\Theta$ theorem of Davis and Kahan (1970),
$\mathrm{dist}\left(r_{P_{n}},\mathrm{spec}(A)\right)\,\left\|\sin\Theta_{X}\right\|_{\mathrm{op}}\leq\left\|H\right\|_{\mathrm{op}}.$
Moreover, for orthogonal projections one has the identity
$\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\mathrm{op}}=\left\|\sin\Theta_{X}\right\|_{\mathrm{op}}$
Therefore, on $\Omega^{X}_{n}$,
$\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}=\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\mathrm{op}}\leq\frac{1}{\mathrm{dist}\left(r_{P_{n}},\mathrm{spec}(A)\right)}\left\|H\right\|_{\mathrm{op}}=\frac{2}{\gamma^{X}_{P_{n}}}\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{op}}.$ (A.21)
The same argument yields
$\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}\leq\frac{2}{\gamma^{Y}_{Q_{n}}}\left\|\hat{C}_{YY}-C_{YY}\right\|_{\mathrm{op}}.$ (A.22)
Since $\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\mathrm{op}}\leq 2$ always, for any $\varepsilon>0$,
$\mathbb{P}\left(\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}>\varepsilon\right)\leq\mathbb{P}\left((\Omega^{X}_{n})^{c}\right)+\mathbb{P}\left(\frac{2}{\gamma^{X}_{P_{n}}}\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{op}}>\varepsilon\right).$
By Assumption 4.1, $\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{op}}/\gamma^{X}_{P_{n}}=o_{p}(1)$, hence both probabilities on the right-hand side converge to 0 and
$\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}=o_{p}(1).$
The same reasoning with (A.22) yields $\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}=o_{p}(1)$.
Next, we present a uniform convergence result. Fix $\varepsilon>0$. Using (A.21) and $\left\|\cdot\right\|_{\mathrm{op}}\leq\left\|\cdot\right\|_{\mathrm{HS}}$,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}>\varepsilon\right)\leq\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}>\frac{\varepsilon\,\gamma^{X}_{P_{n}}}{2}\right).$
By Markov’s inequality,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}>\frac{\varepsilon\,\gamma^{X}_{P_{n}}}{2}\right)\leq\frac{2}{\varepsilon\,\gamma^{X}_{P_{n}}}\sup_{P\in\tilde{\mathcal{P}}_{0}}\operatorname{\mathbb{E}}_{P}\left[\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}\right].$
Under bounded kernels, the Hilbert–Schmidt norm of each centered rank-one summand in $\hat{C}_{XX}-C_{XX}$ is uniformly bounded, hence there exists a constant $C<\infty$ independent of $P\in\tilde{\mathcal{P}}_{0}$ such that
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\operatorname{\mathbb{E}}_{P}\left[\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}\right]\leq\frac{C}{\sqrt{n_{1}}}.$
Therefore,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}>\varepsilon\right)\leq\frac{2C}{\varepsilon\,\gamma^{X}_{P_{n}}\sqrt{n_{1}}}\longrightarrow 0$
by Assumption 4.5. The same argument applies to $\Delta_{Y,Q_{n}}$. Combining this with (A.20) and $\sqrt{n_{1}}\left\|\Sigma_{n_{1}}\right\|=O_{\tilde{\mathcal{P}}_{0}}(1)$ (which follows from boundedness of the kernels and the summands defining $\Sigma_{n_{1}}$), we obtain the claimed uniform convergence. ∎
Lemma B.6.
Let $\Pi_{X}$ and $\Pi_{Y}$ be the orthogonal projectors onto $\overline{\operatorname{range}\,C_{XX}}$ and $\overline{\operatorname{range}\,C_{YY}}$, respectively. Define
$\bar{T}_{n_{1}}\coloneqq\|\Sigma_{n_{1}}\|^{2}.$
If $P_{n},Q_{n}\to\infty$, then under $X\mathrel{\perp\mkern-10.0mu\perp}Y\mid Z$, we have
$n_{1}\left|T_{n_{1}}^{(P_{n},Q_{n})}-\bar{T}_{n_{1}}\right|\overset{p}{\longrightarrow}0.$
Furthermore, under Assumptions 2.1 and 4.7, this convergence in probability holds uniformly in $\tilde{\mathcal{P}}_{0}$.
Proof.
For any $g\in\ker C_{XX}$,
$0=\langle C_{XX}g,g\rangle=\operatorname{\mathbb{E}}\left[\langle k_{\mathcal{X}}(X,\cdot),g\rangle^{2}\right]=\operatorname{\mathbb{E}}\left[g(X)^{2}\right]$
so $g(X)=0$ almost surely. Hence
$\langle g,\epsilon_{X\mid Z}^{(i)}\rangle=\langle g,k_{\mathcal{X}}(X_{i},\cdot)\rangle-\langle g,\mu_{X\mid Z}(Z_{i})\rangle=g(X_{i})-\operatorname{\mathbb{E}}\left[g(X_{i})\mid Z_{i}\right]=0,$
and therefore $\epsilon_{X\mid Z}^{(i)}\in(\ker C_{XX})^{\perp}=\overline{\operatorname{range}\,C_{XX}}$ almost surely. The same argument gives $\epsilon_{Y\mid Z}^{(i)}\in\overline{\operatorname{range}\,C_{YY}}$ almost surely. Consequently,
$\Sigma_{n_{1}}=\left(\Pi_{X}\otimes\Pi_{Y}\otimes I\right)\Sigma_{n_{1}}.$ (A.23)
Define
$\Delta_{n_{1}}^{(P_{n},Q_{n})}\coloneqq\Sigma_{n_{1}}-\left(P_{X,P_{n}}\otimes P_{Y,Q_{n}}\otimes I\right)\Sigma_{n_{1}}.$
Since $\Sigma_{n_{1}}=(\Pi_{X}\otimes\Pi_{Y}\otimes I)\Sigma_{n_{1}}$, we have the identity
$\Delta_{n_{1}}^{(P_{n},Q_{n})}$ $=\Bigl((\Pi_{X}-P_{X,P_{n}})\otimes\Pi_{Y}\otimes I+\Pi_{X}\otimes(\Pi_{Y}-P_{Y,Q_{n}})\otimes I$
$\quad+(\Pi_{X}-P_{X,P_{n}})\otimes(\Pi_{Y}-P_{Y,Q_{n}})\otimes I\Bigr)\Sigma_{n_{1}}.$
Equivalently,
$\Delta_{n_{1}}^{(P_{n},Q_{n})}=\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\{V_{i}^{(1)}+V_{i}^{(2)}+V_{i}^{(3)}\right\},$
where
$V_{i}^{(1)}\coloneqq\left((\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}^{(i)}\right)\otimes\left(\Pi_{Y}\epsilon_{Y\mid Z}^{(i)}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}),$
$V_{i}^{(2)}\coloneqq\left(\Pi_{X}\epsilon_{X\mid Z}^{(i)}\right)\otimes\left((\Pi_{Y}-P_{Y,Q_{n}})\epsilon_{Y\mid Z}^{(i)}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}),$
$V_{i}^{(3)}\coloneqq\left((\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}^{(i)}\right)\otimes\left((\Pi_{Y}-P_{Y,Q_{n}})\epsilon_{Y\mid Z}^{(i)}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}).$
By $\left(a+b+c\right)^{2}\leq 3\left(a^{2}+b^{2}+c^{2}\right)$,
$\left\|\Delta_{n_{1}}^{(P_{n},Q_{n})}\right\|^{2}\leq\frac{3}{n_{1}^{2}}\sum_{r=1}^{3}\left\|\sum_{i\in I_{1}}V_{i}^{(r)}\right\|^{2}.$
Under $H_{0}$ we have $\operatorname{\mathbb{E}}\left[\epsilon_{X\mid Z}^{(i)}\mid Z_{i}\right]=0$ and $\operatorname{\mathbb{E}}\left[\epsilon_{Y\mid Z}^{(i)}\mid Z_{i}\right]=0$, hence
$\operatorname{\mathbb{E}}\left[V_{i}^{(r)}\right]=0\qquad(r=1,2,3).$
By independence across $i$,
$\operatorname{\mathbb{E}}\left\|\sum_{i\in I_{1}}V_{i}^{(r)}\right\|^{2}=n_{1}\,\operatorname{\mathbb{E}}\left\|V_{1}^{(r)}\right\|^{2}\qquad(r=1,2,3).$
Therefore,
$\operatorname{\mathbb{E}}\left[n_{1}\left\|\Delta_{n_{1}}^{(P_{n},Q_{n})}\right\|^{2}\right]\leq 3\sum_{r=1}^{3}\operatorname{\mathbb{E}}\left\|V_{1}^{(r)}\right\|^{2}.$
Using $\|u\otimes v\otimes w\|^{2}=\|u\|^{2}\|v\|^{2}\|w\|^{2}$ and $\|\phi_{\mathcal{Z}}(Z_{1})\|^{2}=k_{\mathcal{Z}}(Z_{1},Z_{1})\leq K_{Z}$, and $\left\|\Pi_{Y}\epsilon_{Y\mid Z}^{(1)}\right\|\leq\left\|\epsilon_{Y\mid Z}^{(1)}\right\|$,
$\operatorname{\mathbb{E}}\left\|V_{1}^{(1)}\right\|^{2}\leq K_{Z}\,\operatorname{\mathbb{E}}\left\|(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}^{(1)}\right\|^{2}\left\|\epsilon_{Y\mid Z}^{(1)}\right\|^{2}.$
Diagonal kernel bounds and conditional Jensen yield almost surely
$\left\|\epsilon_{Y\mid Z}^{(1)}\right\|^{2}\leq 2\left\|k_{\mathcal{Y}}(Y_{1},\cdot)\right\|^{2}+2\left\|\mu_{Y\mid Z}(Z_{1})\right\|^{2}\leq 2K_{Y}+2\operatorname{\mathbb{E}}\left[k_{\mathcal{Y}}(Y_{1},Y_{1})\mid Z_{1}\right]\leq 4K_{Y}.$
Hence
$\operatorname{\mathbb{E}}\left\|V_{1}^{(1)}\right\|^{2}\leq 4K_{Y}K_{Z}\,\operatorname{\mathbb{E}}\left\|(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}^{(1)}\right\|^{2}.$
Similarly,
$\operatorname{\mathbb{E}}\left\|V_{1}^{(2)}\right\|^{2}\leq 4K_{X}K_{Z}\,\operatorname{\mathbb{E}}\left\|(\Pi_{Y}-P_{Y,Q_{n}})\epsilon_{Y\mid Z}^{(1)}\right\|^{2},$
and
$\operatorname{\mathbb{E}}\left\|V_{1}^{(3)}\right\|^{2}\leq 4K_{Z}\,\operatorname{\mathbb{E}}\left\|(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}^{(1)}\right\|^{2}\left\|(\Pi_{Y}-P_{Y,Q_{n}})\epsilon_{Y\mid Z}^{(1)}\right\|^{2}.$
By (A.23), $\epsilon_{X\mid Z}^{(1)}\in\operatorname{range}\,\Pi_{X}$ and $\epsilon_{Y\mid Z}^{(1)}\in\operatorname{range}\,\Pi_{Y}$ almost surely. By Lemma B.7,
$\left\|(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}^{(1)}\right\|^{2}\downarrow 0,\qquad\left\|(\Pi_{Y}-P_{Y,Q_{n}})\epsilon_{Y\mid Z}^{(1)}\right\|^{2}\downarrow 0$
almost surely. The bounds $\left\|\epsilon_{X\mid Z}^{(1)}\right\|^{2}\leq 4K_{X}$ and $\left\|\epsilon_{Y\mid Z}^{(1)}\right\|^{2}\leq 4K_{Y}$ provide integrable domination, so by dominated convergence
$\operatorname{\mathbb{E}}\left\|(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}^{(1)}\right\|^{2}\to 0,\qquad\operatorname{\mathbb{E}}\left\|(\Pi_{Y}-P_{Y,Q_{n}})\epsilon_{Y\mid Z}^{(1)}\right\|^{2}\to 0.$
Therefore $\sum_{r=1}^{3}\operatorname{\mathbb{E}}\left\|V_{1}^{(r)}\right\|^{2}\to 0$, and hence
$\operatorname{\mathbb{E}}\left[n_{1}\left\|\Delta_{n_{1}}^{(P_{n},Q_{n})}\right\|^{2}\right]\to 0.$
By Markov’s inequality,
$n_{1}\left\|\Delta_{n_{1}}^{(P_{n},Q_{n})}\right\|^{2}\overset{p}{\longrightarrow}0,\qquad\sqrt{n_{1}}\left\|\Delta_{n_{1}}^{(P_{n},Q_{n})}\right\|\overset{p}{\longrightarrow}0.$
Using
$\bar{T}_{n_{1}}^{(P_{n},Q_{n})}-\bar{T}_{n_{1}}$ $=\left\|\left(P_{X,P_{n}}\otimes P_{Y,Q_{n}}\otimes I\right)\Sigma_{n_{1}}\right\|^{2}-\|\Sigma_{n_{1}}\|^{2}$
$=\left\|\Sigma_{n_{1}}-\Delta_{n_{1}}^{(P_{n},Q_{n})}\right\|^{2}-\|\Sigma_{n_{1}}\|^{2},$
the polarization inequality gives
$\left|\bar{T}_{n_{1}}^{(P_{n},Q_{n})}-\bar{T}_{n_{1}}\right|\leq 2\|\Sigma_{n_{1}}\|\left\|\Delta_{n_{1}}^{(P_{n},Q_{n})}\right\|+\left\|\Delta_{n_{1}}^{(P_{n},Q_{n})}\right\|^{2}.$
Multiplying by $n_{1}$ and using $n_{1}\|\Sigma_{n_{1}}\|^{2}=O_{p}(1)$ together with the preceding convergence yields
$n_{1}\left|\bar{T}_{n_{1}}^{(P_{n},Q_{n})}-\bar{T}_{n_{1}}\right|\overset{p}{\longrightarrow}0.$
Furthermore, under Assumptions 4.2 and 4.7, this convergence in probability holds uniformly in $\tilde{\mathcal{P}}_{0}$ by Lemma B.8 and Lemma S5 in Lundborg et al. (2024). ∎
Lemma B.7.
Assume Assumption 2.1. Let $\Pi_{X}$ be the orthogonal projector onto $\overline{\operatorname{range}\,C_{XX}}$. Then $P_{X,P_{n}}$ converges to $\Pi_{X}$ in the strong operator topology as $P_{n}\to\infty$; equivalently, for every $h\in\mathcal{H}_{\mathcal{X}}$,
$\left\|(\Pi_{X}-P_{X,P_{n}})h\right\|_{\mathcal{H}_{\mathcal{X}}}\longrightarrow 0\quad\text{as }P_{n}\to\infty.$
The analogous statement holds for $C_{YY}$, $\Pi_{Y}$, and $P_{Y,Q_{n}}$.
Proof.
By Assumption 2.1, $\operatorname{\mathbb{E}}\left[k_{\mathcal{X}}(X,X)\right]<\infty$, so $C_{XX}$ is positive, self-adjoint, and trace-class, hence compact. Thus there exist eigenpairs $(\lambda_{j},e_{j})$ with $\lambda_{j}>0$ such that $\{e_{j}\}_{j\geq 1}$ is an orthonormal basis of $\overline{\operatorname{range}\,C_{XX}}=(\ker C_{XX})^{\perp}$. For any $h\in\mathcal{H}_{\mathcal{X}}$, decompose $h=h_{+}+h_{0}$ with $h_{+}\in\overline{\operatorname{range}\,C_{XX}}$ and $h_{0}\in\ker C_{XX}$. Then
$(\Pi_{X}-P_{X,P_{n}})h=h_{+}-\sum_{j=1}^{P_{n}}\langle h_{+},e_{j}\rangle e_{j},$
whence by Parseval’s identity,
$\left\|(\Pi_{X}-P_{X,P_{n}})h\right\|^{2}=\sum_{j>P_{n}}\left|\langle h_{+},e_{j}\rangle\right|^{2}\to 0\quad(P_{n}\to\infty).$
Therefore $P_{X,P_{n}}$ converges to $\Pi_{X}$ in the strong operator topology. The argument for $P_{Y,Q_{n}}\to\Pi_{Y}$ is identical. ∎
Lemma B.8.
Under Assumptions 2.1 and 4.7, we have
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \operatorname{\mathbb{E}}_{P}\left[\left\|\,\left(\Pi_{X}-P_{X,P_{n}}\right)\epsilon_{X\mid Z}\right\|^{2}\right]\xrightarrow[P_{n}\to\infty]{}0.$
Proof.
By $\epsilon_{X\mid Z}=k_{\mathcal{X}}(X,\cdot)-\mu_{X\mid Z}(Z)$ and $\|a-b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$,
$\operatorname{\mathbb{E}}\left[\left\|(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}\right\|^{2}\right]$
$\quad\leq 2\,\operatorname{\mathbb{E}}\left[\left\|(\Pi_{X}-P_{X,P_{n}})k_{\mathcal{X}}(X,\cdot)\right\|^{2}\right]+2\,\operatorname{\mathbb{E}}\left[\left\|(\Pi_{X}-P_{X,P_{n}})\mu_{X\mid Z}(Z)\right\|^{2}\right].$
By conditional Jensen,
$\operatorname{\mathbb{E}}\left[\left\|(\Pi_{X}-P_{X,P_{n}})\mu_{X\mid Z}(Z)\right\|^{2}\right]\leq\operatorname{\mathbb{E}}\left[\left\|(\Pi_{X}-P_{X,P_{n}})k_{\mathcal{X}}(X,\cdot)\right\|^{2}\right].$
Hence
$\operatorname{\mathbb{E}}\left[\left\|(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}\right\|^{2}\right]\leq 4\,\operatorname{\mathbb{E}}\left[\left\|(\Pi_{X}-P_{X,P_{n}})k_{\mathcal{X}}(X,\cdot)\right\|^{2}\right].$
Since $k_{\mathcal{X}}(X,\cdot)\in\overline{\operatorname{range}\,C_{XX}}$ a.s., Parseval gives
$\operatorname{\mathbb{E}}\left\|(\Pi_{X}-P_{X,P_{n}})k_{\mathcal{X}}(X,\cdot)\right\|^{2}=\operatorname{\mathbb{E}}\left[\sum_{j>m}\langle k_{\mathcal{X}}(X,\cdot),e_{j}\rangle^{2}\right]=\sum_{j>m}\lambda_{j}.$
Let $\{\tilde{e}_{p}\}_{p\geq 1}$ be as in Assumption 4.7. By the reproducing property and Ky Fan’s variational principle (Fan, 1951),
$\sum_{j>m}\lambda_{j}\leq\sum_{p>m}\operatorname{\mathbb{E}}\left[\tilde{e}_{p}(X)^{2}\right].$
Taking $\sup_{P\in\tilde{\mathcal{P}}_{0}}$ and using Assumption 4.7 yields
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\operatorname{\mathbb{E}}_{P}\left[\left\|(\Pi_{X}-P_{X,P_{n}})k_{\mathcal{X}}(X,\cdot)\right\|^{2}\right]\xrightarrow[m\to\infty]{}0,$
and the asserted uniform convergence for $(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}$ follows from the factor $4$ bound above. ∎
B.4 Auxiliary Lemmas for Appendix B
Lemma B.9.
Write $V=(X,Y,Z)$ with law $P_{V}=P_{XYZ}$, and define
$h(v,v^{\prime})\coloneqq\left\langle\Psi(v),\Psi(v^{\prime})\right\rangle,\qquad\left(T_{h}f\right)(v)\coloneqq\int h(v,v^{\prime})f(v^{\prime})\,P_{V}(\mathop{}\!\mathrm{d}v^{\prime}),$
where $\Psi(v)\coloneqq(k_{\mathcal{X}}(x,\cdot)-\mu_{X\mid Z}(z))\otimes(k_{\mathcal{Y}}(y,\cdot)-\mu_{Y\mid Z}(z))\otimes k_{\mathcal{Z}}(z,\cdot)$ for $v=(x,y,z)$. Let $C\coloneqq\operatorname{\mathbb{E}}\left[\Psi\otimes\Psi\right]=\colon\mathcal{H}_{XYZ}\to\mathcal{H}_{XYZ}$. Then under Assumption 2.1, $T_{h}$ and $C$ are self-adjoint, compact, and they have exactly the same nonzero eigenvalues (with the same multiplicities). In particular, if $Ce_{\ell}=\xi_{\ell}e_{\ell}$ with $\xi_{\ell}>0$, then $\xi_{\ell}$ is also a nonzero eigenvalue of $T_{h}$.
Proof.
Define $J\colon\mathcal{H}_{XYZ}\to L^{2}(P_{V})$ by $(Jx)(v)\coloneqq\langle x,\Psi(v)\rangle$. Then $T_{h}=JJ^{\ast}$ and $C=J^{\ast}J$ by direct computation:
$\left(T_{h}f\right)(v)=\left\langle\Psi(v),J^{\ast}f\right\rangle,\qquad J^{\ast}Jx=\operatorname{\mathbb{E}}\left[\left\langle x,\Psi\right\rangle\,\Psi\right].$
Hence both $T_{h}$ and $C$ are self-adjoint. Since $\operatorname{\mathbb{E}}[\|\Psi\|^{2}]<\infty$, $J$ is Hilbert–Schmidt, so $J$, $^{\ast}, JJ$$^{\ast}, JJ$$^{\ast}J$ are compact (Conway, 2007, VI.3, Proposition. 3.5).
Fix $\lambda>0$. If $Ce=\lambda e$, then
$T_{h}\left(Je\right)=JJ^{\ast}Je=J\left(Ce\right)=\lambda\,Je,$
and $Je\neq 0$ because otherwise $Ce=J^{\ast}Je=0$. Thus $\lambda$ is a nonzero eigenvalue of $T_{h}$. Conversely, if $T_{h}f=\lambda f$, then
$C\left(J^{\ast}f\right)=J^{\ast}JJ^{\ast}f=J^{\ast}\left(T_{h}f\right)=\lambda\,J^{\ast}f,$
and $J^{\ast}f\neq 0$, so $\lambda$ is a nonzero eigenvalue of $C$. Moreover,
$J\colon\ker\left(C-\lambda I\right)\longrightarrow\ker(T_{h}-\lambda I),\qquad J^{\ast}\colon\ker(T_{h}-\lambda I)\longrightarrow\ker\left(C-\lambda I\right)$
are inverses up to the factor $\lambda$ on these eigenspaces, since $J^{\ast}J=\lambda I$ and $JJ^{\ast}=\lambda I$ hold after restriction. Therefore the eigenspaces have the same finite dimension, so multiplicities agree. Hence $T_{h}$ and $C$ have exactly the same nonzero eigenvalues. In particular, if $Ce_{\ell}=\xi_{\ell}e_{\ell}$ with $\xi_{\ell}>0$, then $\xi_{\ell}$ is also a nonzero eigenvalue of $T_{h}$. ∎
Lemma B.10.
Fix an index $i$ and let $\hat{e}_{p}\in\mathcal{H}_{\mathcal{X}}$ be constructed from an auxiliary sample that is independent of $\left(X_{i},Z_{i}\right)$. Then
$\zeta^{(i)}_{X\mid Z,p}\coloneqq\hat{e}_{p}(X_{i})-\operatorname{\mathbb{E}}\left[\hat{e}_{p}(X_{i})\mid Z_{i},\hat{e}_{p}\right]=\left\langle\hat{e}_{p},\ k_{\mathcal{X}}(X_{i},\cdot)-\mu_{X\mid Z}(Z_{i})\right\rangle.$
Proof.
By the reproducing property,
$\hat{e}_{p}(X_{i})=\left\langle\hat{e}_{p},\ k_{\mathcal{X}}(X_{i},\cdot)\right\rangle.$
Since $\hat{e}_{p}$ is $\sigma\left(Z_{i},\hat{e}_{p}\right)$-measurable and $k_{\mathcal{X}}(X_{i},\cdot)\in L^{2}\left(\Omega;\mathcal{H}_{\mathcal{X}}\right)$ under $\sup_{x}k_{\mathcal{X}}(x,x)<\infty$, the conditional inner-product/expectation interchange holds:
$\operatorname{\mathbb{E}}\left[\hat{e}_{p}(X_{i})\mid Z_{i},\hat{e}_{p}\right]=\operatorname{\mathbb{E}}\left[\left\langle\hat{e}_{p},\ k_{\mathcal{X}}(X_{i},\cdot)\right\rangle\mid Z_{i},\hat{e}_{p}\right]=\left\langle\hat{e}_{p},\ \operatorname{\mathbb{E}}\left[k_{\mathcal{X}}(X_{i},\cdot)\mid Z_{i},\hat{e}_{p}\right]\right\rangle.$
By independence of $\hat{e}_{p}$ and $\left(X_{i},Z_{i}\right)$,
$\operatorname{\mathbb{E}}\left[k_{\mathcal{X}}(X_{i},\cdot)\mid Z_{i},\hat{e}_{p}\right]=\operatorname{\mathbb{E}}\left[k_{\mathcal{X}}(X_{i},\cdot)\mid Z_{i}\right]=\mu_{X\mid Z}(Z_{i}).$
Subtracting the two displays yields the claim. ∎
Lemma B.11.
Assume Assumption 2.1. Then, almost surely,
$\sup_{i\in I_{1}}\ \sum_{p=1}^{P_{n}}\left(\zeta^{(i)}_{X\mid Z,p}\right)^{2}\leq 4K_{X},\qquad\sup_{i\in I_{1}}\ \sum_{q=1}^{Q_{n}}\left(\zeta^{(i)}_{Y\mid Z,q}\right)^{2}\leq 4K_{Y}.$
Proof.
By Bessel’s inequality (Conway, 2007, Corollary 4.10) and Lemma B.10, we have
$\sum_{p=1}^{P_{n}}\left(\zeta^{(i)}_{X\mid Z,p}\right)^{2}=\sum_{p=1}^{P_{n}}\left\langle\hat{e}_{p},\ \epsilon_{X\mid Z}^{(i)}\right\rangle^{2}\leq\left\|\epsilon_{X\mid Z}^{(i)}\right\|^{2}.$
We first obtain a bound for $\left\|\epsilon_{X\mid Z}^{(i)}\right\|^{2}$. By the Hilbert space inequality $\|a-b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$,
$\left\|\epsilon_{X\mid Z}^{(i)}\right\|^{2}\leq 2\left\|k_{\mathcal{X}}(X_{i},\cdot)\right\|^{2}+2\left\|\mu_{X\mid Z}(Z_{i})\right\|^{2}\leq 4K_{X}\quad\text{a.s.}$
and therefore
$\sum_{p=1}^{P_{n}}\left(\zeta^{(i)}_{X\mid Z,p}\right)^{2}\leq 4K_{X}\quad\text{a.s.}$
Taking the supremum over the finite set $I_{1}$ preserves the bound. The proof for $Y$ is identical with $(\mathcal{X},K_{X})$ replaced by $(\mathcal{Y},K_{Y})$. ∎
Lemma B.12.
Assume the kernel $k_{\mathcal{Z}}$ is bounded:
$\sup_{z\in\mathcal{Z}}k_{\mathcal{Z}}(z,z)\leq K_{Z}<\infty.$
Then
$\sup_{z,z^{\prime}\in\mathcal{Z}}\left|k_{\mathcal{Z}}(z,z^{\prime})\right|\leq K_{Z}.$
Proof.
By the reproducing property, for all $z,z^{\prime}\in\mathcal{Z}$,
$k_{\mathcal{Z}}(z,z^{\prime})=\left\langle k_{\mathcal{Z}}(z,\cdot),\,k_{\mathcal{Z}}(z^{\prime},\cdot)\right\rangle.$
Hence, by Cauchy–Schwarz in $\mathcal{H}_{\mathcal{Z}}$ and the identity $\left\|k_{\mathcal{Z}}(z,\cdot)\right\|^{2}=k_{\mathcal{Z}}(z,z)$,
$\left|k_{\mathcal{Z}}(z,z^{\prime})\right|\leq\left\|k_{\mathcal{Z}}(z,\cdot)\right\|\left\|k_{\mathcal{Z}}(z^{\prime},\cdot)\right\|=(k_{\mathcal{Z}}(z,z)\,k_{\mathcal{Z}}(z^{\prime},z^{\prime}))^{1/2}\leq K_{Z}.$
Taking the supremum over $z,z^{\prime}\in\mathcal{Z}$ yields the claim. ∎
B.5 Lemmas for Uniform Convergence in Probability
For $P,Q\in\mathbb{N}$ define
$\Psi_{i}^{(P,Q)}\coloneqq\left(P_{X,P}\epsilon_{X\mid Z}^{(i)}\right)\otimes\left(P_{Y,Q}\epsilon_{Y\mid Z}^{(i)}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}),\quad\Psi_{i}^{\sharp,(P,Q)}\coloneqq\left(\hat{P}_{X,P}\epsilon_{X\mid Z}^{(i)}\right)\otimes\left(\hat{P}_{Y,Q}\epsilon_{Y\mid Z}^{(i)}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}),$
By the same argument in the proof of Lemma B.6, the residual functions lie in the ranges of the population projectors:
$\epsilon_{X\mid Z}^{(i)}\in\operatorname{ran}\left(\Pi_{X}\right),\quad\epsilon_{Y\mid Z}^{(i)}\in\operatorname{ran}\left(\Pi_{Y}\right)\ \ \Longrightarrow\ \ \Pi_{X}\epsilon_{X\mid Z}^{(i)}=\epsilon_{X\mid Z}^{(i)},\quad\Pi_{Y}\epsilon_{Y\mid Z}^{(i)}=\epsilon_{Y\mid Z}^{(i)}.$
Hence
$\Psi_{i}^{(\infty)}=\left(\Pi_{X}\epsilon_{X\mid Z}^{(i)}\right)\otimes\left(\Pi_{Y}\epsilon_{Y\mid Z}^{(i)}\right)\otimes\phi_{\mathcal{Z}}(Z_{i})=\epsilon_{X\mid Z}^{(i)}\otimes\epsilon_{Y\mid Z}^{(i)}\otimes\phi_{\mathcal{Z}}(Z_{i})\ \eqqcolon\ \Psi_{i}.$
Define the empirical covariance operators
$C_{n}^{(P,Q)}\coloneqq\frac{1}{n_{1}}\sum_{i\in I_{1}}\Psi_{i}^{(P,Q)}\otimes\Psi_{i}^{(P,Q)},\qquad C_{n}^{\sharp,(P,Q)}\coloneqq\frac{1}{n_{1}}\sum_{i\in I_{1}}\Psi_{i}^{\sharp,(P,Q)}\otimes\Psi_{i}^{\sharp,(P,Q)},$
$\hat{C}_{n}^{(P,Q)}\coloneqq\frac{1}{n_{1}}\sum_{i\in I_{1}}\widehat{\Psi}_{i}^{(P,Q)}\otimes\widehat{\Psi}_{i}^{(P,Q)}.$
Lemma B.13.
Let $P_{n},Q_{n}\to\infty$. Then, for every $\varepsilon>0$,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\left\|\hat{C}_{n}^{(P_{n},Q_{n})}-C_{P}\right\|_{\operatorname{tr}}>\varepsilon\right)\longrightarrow 0.$
Proof.
Fix $P\in C$. By the triangle inequality,
$\left\|\hat{C}_{n}^{(P_{n},Q_{n})}-C_{P}\right\|_{\operatorname{tr}}$
$\leq\left\|\hat{C}_{n}^{(P_{n},Q_{n})}-C_{n}^{\sharp,(P_{n},Q_{n})}\right\|_{\operatorname{tr}}+\left\|C_{n}^{\sharp,(P_{n},Q_{n})}-C_{n}^{(P_{n},Q_{n})}\right\|_{\operatorname{tr}}+\left\|C_{n}^{(P_{n},Q_{n})}-C_{n}^{(\infty)}\right\|_{\operatorname{tr}}+\left\|C_{n}^{(\infty)}-C_{P}\right\|_{\operatorname{tr}}.$
Step 1: truncation error. For any $a,b$ in a Hilbert space,
$\left\|a\otimes a-b\otimes b\right\|_{\operatorname{tr}}=\left\|a+b\right\|_{\textrm{HS}}\left\|a-b\right\|_{\textrm{HS}}.$
Apply this with $a=\Psi_{i}^{(P_{n},Q_{n})}$, $b=\Psi_{i}^{(\infty)}$ pointwise and use Cauchy–Schwarz over $i$ to obtain
$\left\|C_{n}^{(P_{n},Q_{n})}-C_{n}^{(\infty)}\right\|_{\operatorname{tr}}\leq\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\Psi_{i}^{(P_{n},Q_{n})}+\Psi_{i}^{(\infty)}\right\|_{\textrm{HS}}^{2}\right)^{1/2}\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\Psi_{i}^{(P_{n},Q_{n})}-\Psi_{i}^{(\infty)}\right\|_{\textrm{HS}}^{2}\right)^{1/2}.$
By the envelopes and contraction of projections,
$\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\Psi_{i}^{(P_{n},Q_{n})}+\Psi_{i}^{(\infty)}\right\|_{\textrm{HS}}^{2}\leq 16K_{X}K_{Y}K_{Z},$
and
$\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\Psi_{i}^{(P_{n},Q_{n})}-\Psi_{i}^{(\infty)}\right\|_{\textrm{HS}}^{2}\leq 8K_{Z}\left(K_{Y}A_{X,n}(P_{n})+K_{X}A_{Y,n}(Q_{n})\right),$
where $A_{X,n}(P)\coloneqq n^{-1}\sum_{i=1}^{n}\left\|(\Pi_{X}-P_{X,P})\epsilon_{X\mid Z}^{(i)}\right\|^{2}$ and $A_{Y,n}(Q)\coloneqq n^{-1}\sum_{i=1}^{n}\left\|(\Pi_{Y}-P_{Y,Q})\epsilon_{Y\mid Z}^{(i)}\right\|^{2}$. By Lemma B.8,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\operatorname{\mathbb{E}}_{P}\left\|(\Pi_{X}-P_{X,P_{n}})\epsilon_{X\mid Z}^{(1)}\right\|^{2}\to 0,\qquad\sup_{P\in\tilde{\mathcal{P}}_{0}}\operatorname{\mathbb{E}}_{P}\left\|(\Pi_{Y}-P_{Y,Q_{n}})\epsilon_{Y\mid Z}^{(1)}\right\|^{2}\to 0.$
Thus, Chebyshev’s inequality yields
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\left\|C_{n}^{(P_{n},Q_{n})}-C_{n}^{(\infty)}\right\|_{\operatorname{tr}}>\varepsilon\right)\ \longrightarrow\ 0.$
Step 2: estimated projector error. Applying the same rank–one identity with $a=\Psi_{i}^{\sharp,(P_{n},Q_{n})}$, $b=\Psi_{i}^{(P_{n},Q_{n})}$ and using the envelopes gives the deterministic bound
$\left\|C_{n}^{\sharp,(P_{n},Q_{n})}-C_{n}^{(P_{n},Q_{n})}\right\|_{\operatorname{tr}}\leq 32\,K_{X}K_{Y}K_{Z}\left(\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\operatorname{op}}+\left\|\hat{P}_{Y,Q_{n}}-P_{Y,Q_{n}}\right\|_{\operatorname{op}}\right),$
which converges to 0 in $\mathbb{P}_{P}$ uniformly over $\tilde{\mathcal{P}}_{0}$ by the same argument in Lemma B.5.
Step 3: regression–induced bias. Write
$\widehat{\Psi}_{i}^{(P_{n},Q_{n})}-\Psi_{i}^{\sharp,(P_{n},Q_{n})}=\left(\hat{P}_{X,P_{n}}b_{X,i}\right)\otimes\left(\hat{P}_{Y,Q_{n}}\epsilon_{Y\mid Z}^{(i)}\right)\otimes\phi_{\mathcal{Z}}(Z_{i})+\left(\hat{P}_{X,P_{n}}\epsilon_{X\mid Z}^{(i)}\right)\otimes\left(\hat{P}_{Y,Q_{n}}b_{Y,i}\right)\otimes\phi_{\mathcal{Z}}(Z_{i})$
$\qquad\qquad+\left(\hat{P}_{X,P_{n}}b_{X,i}\right)\otimes\left(\hat{P}_{Y,Q_{n}}b_{Y,i}\right)\otimes\phi_{\mathcal{Z}}(Z_{i}).$
Then
$\left\|\frac{1}{n_{1}}\sum_{i\in I_{1}}\left(\widehat{\Psi}_{i}^{(P_{n},Q_{n})}\otimes\widehat{\Psi}_{i}^{(P_{n},Q_{n})}-\Psi_{i}^{\sharp,(P_{n},Q_{n})}\otimes\Psi_{i}^{\sharp,(P_{n},Q_{n})}\right)\right\|_{\operatorname{tr}}$
$\leq\left\|\frac{1}{n_{1}}\sum_{i\in I_{1}}\left(\widehat{\Psi}_{i}^{(P_{n},Q_{n})}-\Psi_{i}^{\sharp,(P_{n},Q_{n})}\right)\right\|_{\textrm{HS}}\cdot M_{n},$
with
$M_{n}\coloneqq 2\left\|\frac{1}{n_{1}}\sum_{i\in I_{1}}\Psi_{i}^{\sharp,(P_{n},Q_{n})}\right\|+\left\|\frac{1}{n_{1}}\sum_{i\in I_{1}}\left(\widehat{\Psi}_{i}^{(P_{n},Q_{n})}-\Psi_{i}^{\sharp,(P_{n},Q_{n})}\right)\right\|_{\textrm{HS}}.$
By the envelopes, $\left\|\frac{1}{n_{1}}\sum_{i}\Psi_{i}^{\sharp,(P_{n},Q_{n})}\right\|\leq 4\sqrt{K_{X}K_{Y}K_{Z}}$ deterministically. Moreover, by Cauchy–Schwarz and contraction,
$\left\|\frac{1}{n_{1}}\sum_{i\in I_{1}}\left(\widehat{\Psi}_{i}^{(P_{n},Q_{n})}-\Psi_{i}^{\sharp,(P_{n},Q_{n})}\right)\right\|_{\textrm{HS}}$
$\leq\sqrt{K_{Z}}\Bigg[\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\hat{P}_{X,P_{n}}b_{X,i}\right\|^{2}\right)^{1/2}\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\epsilon_{Y\mid Z}^{(i)}\right\|^{2}\right)^{1/2}$
$\qquad+\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\epsilon_{X\mid Z}^{(i)}\right\|^{2}\right)^{1/2}\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\hat{P}_{Y,Q_{n}}b_{Y,i}\right\|^{2}\right)^{1/2}$
$\qquad+\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\hat{P}_{X,P_{n}}b_{X,i}\right\|^{2}\right)^{1/2}\left(\frac{1}{n_{1}}\sum_{i\in I_{1}}\left\|\hat{P}_{Y,Q_{n}}b_{Y,i}\right\|^{2}\right)^{1/2}\Bigg],$
which converges to 0 in $\mathbb{P}_{P}$ uniformly over $\tilde{\mathcal{P}}_{0}$. Hence
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\ \mathbb{P}_{P}\left(\left\|\hat{C}_{n}^{(P_{n},Q_{n})}-C_{n}^{\sharp,(P_{n},Q_{n})}\right\|_{\operatorname{tr}}>\varepsilon\right)\ \longrightarrow\ 0.$
Step 4: conclusion. Combining Steps 1–3 with Theorem 2 in Lundborg, Shah and Peters (2022) and Lemma S5 in Lundborg et al. (2024) proves the stated uniform convergence in probability. ∎
Lemma B.14.
Define the population and empirical uncentered covariance tensors by
$C_{XX}\coloneqq\operatorname{\mathbb{E}}\left[\phi_{\mathcal{X}}(X)\otimes\phi_{\mathcal{X}}(X)\right],\qquad\hat{C}_{XX}\coloneqq\frac{1}{n_{2}}\sum_{i\in I_{2}}\phi_{\mathcal{X}}(X_{i})\otimes\phi_{\mathcal{X}}(X_{i}).$
Then, for every $\varepsilon>0$,
$\mathbb{P}\left(\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}>\varepsilon\,\right)\leq\frac{K_{X}^{2}}{n\,\varepsilon^{2}},$
and hence
$\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}=O_{\mathcal{P}}\left(n_{2}^{-1/2}\right),\qquad\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}^{2}=O_{\mathcal{P}}\left(n_{2}^{-1}\right).$
The same statements apply to $\hat{C}_{YY}$ and $C_{YY}$.
Proof.
Work in the real Hilbert space $\mathrm{HS}$ with inner product characterized on elementary tensors by $\langle u_{1}\otimes v_{1},\,u_{2}\otimes v_{2}\rangle=\langle u_{1},u_{2}\rangle\langle v_{1},v_{2}\rangle$, and induced norm denoted $\|\cdot\|$. For $i\in I_{2}$ define
$Z_{i}\coloneqq\phi_{\mathcal{X}}(X_{i})\otimes\phi_{\mathcal{X}}(X_{i}),\qquad\bar{Z}\coloneqq\frac{1}{n_{2}}\sum_{i\in I_{2}}Z_{i},\qquad m\coloneqq\operatorname{\mathbb{E}}[Z_{1}]=C_{XX}.$
By the tensor norm identity $\left\|u\otimes v\right\|_{\textrm{HS}}=\|u\|\|v\|$ and the reproducing property,
$\|Z_{1}\|_{\textrm{HS}}=\left\|\phi_{\mathcal{X}}(X_{1})\right\|^{2}=k_{\mathcal{X}}\left(X_{1},X_{1}\right)\leq K_{X}\quad\text{a.s.}$
Hence $Z_{1}\in L^{2}(\Omega;\textrm{HS})$ and $m$ is well-defined as a Bochner expectation.
Since $\{Z_{i}\}_{i\in I_{2}}$ are i.i.d., we have
$\operatorname{\mathbb{E}}\left\|\bar{Z}-m\right\|^{2}=\frac{1}{n_{2}}\,\operatorname{\mathbb{E}}\|Z_{1}-m\|^{2}=\frac{1}{n_{2}}\left(\operatorname{\mathbb{E}}\|Z_{1}\|^{2}-\|m\|\|^{2}\right)\leq\frac{1}{n_{2}}\,\operatorname{\mathbb{E}}\|Z_{1}\|^{2}.$
Moreover, $\|Z_{1}\|^{2}\leq K_{X}^{2}$ a.s., so $\operatorname{\mathbb{E}}\|Z_{1}\|^{2}\leq K_{X}^{2}$. Therefore
$\operatorname{\mathbb{E}}\left\|\hat{C}_{XX}-C_{XX}\right\|^{2}=\operatorname{\mathbb{E}}\left\|\bar{Z}-m\right\|^{2}\leq\frac{K_{X}^{2}}{n}.$
Chebyshev’s inequality yields, for every $\varepsilon>0$,
$\mathbb{P}\left(\left\|\hat{C}_{XX}-C_{XX}\right\|>\varepsilon\right)\leq\varepsilon^{-2}\operatorname{\mathbb{E}}\left\|\hat{C}_{XX}-C_{XX}\right\|^{2}\leq\frac{K_{X}^{2}}{n\,\varepsilon^{2}}.$
This proves the stated probabilistic bound and the rates. The proof for $(\hat{C}_{YY},C_{YY})$ is identical with $K_{Y}$ in place of $K_{X}$. ∎
Lemma B.15.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. Let $\{X_{n}\}_{n\geq 1}$ be real-valued random variables and $\{\mathcal{F}_{n}\}_{n\geq 1}$ sub-$\sigma$-algebras of $\mathcal{F}$. Assume that for each $n$,
$\left|X_{n}\right|=Y_{n}+Z_{n}$
for some real-valued random variables $Y_{n},Z_{n}$. If $Y_{n}=o_{p}(1)$ and $\operatorname{\mathbb{E}}\left[Z_{n}\mid\mathcal{F}_{n}\right]=o_{p}(1)$, then $X_{n}=o_{p}(1)$.
Remark B.1.
This is the pointwise version of Lemma S6 in Lundborg et al. (2024). The proof is analogous to theirs and is therefore omitted.
Appendix C Characteristic kernels for Random Objects
We present (i) concise proofs for the main results of Section 5 and (ii) a brief discussion of implications for weak Fréchet regression in Bhattacharjee, Li and Xue (2025).
C.1 Proof of Main Theorem in Section 5
Throughout, we write the kernel mean embedding for a measure $P$ with respect to a kernel $k$ as $\mu^{k}(P)$ to make the dependence on both the kernel and the measure explicit. We first establish a key lemma that is foundational for the subsequent results: the property of being characteristic is preserved when a kernel is restricted to a Borel subset.
Lemma C.1.
Let $\mathcal{N},\mathcal{M}$ be Polish spaces and let $\phi\colon\mathcal{M}\to\mathcal{N}$ be an injective Borel measurable map. Let $k_{\mathcal{N}}\colon\mathcal{N}\times\mathcal{N}\to\mathbb{R}$ be a kernel that is characteristic (more generally, it suffices that $k_{\mathcal{N}}$ is characteristic on $\phi(\mathcal{M})$). Define the pull-back kernel
$(\phi^{\ast}k_{\mathcal{N}})(u,v)\coloneqq k_{\mathcal{N}}(\phi(u),\phi(v)),\qquad u,v\in\mathcal{M}.$
Consider the class of probability measures
$\mathcal{P}_{\phi^{\ast}k_{\mathcal{N}}}(\mathcal{M})\coloneqq\left\{P\in\mathcal{P}(\mathcal{M})\ \colon\ \int_{\mathcal{M}}\sqrt{(\phi^{\ast}k_{\mathcal{N}})(x,x)}\,P(\mathop{}\!\mathrm{d}x)<\infty\right\}.$
Then $\phi^{\ast}k_{\mathcal{N}}$ is characteristic on $\mathcal{M}$ with respect to $\mathcal{P}_{\phi^{\ast}k_{\mathcal{N}}}(\mathcal{M})$.
Remark C.1.
The class $\mathcal{P}_{\phi^{\ast}k_{\mathcal{N}}}(\mathcal{M})$ is precisely the domain on which the corresponding Bochner integrals are well-defined. If the kernel $\phi^{\ast}k_{\mathcal{N}}$ is bounded, then $\mathcal{P}_{\phi^{\ast}k_{\mathcal{N}}}(\mathcal{M})=\mathcal{P}(\mathcal{M})$, the set of all Borel probability measures on $\mathcal{M}$ (Sriperumbudur et al., 2010, Proposition 2).
Remark C.2.
In Lemma C.1, the assumption that $\mathcal{M}$ is Polish is used only to invoke standard results ensuring that $\phi(\mathcal{M})$ is Borel in $\mathcal{N}$ and that $\phi^{-1}\colon\phi(\mathcal{M})\to\mathcal{M}$ is Borel measurable. The argument itself requires only these measurability properties, not completeness of $\mathcal{M}$. In applications where the index set is not Polish (for instance, certain graph domains), one may instead verify directly that $\phi(\mathcal{M})$ is Borel and that $\phi$ is a Borel isomorphism onto its image; under these conditions, the conclusion of Lemma C.1 remains valid.
Proof of Lemma C.1.
By Srivastava (1998, Proposition 4.5.1 and Theorem 4.5.4), the injective Borel map $\phi\colon\mathcal{M}\to\mathcal{N}$ is a Borel isomorphism from $\mathcal{M}$ onto its image $\phi(\mathcal{M})$—that is, $\phi$ is a bijection onto $\phi(\mathcal{M})$ such that both $\phi$ and its inverse $\phi^{-1}\colon\phi(\mathcal{M})\to\mathcal{M}$ are Borel measurable—and $\phi(\mathcal{M})$ is a Borel subset of $\mathcal{N}$. Let $k_{\phi(\mathcal{M})}$ denote the restriction of $k_{\mathcal{N}}$ to $\phi(\mathcal{M})\times\phi(\mathcal{M})$, i.e.,
$k_{\phi(\mathcal{M})}(u,v)\coloneqq k_{\mathcal{N}}(u,v),\qquad u,v\in\phi(\mathcal{M}).$
First, $k_{\phi(\mathcal{M})}$ is characteristic on $\phi(\mathcal{M})$. Indeed, let $\nu_{1},\nu_{2}$ be probability measures on $\phi(\mathcal{M})$ whose kernel mean embeddings with respect to $k_{\phi(\mathcal{M})}$ coincide. Extend $\nu_{i}$ to probability measures on $\mathcal{N}$ by $\nu_{i}(B)\coloneqq\nu_{i}(B\cap\phi(\mathcal{M}))$ for all Borel $B\subset\mathcal{N}$ (zero extension outside $\phi(\mathcal{M})$); this is legitimate since $\phi(\mathcal{M})$ is Borel in $\mathcal{N}$ (Srivastava, 1998, Theorem 4.5.4). Then the corresponding embeddings with respect to $k_{\mathcal{N}}$ also coincide by Corollary 5.8 in Paulsen and Raghupathi (2016). As $k_{\mathcal{N}}$ is characteristic on $\mathcal{N}$ (or at least on $\phi(\mathcal{M})$), we conclude $\nu_{1}=\nu_{2}$ on $\phi(\mathcal{M})$.
Now let $P,Q\in\mathcal{P}_{\phi^{\ast}k_{\mathcal{N}}}(\mathcal{M})$ and assume their kernel mean embeddings with respect to $\phi^{\ast}k_{\mathcal{N}}$ coincide. For $R\in\{P,Q\}$, define the push-forward $\phi_{\sharp}R\coloneqq R\circ\phi^{-1}$, a probability measure on $\phi(\mathcal{M})$. Since
$\int_{\phi(\mathcal{M})}\sqrt{k_{\phi(\mathcal{M})}(v,v)}(\phi_{\sharp}R)(\mathop{}\!\mathrm{d}v)$ $=\int_{\mathcal{M}}\sqrt{k_{\mathcal{N}}(\phi(u),\phi(u))}\,R(\mathop{}\!\mathrm{d}u)$
$=\int_{\mathcal{M}}\sqrt{(\phi^{\ast}k_{\mathcal{N}})(u,u)}\,R(\mathop{}\!\mathrm{d}u)<\infty,$
the embeddings $\mu^{k_{\phi(\mathcal{M})}}(\phi_{\sharp}R)$ are well-defined for $R=P,Q$.
Since elements of the associated RKHSs are determined by their pointwise values, we evaluate the embeddings at $w\in\mathcal{M}$, and obtain for every $w\in\mathcal{M}$
$\mu^{\phi^{\ast}k_{\mathcal{N}}}(R)(w)$ $=\int_{\mathcal{M}}(\phi^{\ast}k_{\mathcal{N}})(u,w)\,R(\mathop{}\!\mathrm{d}u)$
$=\int_{\mathcal{M}}k_{\mathcal{N}}(\phi(u),\phi(w))\,R(\mathop{}\!\mathrm{d}u)$
$=\int_{\phi(\mathcal{M})}k_{\mathcal{N}}(z,\phi(w))\,\phi_{\sharp}R(\mathop{}\!\mathrm{d}z)$
$=\int_{\phi(\mathcal{M})}k_{\phi(\mathcal{M})}(z,\phi(w))\,\phi_{\sharp}R(\mathop{}\!\mathrm{d}z)$
$=\mu^{k_{\phi(\mathcal{M})}}(\phi_{\sharp}R)(\phi(w)).$
Hence $\mu^{k_{\phi(\mathcal{M})}}(\phi_{\sharp}P)\circ\phi=\mu^{k_{\phi(\mathcal{M})}}(\phi_{\sharp}Q)\circ\phi$ on $\mathcal{M}$. Since $\phi$ is surjective onto $\phi(\mathcal{M})$, we get $\mu^{k_{\phi(\mathcal{M})}}(\phi_{\sharp}P)=\mu^{k_{\phi(\mathcal{M})}}(\phi_{\sharp}Q)$ on $\phi(\mathcal{M})$. By the characteristicness of $k_{\phi(\mathcal{M})}$ on $\phi(\mathcal{M})$, we conclude $\phi_{\sharp}P=\phi_{\sharp}Q$.
Finally, since $\phi$ is a Borel isomorphism from $\mathcal{M}$ onto $\phi(\mathcal{M})$, for every Borel $B\in\mathcal{B}(\mathcal{M})$,
$P(B)=\phi_{\sharp}P(\phi(B))=\phi_{\sharp}Q(\phi(B))=Q(B),$
so $P=Q$. Therefore the kernel mean embedding associated with $\phi^{\ast}k_{\mathcal{N}}$ is injective on $\mathcal{P}_{\phi^{\ast}k_{\mathcal{N}}}(\mathcal{M})$, as claimed. ∎
Proof of Theorem 5.1.
Fix a base point $x_{0}\in\mathcal{M}$ and define
$\tilde{k}_{\rho}(u,v)\coloneqq\frac{1}{2}\left\{\rho(u,x_{0})+\rho(v,x_{0})-\rho(u,v)\right\},\qquad u,v\in\mathcal{M}.$
By Lemma 12 of Sejdinovic et al. (2013), $\tilde{k}_{\rho}$ is symmetric and positive definite; hence there exists an RKHS $\widetilde{\mathcal{H}}$ with feature map $\psi(u)=\tilde{k}_{\rho}(u,\cdot)\in\widetilde{\mathcal{H}}$. In fact, using $\|\psi(u)-\psi(v)\|_{\widetilde{\mathcal{H}}}^{2}=\rho(u,v)$, the continuity of $\rho$ yields the continuity of $\psi$, which immediately implies Borel measurability. Since $\mathcal{M}$ is Polish, $\widetilde{\mathcal{H}}$ is separable (Owhadi and Scovel, 2017, Theorem 2.4), so it is a separable Hilbert space.
Proposition 14 of Sejdinovic et al. (2013) further yields that $\psi$ is injective and
$\rho(u,v)=\left\lVert\psi(u)-\psi(v)\right\rVert_{\widetilde{\mathcal{H}}}^{2},\qquad u,v\in\mathcal{M}.$
For any $\gamma>0$ consider the Gaussian kernel on the RKHS $\widetilde{\mathcal{H}}$,
$\tilde{k}_{\gamma}(x,y)\coloneqq\exp\left(-\gamma\,\left\lVert x-y\right\rVert_{\widetilde{\mathcal{H}}}^{2}\right),\qquad x,y\in\widetilde{\mathcal{H}}$.
Define its pull-back along $\psi$ by
$(\psi^{\ast}\tilde{k}_{\gamma})(u,v)\coloneqq\tilde{k}_{\gamma}(\psi(u),\psi(v)),\qquad u,v\in\mathcal{M}.$
Then, using $\rho(u,v)=\left\lVert\psi(u)-\psi(v)\right\rVert_{\widetilde{\mathcal{H}}}^{2}$,
$(\psi^{\ast}\tilde{k}_{\gamma})(u,v)=\exp\left(-\gamma\,\left\lVert\psi(u)-\psi(v)\right\rVert_{\widetilde{\mathcal{H}}}^{2}\right)=\exp\left(-\gamma\,\rho(u,v)\right)\eqqcolon k_{\rho}(u,v).$
Thus $k_{\rho}=\psi^{\ast}\tilde{k}_{\gamma}$ is precisely the pull-back of the Gaussian kernel on $\widetilde{\mathcal{H}}$ along $\psi$.
By Theorem 9 of Wynne and Duncan (2022), Gaussian kernels on separable Hilbert spaces are characteristic; hence $\tilde{k}_{\gamma}$ is characteristic on $\widetilde{\mathcal{H}}$. Since $\psi$ is injective and Borel measurable between Polish spaces, Lemma C.1 applies and yields that the pull-back $k_{\rho}=\psi^{\ast}\tilde{k}_{\gamma}$ is characteristic on $\mathcal{M}$. As $\gamma>0$ was arbitrary, $k(u,v)=\exp\left(-\gamma\,\rho(u,v)\right)$ is characteristic for every $\gamma>0$. ∎
Remark C.3.
In Theorem 5.1, the continuity of $\rho$ was used only to guarantee separability of the Hilbert space underlying the Gaussian kernel constructed from the negative-type structure. This continuity assumption can be replaced by a purely metric assumption on $(\mathcal{M},\rho)$.
Indeed, suppose $\rho$ is a semimetric of negative type on $\mathcal{M}$ in the sense of Definition 5.1, and define
$d_{\rho}(u,v)\coloneqq\rho(u,v)^{1/2},\qquad u,v\in\mathcal{M}.$
By the standard representation theorem for negative-type semimetrics, there exist a real Hilbert space $\mathcal{H}$ and a map $\psi\colon\mathcal{M}\to\mathcal{H}$ such that
$\rho(u,v)=\left\|\psi(u)-\psi(v)\right\|_{\mathcal{H}}^{2}\quad\text{for all }u,v\in\mathcal{M}.$
In particular,
$d_{\rho}(u,v)=\left\|\psi(u)-\psi(v)\right\|_{\mathcal{H}},$
so $\psi$ is an isometric embedding of $(\mathcal{M},d_{\rho})$ into $\mathcal{H}$. Hence $\psi$ is a homeomorphism from $(\mathcal{M},d_{\rho})$ onto its image $\psi(\mathcal{M})$, and the subspace topology on $\psi(\mathcal{M})$ induced by $\mathcal{H}$ coincides with the metric topology of $d_{\rho}$.
If, in addition, $(\mathcal{M},d_{\rho})$ is separable, then $\psi(\mathcal{M})$ is separable, and so is the closed linear span $\overline{\operatorname{span}(\psi(\mathcal{M}))}$. Therefore the Gaussian kernel on this separable Hilbert space is characteristic, and its pull-back along $\psi$ coincides with $k_{\rho}(u,v)=\exp(-\gamma\,\rho(u,v))$, which is thus characteristic on $\mathcal{M}$. Consequently, in Theorem 5.1 the continuity of $\rho$ can be replaced by the assumption that $(\mathcal{M},d_{\rho})$ is a separable metric space.
Proof of Corollary 5.2.
Since $f$ is completely monotone on $(0,\infty)$ and admits the finite limit $f(0)=\lim_{t\downarrow 0}f(t)<\infty$, Bernstein’s theorem (Schilling, Song and Vondracek, 2012, Theorem 1.4) yields a unique positive Borel measure $\mu$ on $0,\infty)$ such that, for all $t>0$,
$f(t)=\int_{0}^{\infty}e^{-st}\mu(\mathop{}\!\mathrm{d}s).$
By monotone convergence,
$\lim_{t\downarrow 0}f(t)=\lim_{t\downarrow 0}\int_{0}^{\infty}e^{-st}\mu(\mathop{}\!\mathrm{d}s)=\int_{0}^{\infty}1\mu(\mathop{}\!\mathrm{d}s)=\mu([0,\infty))=f(0),$
hence $\mu([0,\infty))<\infty$. Non-constancy of $f$ implies $\mu((0,\infty))>0$; otherwise $\mu=c\,\delta_{0}$ and $f\equiv c$.
For all $u,v\in\mathcal{M}$ and $s\geq 0$, $0\leq e^{-s\rho(u,v)}\leq 1$, so with $\mu([0,\infty))<\infty$,
$0\leq k_{f}(u,v)=\int_{0}^{\infty}e^{-s\rho(u,v)}\mu(\mathop{}\!\mathrm{d}s)\leq\mu([0,\infty))=f(0)<\infty,$
and $k_{f}$ is bounded. Continuity of $\rho$ and of $f$ on $0,\infty)$ yields continuity of $(u,v)\mapsto f\left(\rho(u,v)\right)$. For each fixed $s\geq 0$, the kernel $k_{s}(u,v)\coloneqq e^{-s\,\rho(u,v)}$ is positive definite on $\mathcal{M}$ because $\rho$ is of negative type by Schoenberg’s theorem (Berg, Christensen and Ressel, 1984, Theorem 2.2). A nonnegative mixture of positive definite kernels is positive definite, hence
$k_{f}(u,v)=\int_{0}^{\infty}k_{s}(u,v)\mu(\mathop{}\!\mathrm{d}s)$
is positive definite.
Let $\mathcal{H}_{k}$ be the RKHS of a bounded positive definite kernel $k$ on $\mathcal{M}$, and let $\mu_{P}^{(k)}\in\mathcal{H}_{k}$ denote the kernel mean of a probability measure $P$ on $\mathcal{M}$. Set
$\mathrm{MMD}_{k}(P,Q)\coloneqq\left\|\mu_{P}^{(k)}-\mu_{Q}^{(k)}\right\|_{\mathcal{H}_{k}}.$
Using the quadratic form for $\mathrm{MMD}_{k}^{2}$ together with Tonelli’s theorem (justified by $0\leq k_{s}\leq 1$ and finiteness of $\mu$), we obtain, for all probability measures $P,Q$ on $\mathcal{M}$,
$\left(\mathrm{MMD}_{k_{f}}(P,Q)\right)^{2}=\int_{0}^{\infty}\left(\mathrm{MMD}_{k_{s}}(P,Q)\right)^{2}\mu(\mathop{}\!\mathrm{d}s).$
By Theorem 5.1, $k_{s}$ is characteristic for every $s>0$; in particular, $P\neq Q$ implies $\mathrm{MMD}_{k_{s}}(P,Q)>0$. Since $\mu((0,\infty))>0$, it follows that $P\neq Q$ implies
$\left(\mathrm{MMD}_{k_{f}}(P,Q)\right)^{2}=\int_{0}^{\infty}\left(\mathrm{MMD}_{k_{s}}(P,Q)\right)^{2}\mu(\mathop{}\!\mathrm{d}s)>0,$
whereas $P=Q$ gives $\mathrm{MMD}_{k_{s}}(P,Q)=0$ for all $s\geq 0$ and hence $\mathrm{MMD}_{k_{f}}(P,Q)=0$. Therefore the kernel $k_{f}$ is characteristic on $\mathcal{M}$. ∎
Remark C.4.
The following are equivalent: (A) $f$ is completely monotone on $\left(0,\infty\right)$ and has the finite limit $f(0)\coloneqq\lim_{t\downarrow 0}f(t)$. (B) There exists a unique finite positive Borel measure $\mu$ on $0,\infty)$ with $\mu((0,\infty))>0$ such that $f(t)=\int_{0}^{\infty}e^{-st}\mu(\mathop{}\!\mathrm{d}s)$ for all $t\in[0,\infty)$. The direction $(A)\Rightarrow(B)$ is already contained in the proof of Corollary 5.6. For completeness, $(B)\Rightarrow(A)$ follows as: (i) Continuity at 0: by monotone convergence, $\lim_{t\downarrow 0}f(t)=\int_{0}^{\infty}1\mu(\mathop{}\!\mathrm{d}s)=\mu([0,\infty))<\infty$. (ii) Complete monotonicity: for $t>0$ and $n\in\mathbb{N}_{0}$,
$f^{(n)}(t)=\int_{0}^{\infty}\frac{\partial^{n}}{\partial t^{n}}e^{-st}\mu(\mathop{}\!\mathrm{d}s)=(-1)^{n}\int_{0}^{\infty}s^{n}e^{-st}\mu(\mathop{}\!\mathrm{d}s),$
where differentiation under the integral is justified by dominated convergence (fix $t_{0}>0$ and bound by $s^{n}e^{-st_{0}/2}\in L^{1}(\mu)$). Hence $(-1)^{n}f^{(n)}(t)\geq 0$. (iii) Non-constancy: if $\mu((0,\infty))=0$ then $\mu=c\,\delta_{0}$ and $f\equiv c$; if $\mu((0,\infty))>0$ then for $0<t_{1}<t_{2}$, $f(t_{1})-f(t_{2})=\int_{(0,\infty)}\left(e^{-st_{1}}-e^{-st_{2}}\right)\mu(\mathop{}\!\mathrm{d}s)>0$. Uniqueness of $\mu$ is part of Bernstein’s theorem.
Proof of Corollary 5.3.
Since each $\mathcal{N}_{\ell}$ is Polish, the finite product $\mathcal{N}=\prod_{\ell=1}^{L}\mathcal{N}_{\ell}$ is Polish. By the assumptions of Corollary 5.2, for every $\ell$ the kernel
$k^{(\ell)}(u_{\ell},v_{\ell})=f_{\ell}(\rho_{\ell}(u_{\ell},v_{\ell})),\qquad u_{\ell},v_{\ell}\in\mathcal{N}_{\ell},$
is bounded, continuous, positive definite, and characteristic on $\mathcal{N}_{\ell}$.
Boundedness and continuity of
$k_{\otimes,L}(x,y)=\prod_{\ell=1}^{L}k^{(\ell)}(u_{\ell},v_{\ell}),\qquad u=(u_{1},\dots,u_{L}),\ v=(v_{1},\dots,v_{L}),$
follow from being a finite product of bounded continuous functions. For positive definiteness, Lemma 4.6 in Steinwart and Christmann (2008) yields that the tensor product of kernels is again a positive definite kernel on the product space; hence $k_{\otimes,L}$ is positive definite.
It remains to prove that $k_{\otimes,L}$ is characteristic. For each $\ell$, by Corollary 5.2 there exists a finite positive Borel measure $\mu_{\ell}$ on $0,\infty)$ with $\mu_{\ell}\left((0,\infty)\right)>0$ such that
$f_{\ell}(t)=\int_{0}^{\infty}e^{-st}\mu_{\ell}(\mathop{}\!\mathrm{d}s)\qquad(t>0).$
Let $\mu\coloneqq\bigotimes_{\ell=1}^{L}\mu_{\ell}$ on $0,\infty)^{L}$. Tonelli’s theorem gives
$k_{\otimes,L}(u,v)=\prod_{\ell=1}^{L}\int_{0}^{\infty}e^{-s_{\ell}\rho_{\ell}(u_{\ell},v_{\ell})}\mu_{\ell}(\mathop{}\!\mathrm{d}s_{\ell})=\int_{[0,\infty)^{L}}\exp\left(-\sum_{\ell=1}^{L}s_{\ell}\rho_{\ell}(u_{\ell},v_{\ell})\right)\mu(\mathop{}\!\mathrm{d}s),$
where $s=\left(s_{1},\dots,s_{L}\right)\in[0,\infty)^{L}$. Define
$\rho_{s}(x,y)\coloneqq\sum_{\ell=1}^{L}s_{\ell}\,\rho_{\ell}(u_{\ell},v_{\ell}),$
and the kernel $k_{s}(x,y)\coloneqq\exp\left(-\,\rho_{s}(x,y)\right)$. For $s=(s_{1},\dots,s_{L})\in[0,\infty)^{L}$, define
$\rho_{s}(x,y)\coloneqq\sum_{\ell=1}^{L}s_{\ell}\,\rho_{\ell}(u_{\ell},v_{\ell}),\qquad k_{s}(x,y)\coloneqq\exp\left(-\rho_{s}(x,y)\right).$
Each $\rho_{\ell}$ is a continuous semimetric of negative type, hence any non-negative linear combination $\rho_{s}$ is again a continuous semimetric of negative type. Moreover, if $s\in(0,\infty)^{L}$, then
$\rho_{s}(x,y)=0\ \Longrightarrow\ \rho_{\ell}(u_{\ell},v_{\ell})=0\ \text{for all }\ell\ \Longrightarrow\ x=y,$
so $\rho_{s}$ is a (negative-type) semimetric on $\mathcal{N}$ separating points (symmetry $\rho_{s}(x,y)=\rho_{s}(y,x)$ being clear). By Theorem 5.1, it follows that for every $s\in(0,\infty)^{L}$ the kernel $k_{s}(x,y)=\exp\left(-\rho_{s}(x,y)\right)$ is characteristic on $\mathcal{N}$.
Since each $\mu_{\ell}((0,\infty))>0$, we have
$\mu((0,\infty)^{L})=\prod_{\ell=1}^{L}\mu_{\ell}((0,\infty))>0.$
Using the quadratic form representation of $\mathrm{MMD}_{k_{\otimes,L}}^{2}$ and Tonelli’s theorem as in Corollary 5.2, we obtain
$\left(\mathrm{MMD}_{k_{\otimes,L}}(P,Q)\right)^{2}=\int_{[0,\infty)^{L}}\left(\mathrm{MMD}_{k_{s}}(P,Q)\right)^{2}\mu(\mathop{}\!\mathrm{d}s).$
If $P\neq Q$, then for all $s\in(0,\infty)^{L}$ the kernel $k_{s}$ is characteristic, hence $\mathrm{MMD}_{k_{s}}(P,Q)>0$ on $(0,\infty)^{L}$; since $\mu((0,\infty)^{L})>0$, the above integral is strictly positive. If $P=Q$, the integrand vanishes for all $s$ and the integral is zero. Therefore $k_{\otimes,L}$ is characteristic on $\mathcal{N}$. ∎
C.2 Examples: Negative-Type Semimetrics on the Hilbert Sphere and Metric-Valued Function Spaces for Random Objects
C.2.1 Negative-Type Semimetrics on the Hilbert sphere and Fisher Rao distance
Let $\mathcal{H}$ be a real Hilbert space and define the Hilbert sphere and its geodesic distance as
$\mathbb{S}(\mathcal{H})\coloneqq\left\{x\in\mathcal{H}\colon\ \|x\|_{\mathcal{H}}=1\right\},\qquad d_{\mathbb{S}}(u,v)=\arccos(\langle u,v\rangle_{\mathcal{H}}),\qquad u,v\in\mathbb{S}(\mathcal{H}).$
Proposition C.2.
For every $0<q\leq 1$, the semimetric $d_{\mathbb{S}}(u,v)^{q}$ on $\mathbb{S}(\mathcal{H})$ is of negative type.
Proof.
It suffices to prove the case $q=1$; the case $0<q<1$ then follows from Berg, Christensen and Ressel (1984, Corollary 2.10). Fix $x_{1},\dots,x_{n}\in\mathbb{S}(\mathcal{H})$ and $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$ with $\sum_{i=1}^{n}\alpha_{i}=0$. Using
$\arccos t=\frac{\pi}{2}-\arcsin t=\frac{\pi}{2}-\sum_{k=0}^{\infty}c_{k}\,t^{\,2k+1}\kern 5.0pt(|t|\leq 1),\kern 5.0ptc_{k}\coloneqq\frac{(2k)!}{4^{k}(k!)^{2}(2k+1)}>0,$
we obtain
$\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}d_{\mathbb{S}}(x_{i},x_{j})=\underbrace{\frac{\pi}{2}\left(\sum_{i=1}^{n}\alpha_{i}\right)\left(\sum_{j=1}^{n}\alpha_{j}\right)}_{=\,0}-\sum_{k=0}^{\infty}c_{k}\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}\,\langle x_{i},x_{j}\rangle^{2k+1}.$
For each $m\in\mathbb{N}$, the Gram matrix $G=(\langle x_{i},x_{j}\rangle)_{i,j}$ is positive semidefinite. By the Schur product theorem, all Hadamard powers $G^{\circ m}$ are again positive semidefinite (Horn and Johnson, 1985, Theorem 7.5.3(a)). Since $\bigl(G^{\circ m}\bigr)_{ij}=\langle x_{i},x_{j}\rangle^{m}$, it follows that
$\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}\,\langle x_{i},x_{j}\rangle^{2k+1}\geq 0,\quad\text{for every $k\geq 0$.}$
Since all coefficients $c_{k}>0$ and $|\langle x_{i},x_{j}\rangle|\leq 1$, the series is absolutely convergent and termwise nonpositive, whence $\sum_{i,j}\alpha_{i}\alpha_{j}d_{\mathbb{S}}(x_{i},x_{j})\leq 0$. Thus $d_{\mathbb{S}}$ is of negative type. Finally, if a measurable semimetric $\rho$ is of negative type, then $\rho^{\,q}$ is also of negative type for every $q\in(0,1]$ (Berg, Christensen and Ressel, 1984, Corollary 2.10). Applying this to $\rho=d_{\mathbb{S}}$ yields the claim for all $0<q\leq 1$. ∎
Corollary C.3.
Let $(\Omega,\mu)$ be a measure space and set
$\mathfrak{D}\coloneqq\left\{p\geq 0\,\colon\int_{\Omega}p\,\mathop{}\!\mathrm{d}\mu=1\right\}.$
Define the square-root embedding $\Phi\colon\mathfrak{D}\to\mathbb{S}(L^{2}(\mu))$ by $\Phi(p)=\sqrt{p}$. With the Fisher–Rao distance
$d_{\mathrm{FR}}(p,q)\coloneqq\arccos\left(\left\langle\sqrt{p},\sqrt{q}\right\rangle_{L^{2}}\right),$
the function $d_{\mathrm{FR}}^{\,q}$ is of negative type on $\mathfrak{D}$ for every $0<q\leq 1$.
Proof.
For all $p,q\in\mathfrak{D}$,
$\left\|\Phi(p)\right\|_{L^{2}(\mu)}^{2}=\int_{\Omega}p\,\mathop{}\!\mathrm{d}\mu=1,\quad d_{\mathrm{FR}}(p,q)=\arccos(\left\langle\Phi(p),\Phi(q)\right\rangle_{L^{2}})=d_{\mathbb{S}}(\Phi(p),\Phi(q)),$ (A.24)
so $\Phi$ is an isometric embedding from $(\mathfrak{D},d_{\mathrm{FR}})$ into the unit sphere $(\mathbb{S}(L^{2}(\mu)),d_{\mathbb{S}})$. By Proposition C.2, $d_{\mathbb{S}}^{\,q}$ is of negative type for all $0<q\leq 1$. Negative type is preserved under isometric pullbacks, hence $d_{\mathrm{FR}}^{\,q}$ is of negative type on $\mathfrak{D}$. ∎
In view of Proposition C.2 and Corollary C.3, when the underlying Hilbert space $\mathcal{H}$ (in particular, $L^{2}(\mu)$) is separable, Lemma C.1 applies. Consequently, one can naturally construct a characteristic kernel by pulling back a kernel from the Hilbert sphere via the square-root embedding.
C.2.2 The $L^{p}$ space for time-varying random objects
Let $(\mathcal{T},\rho_{\mathcal{T}})$ be a separable metric space equipped with a finite Radon measure $\nu$, and let $(\mathcal{M},d_{\mathcal{M}})$ be a metric space. For a fixed $1\leq p<\infty$, we consider the class of $\nu$-measurable maps $f\colon\mathcal{T}\to\mathcal{M}$, denoted by $\mathsf{Meas}_{\nu}(\mathcal{T},\mathcal{M})$.
Definition C.1 (Metric-valued $L^{p}$ space).
We define the metric-valued $L^{p}$ space as the set of equivalence classes of measurable maps,
$L^{p}(\mathcal{T},\mathcal{M})\coloneqq\left\{f\in\mathsf{Meas}_{\nu}(\mathcal{T},\mathcal{M}):\int_{\mathcal{T}}d_{\mathcal{M}}(f(t),o)^{p}\,\nu(\mathop{}\!\mathrm{d}t)<\infty\ \text{for some }o\in\mathcal{M}\right\}\big/\sim,$
where the equivalence relation is given by $f\sim g$ if and only if $d_{\mathcal{M}}(f(t),g(t))=0$ for $\nu$-almost every $t\in\mathcal{T}$. The space is equipped with the distance
$D_{p}(f,g)\coloneqq\left(\int_{\mathcal{T}}d_{\mathcal{M}}(f(t),g(t))^{p}\,\nu(\mathop{}\!\mathrm{d}t)\right)^{1/p}.$
We establish the following properties for the space $(L^{p}(\mathcal{T},\mathcal{M}),D_{p})$:
  1. 1.
    It is a metric space.
  2. 2.
    The space is well-defined regardless of the choice of $o\in\mathcal{M}$.
  3. 3.
    It is complete if $\mathcal{M}$ is complete.
  4. 4.
    It is separable if $\mathcal{M}$ is separable.
  5. 5.
    $D_{p}^{\,q}$ is of negative type for any $q\in(0,p]$ if $d_{\mathcal{M}}^{\,p}$ is of negative type.
Consequently, for a Polish space $\mathcal{M}$ where $d_{\mathcal{M}}^{\,p}$ is of negative type, the distance-induced exponential kernel $\exp(-\gamma D_{p}(f,g)^{q})$ and its generalization for any $q\in(0,p]$ are characteristic on this space.
Proposition C.4.
$D_{p}$ is a metric on $L^{p}(\mathcal{T},\mathcal{M})$.
Proof.
Non-negativity and symmetry are immediate. For the triangle inequality, for all $t$ we have $d_{\mathcal{M}}(f(t),h(t))\leq d_{\mathcal{M}}(f(t),g(t))+d_{\mathcal{M}}(g(t),h(t))$. Taking the $L^{p}$-norm in $t$ (Minkowski’s inequality) yields $D_{p}(f,h)\leq D_{p}(f,g)+D_{p}(g,h)$. Finally, if $D_{p}(f,g)=0$, then $d_{\mathcal{M}}(f(t),g(t))=0$ $\nu$-a.e.; by the definition of the quotient, this means $f=g$ in $L^{p}(\mathcal{T},\mathcal{M})$. ∎
Proposition C.5 (Independence of the base point).
Assume $\nu(\mathcal{T})<\infty$. If
$\int d_{\mathcal{M}}(f(t),o)^{p}\,\nu(\mathop{}\!\mathrm{d}t)<\infty,$ (A.25)
for some $o\in\mathcal{M}$, then for every $o^{\prime}\in\mathcal{M}$, $\int d_{\mathcal{M}}(f(t),o^{\prime})^{p}\,\nu(\mathop{}\!\mathrm{d}t)<\infty$. Thus the definition of $L^{p}(\mathcal{T},\mathcal{M})$ does not depend on the base point $o$.
Proof.
For all $t$, $d_{\mathcal{M}}(f(t),o^{\prime})\leq d_{\mathcal{M}}(f(t),o)+d_{\mathcal{M}}(o,o^{\prime})$. Hence, using $(a+b)^{p}\leq 2^{p-1}(a^{p}+b^{p})$,
$\int_{\mathcal{T}}d_{\mathcal{M}}(f(t),o^{\prime})^{p}\,\nu(\mathop{}\!\mathrm{d}t)\leq 2^{p-1}\int_{\mathcal{T}}d_{\mathcal{M}}(f(t),o)^{p}\,\nu(\mathop{}\!\mathrm{d}t)+2^{p-1}d_{\mathcal{M}}(o,o^{\prime})^{p}\,\nu(\mathcal{T})<\infty.$
Lemma C.6.
Let $f_{n}\colon\mathcal{T}\to\mathcal{M}$ be $\nu$-measurable and suppose $f_{n}(t)\to f(t)$ for $\nu$-a.e. $t$. Then $f$ is $\nu$-measurable.
Proof.
Define
$A\coloneqq\bigl\{t\in\mathcal{T}:(f_{n}(t))_{n\geq 1}\ \text{converges in }\mathcal{M}\bigr\}.$
Since $t\mapsto d_{\mathcal{M}}(f_{n}(t),f_{m}(t))$ is measurable for all $n,m$ and
$A=\bigcap_{k=1}^{\infty}\ \bigcup_{N=1}^{\infty}\ \bigcap_{n,m\geq N}\Bigl\{t:\ d_{\mathcal{M}}(f_{n}(t),f_{m}(t))<\tfrac{1}{k}\Bigr\},$
the set $A$ is measurable.
Fix $x_{0}\in\mathcal{M}$ and define $g\,\colon\mathcal{T}\to\mathcal{M}$ by
$g(t)=\begin{cases}\lim_{n\to\infty}f_{n}(t),&t\in A,\\ o,&t\notin A.\end{cases}$
We claim that $g$ is measurable. Let $F\subset\mathcal{M}$ be closed and set $h_{n}(t)\coloneqq d_{\mathcal{M}}(f_{n}(t),F)$ and $h(t)\coloneqq\liminf_{n\to\infty}h_{n}(t)$. Each $h_{n}$ is measurable (distance to a closed set is continuous), hence $h$ is measurable. For $t\in A$ we have $g(t)=\lim_{n}f_{n}(t)$, so $d_{\mathcal{M}}(g(t),F)=\lim_{n}h_{n}(t)=h(t)$. Therefore
$\{t:g(t)\in F\}=(A\cap\{h=0\})\cup(A^{c}\cap\{o\in F\}),$
which is measurable. As closed sets generate the Borel $\sigma$-algebra on $\mathcal{M}$, $g$ is Borel measurable.
By assumption there exists a $\nu$-null set $N$ such that $f_{n}(t)\to f(t)$ for all $t\in\mathcal{T}\setminus N$. In particular $\mathcal{T}\setminus N\subset A$ and $g(t)=f(t)$ for $t\in\mathcal{T}\setminus N$. Modifying $f$ on the null set $N$ (which does not affect $\nu$-measurability in the completion), we may assume $f=g$. Hence $f$ is $\nu$-measurable. ∎
Proposition C.7 (Completeness).
Assume $\nu(\mathcal{T})<\infty$ and that $(\mathcal{M},d_{\mathcal{M}})$ is complete. Then $\left(L^{p}(\mathcal{T},\mathcal{M}),D_{p}\right)$ is complete.
Proof.
Let $\{f_{n}\}$ be a $D_{p}$-Cauchy sequence. Choose a subsequence $\{f_{n_{k}}\}$ with $D_{p}(f_{n_{k+1}},f_{n_{k}})<2^{-k}$. Define $h_{k}(t)\coloneqq d_{\mathcal{M}}(f_{n_{k+1}}(t),f_{n_{k}}(t))\geq 0$. By Hölder and $\nu(\mathcal{T})<\infty$,
$\int_{\mathcal{T}}h_{k}\,\mathop{}\!\mathrm{d}\nu\leq\nu(\mathcal{T})^{1-1/p}\left(\int_{\mathcal{T}}h_{k}^{p}\,\mathop{}\!\mathrm{d}\nu\right)^{1/p}=\nu(\mathcal{T})^{1-1/p}D_{p}(f_{n_{k+1}},f_{n_{k}})\leq\nu(\mathcal{T})^{1-1/p}2^{-k}.$
Thus $\sum_{k}\int h_{k}\,\mathop{}\!\mathrm{d}\nu<\infty$, and by Tonelli, $\sum_{k}h_{k}(t)<\infty$ for $\nu$-a.e. $t$. Hence $\{f_{n_{k}}(t)\bigr\}$ is Cauchy in $\mathcal{M}$ for $\nu$-a.e. $t$, and completeness of $\mathcal{M}$ yields a pointwise limit $f(t)$. By Lemma C.6, $f$ is $\nu$-measurable.
We claim $D_{p}(f_{n_{k}},f)\to 0$. For a.e. $t$, $d_{\mathcal{M}}\left(f_{n_{k}}(t),f(t)\right)\leq\sum_{j\geq k}h_{j}(t)$. By Minkowski’s inequality for finite sums and then by monotone convergence,
$D_{p}(f_{n_{k}},f)=\left\|d_{\mathcal{M}}(f_{n_{k}},f)\right\|_{L^{p}}\leq\left\|\sum_{j\geq k}h_{j}\right\|_{L^{p}}\leq\sum_{j\geq k}\left\|h_{j}\right\|_{L^{p}}=\sum_{j\geq k}D_{p}(f_{n_{j+1}},f_{n_{j}})\xrightarrow[k\to\infty]{}0.$
Finally, since $\{f_{n}\}$ is Cauchy, for any $\varepsilon>0$ choose $K$ such that $D_{p}(f_{n},f_{n_{k}})<\varepsilon/2$ for all $n\geq N$, and pick $K$ large so that $D_{p}(f_{n_{k}},f)<\varepsilon/2$. Then $D_{p}(f_{n},f)\leq D_{p}(f_{n},f_{n_{k}})+D_{p}(f_{n_{k}},f)<\varepsilon$ for $n\geq N$. ∎
Proposition C.8 (Separability).
Assume $(\mathcal{T},\rho_{\mathcal{T}})$ is a separable metric space and $\nu$ is a finite Radon measure on $\mathcal{T}$. If $(\mathcal{M},d_{\mathcal{M}})$ is separable, then $L^{p}(\mathcal{T},\mathcal{M})$ equipped with $D_{p}$ is separable for every $1\leq p<\infty$.
Proof.
Since $\mathcal{M}$ is separable, fix a countable dense subset $D=\{y_{m}\}_{m\geq 1}\subset\mathcal{M}$. Because $\mathcal{T}$ is separable metric, it is second countable; let $\mathcal{B}=\{B_{n}\}_{n\geq 1}$ be a countable base of open sets in $\mathcal{T}$. Let $\mathcal{A}=\mathrm{alg}(\mathcal{B})$ be the algebra generated by $\mathcal{B}$ (finite unions and relative complements in $\mathcal{T}$); then $\mathcal{A}$ is countable. Define
$\mathfrak{S}\coloneqq\left\{\sum_{j=1}^{k}y_{j}\mathds{1}_{A_{j}}:k\in\mathbb{N},y_{j}\in D,A_{j}\in\mathcal{A}~~\text{pairwise disjoint, value }o\in D\text{ on }\mathcal{T}\setminus\textstyle\bigcup_{j}A_{j}\right\}.$
Clearly $\mathfrak{S}$ is countable. It remains to show that $\mathfrak{S}$ is dense in $L^{p}(\mathcal{T},\mathcal{M})$. Here, since $(\mathcal{M},d_{\mathcal{M}})$ is merely a metric space (not a vector space), the notation $\sum_{j=1}^{k}y_{j}\,\mathds{1}_{A_{j}}$ denotes the piecewise-constant map $t\mapsto y_{j}$ for $t\in A_{j}$ (and $o$ otherwise); the pairwise disjointness of $\{A_{j}\}$ ensures that this is well-defined, even without the linear structure of a vector space.
Density. Fix $f\in L^{p}(\mathcal{T},\mathcal{M})$ and $o\in\mathcal{M}$ with $\int_{\mathcal{T}}d_{\mathcal{M}}(f(t),o)^{p}\,\nu(\mathop{}\!\mathrm{d}t)<\infty$. Let $\varepsilon>0$ be arbitrary.
By monotone convergence, choose $R>0$ so that
$\int_{\mathcal{T}\setminus E_{R}}d_{\mathcal{M}}(f(t),o)^{p}\,\nu(\mathop{}\!\mathrm{d}t)<\left(\frac{\varepsilon}{4}\right)^{p},\qquad E_{R}\coloneqq\left\{t\in\mathcal{T}:\ d_{\mathcal{M}}(f(t),o)\leq R\right\}.$ (A.26)
Fix $\delta>0$ to be specified later. Because $D$ is dense, $\{B_{\mathcal{M}}(y_{m},\delta)\}_{m\geq 1}$ covers $\mathcal{M}$. Define $A^{\prime}_{m}\coloneqq f^{-1}\left(B_{\mathcal{M}}\left(y_{m},\delta\right)\right)\cap E_{R}$ (measurable), and define recursively
$A_{1}\coloneqq A^{\prime}_{1},\qquad A_{m}\coloneqq A^{\prime}_{m}\setminus\bigcup_{j=1}^{m-1}A_{j}\quad(m\geq 2).$
This makes the family $\{A_{m}\}_{m\geq 1}$ pairwise disjoint, and
$E_{R}=\bigsqcup_{m\geq 1}A_{m},$
where the union is disjoint.
Pick $N$ so that
$\nu\left(E_{R}\setminus\bigcup_{m=1}^{N}A_{m}\right)<\eta,$ (A.27)
where $\eta>0$ will be chosen later. Define the finite-valued map
$s_{R,\delta,N}(t)\coloneqq\begin{cases}y_{m},&t\in A_{m},\ m=1,\dots,N,\\ o,&t\in\mathcal{T}\setminus\bigcup_{m=1}^{N}A_{m}.\end{cases}$
Then, by (A.26), (A.27) and the construction of $\{A_{m}\}$,
$D_{p}\left(f,s_{R,\delta,N}\right)^{p}$ $\leq\int_{\bigcup_{m=1}^{N}A_{m}}d_{\mathcal{M}}\left(f(t),y_{m}\right)^{p}\,\nu(\mathop{}\!\mathrm{d}t)$
$\quad+\int_{E_{R}\setminus\bigcup_{m=1}^{N}A_{m}}d_{\mathcal{M}}(f(t),o)^{p}\,\nu(\mathop{}\!\mathrm{d}t)$
$\quad+\int_{\mathcal{T}\setminus E_{R}}d_{\mathcal{M}}(f(t),o)^{p}\,\nu(\mathop{}\!\mathrm{d}t)$
$\leq\delta^{p}\,\nu(\mathcal{T})+R^{p}\eta+\left(\frac{\varepsilon}{4}\right)^{p}.$
As $\nu(\mathcal{T})<\infty$, choosing $\delta,\eta$ sufficiently small gives $D_{p}\left(f,s_{R,\delta,N}\right)<\varepsilon/2$.
Fix a countable base $\{B_{n}\}_{n\geq 1}$ of open sets in $\mathcal{T}$ and let $\mathcal{A}\coloneqq\mathrm{alg}\left(\{B_{n}\}\right)$ be the algebra they generate (finite unions and relative complements in $\mathcal{T}$); then $\mathcal{A}$ is countable. Given a measurable $A\subset\mathcal{T}$ and $\theta>0$, inner and outer regularity of $\nu$ provide a compact $K\subset A$ and an open $O\supset A$ with
$\nu\left(A\setminus K\right)<\frac{\theta}{2},\qquad\nu\left(O\setminus A\right)<\frac{\theta}{2}.$
Because $\{B_{n}\}$ is a base and $K\subset O$, there are indices $n_{1},\dots,n_{J}$ with
$K\subset U\coloneqq\bigcup_{j=1}^{J}B_{n_{j}}\subset O,$
hence $U\in\mathcal{A}$ and
$\nu(A\triangle U)\leq\nu(A\setminus K)+\nu(O\setminus A)<\theta,$
where $\triangle$ denotes the symmetric difference.
Apply this with $A=A_{m}$ to obtain $S_{m}\in\mathcal{A}$ such that $\nu(A_{m}\triangle S_{m})<\theta$ for $m=1,\dots,N$. Define
$\widetilde{S}_{1}\coloneqq S_{1},\qquad\widetilde{S}_{m}\coloneqq S_{m}\setminus\bigcup_{j<m}S_{j}\quad(m\geq 2),$
so that $\widetilde{S}_{1},\dots,\widetilde{S}_{n}\in\mathcal{A}$ are pairwise disjoint. Define $E_{m}\coloneqq A_{m}\triangle S_{m}$. Then, for each $m$,
$A_{m}\triangle\widetilde{S}_{m}\subset E_{m}\cup\left(S_{m}\cap\bigcup_{j<m}S_{j}\right),$
and therefore
$\nu\left(A_{m}\triangle\widetilde{S}_{m}\right)\leq\nu(E_{m})\ +\ \sum_{j<m}\nu\left(S_{m}\cap S_{j}\right).$
Moreover, if $x\in S_{m}\cap S_{j}$ and $x\notin E_{m}\cup E_{j}$, then $x\in A_{m}\cap A_{j}=\varnothing$ (since the $A_{r}$ are disjoint), which is impossible; hence
$S_{m}\cap S_{j}\subset E_{m}\cup E_{j}\implies\nu(S_{m}\cap S_{j})\leq\nu(E_{m})+\nu(E_{j}).$
Summing in $m$ yields
$\sum_{m=1}^{N}\nu\left(A_{m}\triangle\widetilde{S}_{m}\right)$ $\leq\sum_{m=1}^{N}\nu(E_{m})+\sum_{m=1}^{N}\sum_{j<m}\bigl(\nu(E_{m})+\nu\left(E_{j}\right)\bigr)$
$=N\,\sum_{m=1}^{N}\nu(E_{m}).$
In particular, if $\nu(E_{m})\leq\theta$ for all $m$, then
$\sum_{m=1}^{N}\nu\left(A_{m}\triangle\widetilde{S}_{m}\right)\leq N^{2}\theta.$
Define $s\in\mathfrak{S}$ by $s(t)=y_{m}$ on $\widetilde{S}_{m}$ $(m=1,\dots,N)$ and $s(t)=o$ elsewhere. Note that whenever $\widetilde{S}_{m}\neq\emptyset$, necessarily $A_{m}\neq\emptyset$, hence there exists $t\in A_{m}$ with $d_{\mathcal{M}}(y_{m},o)\leq d_{\mathcal{M}}(y_{m},f(t))+d_{\mathcal{M}}(f(t),o)<\delta+R$. Therefore $d_{\mathcal{M}}(s_{R,\delta,N}(t),o)\leq R+\delta$ and $d_{\mathcal{M}}(s(t),o)\leq R+\delta$ for all $t$.
Let $U\coloneqq\bigcup_{m=1}^{N}\left(A_{m}\triangle\widetilde{S}_{m}\right)$. Then $s_{R,\delta,N}=s$ on $\mathcal{T}\setminus U$, and on $U$, $d_{\mathcal{M}}\left(s_{R,\delta,N}(t),s(t)\right)\leq 2(R+\delta).$ Hence
$D_{p}\left(s_{R,\delta,N},s\right)^{p}$ $=\int_{U}d_{\mathcal{M}}\left(s_{R,\delta,N}(t),s(t)\right)^{p}\,\nu(\mathop{}\!\mathrm{d}t)\leq\left(2(R+\delta)\right)^{p}\,\nu(U)$
$\leq\left(2(R+\delta)\right)^{p}\,\sum_{m=1}^{N}\nu\left(A_{m}\triangle\widetilde{S}_{m}\right)\leq\left(2(R+\delta)\right)^{p}\,N^{2}\theta.$
Choose $\theta\coloneqq\dfrac{\left(\varepsilon/2\right)^{p}}{\left(2(R+\delta)\right)^{p}N^{2}}$ to get $D_{p}\left(s_{R,\delta,N},s\right)<\varepsilon/2$.
By the triangle inequality,
$D_{p}(f,s)\leq D_{p}(f,s_{R,\delta,N})+D_{p}(s_{R,\delta,N},s)<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon.$
Since $s\in\mathfrak{S}$ and $\varepsilon>0$ was arbitrary, $\mathfrak{S}$ is dense in $L^{p}(\mathcal{T},\mathcal{M})$. As $\mathfrak{S}$ is countable, $L^{p}(\mathcal{T},\mathcal{M})$ is separable. ∎
Corollary C.9 (The Polish property and kernels).
If $\mathcal{M}$ is Polish, then $L^{p}(\mathcal{T},\mathcal{M})$ is Polish. Moreover, if $d_{\mathcal{M}}^{\,p}$ is of negative type, then for any $q\in(0,p]$, $D_{p}^{\,q}$ is of negative type on $L^{p}(\mathcal{T},\mathcal{M})$. Thus, $h\colon[0,\infty)\to\mathbb{R}$ is non-constant, completely monotone on $(0,\infty)$, and satisfies $h(0)=\lim_{t\downarrow 0}h(t)$, then
$k_{h}(f,g)\coloneqq h(D_{p}(f,g)^{q})$
is a bounded continuous characteristic kernel.
Proof.
Since Polishness follows from Theorem C.7 and C.8, to show that the generalized exponential kernel is positive definite and characteristic it suffices, by Theorem 5.4, to prove that the distance $D_{p}^{p}$ is of negative type.
Fix $n\in\mathbb{N}$, $f_{1},\dots,f_{n}\in L^{p}(\mathcal{T},\mathcal{M})$, and $c_{1},\dots,c_{n}\in\mathbb{R}$ with $\sum_{i=1}^{n}c_{i}=0$. Define
$S(t)\coloneqq\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}d_{\mathcal{M}}(f_{i}(t),f_{j}(t))^{p}\qquad(t\in\mathcal{T}).$
Since $d_{\mathcal{M}}^{\,p}$ is of negative type on $\mathcal{M}$, we have $S(t)\leq 0$ for each fixed $t$.
To justify integrability and interchange of the (finite) sum with the integral, fix $o\in\mathcal{M}$ and note that
$d_{\mathcal{M}}(f_{i}(t),f_{j}(t))^{p}\leq 2^{p-1}\left\{d_{\mathcal{M}}(f_{i}(t),o)^{p}+d_{\mathcal{M}}\left(f_{j}(t),o\right)^{p}\right\},$
so each map $t\mapsto d_{\mathcal{M}}(f_{i}(t),f_{j}(t))^{p}$ lies in $L^{1}(\nu)$ (because $f_{i}\in L^{p}(\mathcal{T},\mathcal{M})$). Hence $S\in L^{1}(\nu)$, and by linearity of the Lebesgue integral for finite sums,
$\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}\,D_{p}\left(f_{i},f_{j}\right)^{p}=\sum_{i,j}c_{i}c_{j}\int_{\mathcal{T}}d_{\mathcal{M}}(f_{i}(t),f_{j}(t))^{p}\,\nu(\mathop{}\!\mathrm{d}t)=\int_{\mathcal{T}}S(t)\,\nu(\mathop{}\!\mathrm{d}t)\leq 0.$
Therefore, $D_{p}^{\,p}$ is of negative type on $L^{p}(\mathcal{T},\mathcal{M})$. ∎
C.3 Discussion for Bhattacharjee, Li and Xue (2025)
This section clarifies the connection between characteristic kernels and the weak conditional Fréchet mean highlighted in Remark 5.4. Bhattacharjee, Li and Xue (2025) revisits the conditional Fréchet mean,
$m_{\oplus}(x)\coloneqq\operatorname*{\arg\min}_{y\in\mathcal{Y}}\operatorname{\mathbb{E}}[d_{\mathcal{Y}}(Y,y)^{2}\mid X=x],$
and proposes a weak conditional Fréchet mean by replacing the full conditional expectation with its orthogonal projection onto the RKHS $\mathcal{H}^{(k)}$ associated with a kernel $k$; see Bhattacharjee, Li and Xue (2025) for the complete construction.
They identify a sufficient condition under which the weak conditional Fréchet mean coincides with the conditional Fréchet mean: namely, that $\mathcal{H}^{(k)}$ is dense in $L^{2}\left(P_{X}\right)$ modulo constants. In particular, Bhattacharjee, Li and Xue (2025, Corollary 4) states that, if $\mathcal{H}^{(k)}$ is dense in $L^{2}(P_{X})$ modulo constants, then weak and full conditional Fréchet mean coincide.
To support the density requirement on metric spaces, Bhattacharjee, Li and Xue (2025) point to the construction of cc-universal kernels in Zhang, Xue and Li (2024). While cc-universality ensures uniform density on every compact subset of the response space (and yields rich scalarization families in that setting), cc-universality on a non-compact predictor domain does not, in general, imply $L^{2}$-density modulo constants for a given $P_{X}$ unless one additionally assumes compact support of $P_{X}$ (Sriperumbudur, Fukumizu and Lanckriet, 2011). Thus, cc-universality alone is insufficient to guarantee the required $L^{2}$-density on general non-compact predictor spaces.
An alternative and more direct route is to impose that the predictor kernel $k$ be characteristic. For any bounded measurable positive-definite kernel, Fukumizu, Bach and Jordan (2009, Proposition 5) establishes the equivalence
$\text{$k$ is characteristic}\Longleftrightarrow\mathcal{H}^{(k)}+\mathbb{R}~~\text{is dense in}~~L^{2}(P)~~\text{for every probability measure}~P,$
so that the $L^{2}$ modulo constants" is achieved uniformly over all $P$, without any compact-support assumption. Consequently, a characteristic predictor kernel furnishes a support-free sufficient condition under which the weak and the full conditional Fréchet means coincide, matching precisely the identification needed by Bhattacharjee, Li and Xue (2025).
The characteristic kernels we construct on metric spaces provide the exact $L^{2}$-density (modulo constants) required for the equality of weak and full conditional Fréchet means, and hence are directly applicable to Fréchet regression in the sense of Bhattacharjee, Li and Xue (2025).
Appendix D Uniform Spectral Gap and Growing Gap Condition
D.1 General Results
We consider a parametric distribution for $X$. If the parameter set is compact and the density functions are Hellinger continuous with respect to the parameter $\theta$, it can be shown that a minimum spectral gap and its associated distribution $P_{n}$ exist, satisfying Assumption 4.5.
Let $\Theta_{0}$ be a compact metric space. For each $\theta\in\Theta_{0}$, let $\rho_{\theta}$ be a Borel probability measure on $(\mathcal{X},\mathcal{B})$. Define the integral operator
$T_{\theta}\colon L^{2}(\rho_{\theta})\to L^{2}(\rho_{\theta}),\quad(T_{\theta}f)(x)\coloneqq\int_{\mathcal{X}}k(x,y)\,f(y)\mathop{}\!\mathrm{d}\rho_{\theta}(y).$
Assume that each $T_{\theta}$ is self-adjoint and compact on $L^{2}(\rho_{\theta})$, so that its eigenvalues
$\lambda_{1}(\theta)\geq\lambda_{2}(\theta)\geq\cdots\downarrow 0$
are well-defined, listed with algebraic multiplicity. For a fixed $P\in\mathbb{N}$, define the spectral gap
$\gamma_{P}(\theta)\coloneqq\lambda_{P}(\theta)-\lambda_{P+1}(\theta).$
We impose the following assumptions.
Assumption D.1.
  • (A1)
    There exists a $\sigma$-finite measure $\mu$ on $(\mathcal{X},\mathcal{B})$ such that
    $\rho_{\theta}\ll\mu\quad\text{for all }\theta\in\Theta_{0}.$
    Hence, for each $\theta\in\Theta_{0}$, there exists a density
    $p_{\theta}\coloneqq\frac{\mathop{}\!\mathrm{d}\rho_{\theta}}{\mathop{}\!\mathrm{d}\mu}\geq 0\quad\text{with}\quad\int_{\mathcal{X}}p_{\theta}\mathop{}\!\mathrm{d}\mu=1.$
  • (A2)
    The map
    $\Theta_{0}\ni\theta\ \longmapsto\ \sqrt{p_{\theta}}\in L^{2}(\mu)$
    is continuous, i.e.
    $\lim_{\vartheta\to\theta}\int_{\mathcal{X}}\left(\sqrt{p_{\theta}(x)}-\sqrt{p_{\vartheta}(x)}\right)^{2}\mathop{}\!\mathrm{d}\mu(x)=0\quad\text{for all }\theta\in\Theta_{0}.$
  • (A3)
    The kernel $k$ is bounded.
  • (A4)
    For any $P\in\mathbb{N}$, we assume
    $\lambda_{P}(\theta)>\lambda_{P+1}(\theta)\quad\text{for all }\theta\in\Theta_{0}.$
Theorem D.1 (Uniform spectral gap at level $P$).
Assume Assumption D.1. Then there exists a constant $c_{P}>0$ such that
$\lambda_{P}(\theta)-\lambda_{P+1}(\theta)\geq c_{P}\quad\text{for all }\theta\in\Theta_{0},$
i.e.
$\inf_{\theta\in\Theta_{0}}\bigl(\lambda_{P}(\theta)-\lambda_{P+1}(\theta)\bigr)>0.$
This theorem implies that there exists an increasing sequence of integers $P_{n}\to\infty$ such that
$\sqrt{n}c_{P_{n}}^{X}\to\infty.$
If we further restrict our attention to translation invariant kernels (i.e., $k(x,y)=\kappa(x-y)$), the assumptions can be extended to accommodate location shifts. This is because, for such kernels, the eigenvalues of the integral operator do not depend on the mean parameter (a location shift), as formalized in the following proposition.
Proposition D.2 (Mean invariance for translation invariant kernels).
Let $\rho$ be a Borel probability measure on $\left(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d})\right)$. For each $\mu\in\mathbb{R}^{d}$, define the translated measure $\rho_{\mu}$ by
$\rho_{\mu}(A)\coloneqq\rho(A-\mu),\qquad A\in\mathcal{B}(\mathbb{R}^{d}).$
Let $\kappa\colon\mathbb{R}^{d}\to\mathbb{R}$ be measurable and define the translation invariant kernel
$k(x,y)\coloneqq\kappa(x-y),\qquad x,y\in\mathbb{R}^{d}.$
For each $\mu\in\mathbb{R}^{d}$, define the integral operator
$T_{\mu}:L^{2}(\rho_{\mu})\to L^{2}(\rho_{\mu}),\qquad(T_{\mu}f)(x)\coloneqq\int_{\mathbb{R}^{d}}k(x,y)f(y)\mathop{}\!\mathrm{d}\rho_{\mu}(y),$
and assume $T_{\mu}$ is well-defined and bounded. Define also
$T_{0}:L^{2}(\rho)\to L^{2}(\rho),\qquad\left(T_{0}g\right)(x)\coloneqq\int_{\mathbb{R}^{d}}k(x,y)g(y)\mathop{}\!\mathrm{d}\rho(y).$
Then for each $\mu\in\mathbb{R}^{d}$ there exists a unitary operator
$U_{\mu}:L^{2}(\rho_{\mu})\to L^{2}(\rho),\qquad(U_{\mu}f)(x)\coloneqq f(x+\mu),$
such that
$U_{\mu}T_{\mu}U_{\mu}^{-1}=T_{0}.$
In particular, $T_{\mu}$ and $T_{0}$ are unitarily equivalent for all $\mu$, and hence have identical spectra, including eigenvalues with their algebraic multiplicities. Thus the eigenvalues of $T_{\mu}$ do not depend on the mean parameter $\mu$.
Proof.
Let $f\in L^{2}(\rho_{\mu})$. Then
$\left\lVert U_{\mu}f\right\rVert_{L^{2}(\rho)}^{2}=\int_{\mathbb{R}^{d}}\left\lvert f(x+\mu)\right\rvert^{2}\mathop{}\!\mathrm{d}\rho(x).$
By the definition of $\rho_{\mu}$,
$\int_{\mathbb{R}^{d}}g(y)\mathop{}\!\mathrm{d}\rho_{\mu}(y)=\int_{\mathbb{R}^{d}}g(x+\mu)\mathop{}\!\mathrm{d}\rho(x)\quad\text{for all bounded measurable }g.$
Taking $g(y)=\left\lvert f(y)\right\rvert^{2}$ gives
$\int_{\mathbb{R}^{d}}\left\lvert f(x+\mu)\right\rvert^{2}\mathop{}\!\mathrm{d}\rho(x)=\int_{\mathbb{R}^{d}}\left\lvert f(y)\right\rvert^{2}\mathop{}\!\mathrm{d}\rho_{\mu}(y)=\left\lVert f\right\rVert_{L^{2}(\rho_{\mu})}^{2}.$
Hence $U_{\mu}$ is an isometry.
For surjectivity, let $g\in L^{2}(\rho)$ and define $f\colon\mathbb{R}^{d}\to\mathbb{C}$ by $f(y)\coloneqq g(y-\mu)$. Then
$\int_{\mathbb{R}^{d}}\left\lvert f(y)\right\rvert^{2}d\rho_{\mu}(y)=\int_{\mathbb{R}^{d}}\left\lvert g(x)\right\rvert^{2}d\rho(x)<\infty,$
so $f\in L^{2}(\rho_{\mu})$ and
$(U_{\mu}f)(x)=f(x+\mu)=g(x).$
Thus $U_{\mu}$ is onto. Being an onto isometry, $U_{\mu}$ is unitary.
Let $g\in L^{2}(\rho)$ and put $f\coloneqq U_{\mu}^{-1}g$, so that $f(y)=g(y-\mu)$. Then
$(T_{\mu}f)(x)=\int_{\mathbb{R}^{d}}k(x,y)f(y)\mathop{}\!\mathrm{d}\rho_{\mu}(y)=\int_{\mathbb{R}^{d}}\kappa(x-y)g(y-\mu)\mathop{}\!\mathrm{d}\rho_{\mu}(y).$
Using $\rho_{\mu}(A)=\rho(A-\mu)$ and the change of variables $y=z+\mu$, we obtain
$\left(T_{\mu}f\right)(x)=\int_{\mathbb{R}^{d}}\kappa\left(x-\left(z+\mu\right)\right)g(z)\mathop{}\!\mathrm{d}\rho(z).$
Now
$\left(U_{\mu}T_{\mu}U_{\mu}^{-1}g\right)(x)=\left(U_{\mu}T_{\mu}f\right)(x)=\left(T_{\mu}f\right)(x+\mu)$
and hence
$\left(U_{\mu}T_{\mu}U_{\mu}^{-1}g\right)(x)=\int_{\mathbb{R}^{d}}\kappa\left((x+\mu)-\left(z+\mu\right)\right)g(z)\mathop{}\!\mathrm{d}\rho(z).$
By translation invariance of the kernel, $\kappa\left((x+\mu)-\left(z+\mu\right)\right)=\kappa(x-z)$, so
$\left(U_{\mu}T_{\mu}U_{\mu}^{-1}g\right)(x)=\int_{\mathbb{R}^{d}}\kappa(x-z)g(z)\mathop{}\!\mathrm{d}\rho(z)=\left(T_{0}g\right)(x).$
Thus $U_{\mu}T_{\mu}U_{\mu}^{-1}=T_{0}$.
Since $U_{\mu}$ is unitary, $T_{\mu}$ and $T_{0}$ are unitarily equivalent. Therefore they have identical spectra, including algebraic multiplicities of eigenvalues. In particular, the eigenvalues of $T_{\mu}$ do not depend on $\mu$. ∎
D.2 Proof of Theorem D.1
The proof relies on mapping the operator $T_{\theta}$, which acts on a $\theta$-dependent space $L^{2}(\rho_{\theta})$, to an equivalent operator $B_{\theta}$ on a fixed Hilbert space $L^{2}(\mu)$.
Define the isometry $J_{\theta}:L^{2}(\rho_{\theta})\to L^{2}(\mu)$ by $(J_{\theta}f)(x)\coloneqq\sqrt{p_{\theta}(x)}\,f(x)$, and its adjoint $J_{\theta}^{\ast}:L^{2}(\mu)\to L^{2}(\rho_{\theta})$. Let $B_{\theta}\coloneqq J_{\theta}T_{\theta}J_{\theta}^{\ast}$. This is a self-adjoint Hilbert–Schmidt operator on $L^{2}(\mu)$.
Two key lemmas, (Lemma D.3 and D.4 below), establish the following:
  1. 1.
    The nonzero eigenvalues of $T_{\theta}$ and $B_{\theta}$ coincide, including multiplicities (Lemma D.3).
  2. 2.
    Under assumptions (A1)–(A3), the map $\Theta_{0}\ni\theta\mapsto B_{\theta}$ is continuous in the Hilbert–Schmidt norm, i.e., $\lim_{\vartheta\to\theta}\lVert B_{\theta}-B_{\vartheta}\rVert_{\textrm{HS}}=0$ (Lemma D.4).
Since the Hilbert–Schmidt norm dominates the operator norm ($\lVert\cdot\rVert_{\textrm{op}}\leq\lVert\cdot\rVert_{\textrm{HS}}$), the map $\theta\mapsto B_{\theta}$ is also continuous in the operator norm. By standard perturbation results for compact self-adjoint operators, this implies that, for each fixed $j\in\mathbb{N}$, the eigenvalue map
$\Theta_{0}\ni\theta\ \longmapsto\ \lambda_{j}(\theta)$
is continuous. Consequently, the spectral gap function
$\Theta_{0}\ni\theta\ \longmapsto\ \gamma_{P}(\theta)=\lambda_{P}(\theta)-\lambda_{P+1}(\theta)$
is continuous.
By assumption (A4), we have $\gamma_{P}(\theta)>0$ for every $\theta\in\Theta_{0}$. Since $\Theta_{0}$ is compact (by premise) and $\gamma_{P}$ is continuous, the function $\gamma_{P}$ attains its minimum at some $\theta^{\ast}\in\Theta_{0}$. Since $\Theta_{0}$ is compact (by premise) and $\gamma_{P}$ is continuous, $\gamma_{P}$ attains its minimum at some $\theta^{\ast}\in\Theta_{0}$. By (A4), $\gamma_{P}(\theta)>0$ for all $\theta$, so in particular $\gamma_{P}(\theta^{\ast})>0$. We may set $c_{P}\coloneqq\gamma_{P}(\theta^{\ast})>0$. Hence, the minimum must be positive. ∎
The proof above relies on the following two lemmas.
Lemma D.3.
For each fixed $\theta\in\Theta_{0}$, the nonzero eigenvalues of $T_{\theta}$ and $B_{\theta}$ coincide, including algebraic multiplicities.
Proof.
Let $\lambda\neq 0$.
If $T_{\theta}f=\lambda f$ with $f\neq 0$, then $g\coloneqq J_{\theta}f\neq 0$ (since $J_{\theta}$ is an isometry) and
$B_{\theta}g=J_{\theta}T_{\theta}J_{\theta}^{\ast}J_{\theta}f=J_{\theta}T_{\theta}f=\lambda J_{\theta}f=\lambda g.$
Conversely, if $B_{\theta}g=\lambda g$ with $g\neq 0$, put $f\coloneqq J_{\theta}^{\ast}g$. Then
$T_{\theta}f=T_{\theta}J_{\theta}^{\ast}g=J_{\theta}^{\ast}J_{\theta}T_{\theta}J_{\theta}^{\ast}g=J_{\theta}^{\ast}B_{\theta}g=\lambda J_{\theta}^{\ast}g=\lambda f.$
We must ensure $f\neq 0$. If $f=0$, then $J_{\theta}^{\ast}g=0$. This implies $\langle B_{\theta}g,g\rangle_{\mu}=\langle J_{\theta}T_{\theta}J_{\theta}^{\ast}g,g\rangle_{\mu}=\langle T_{\theta}J_{\theta}^{\ast}g,J_{\theta}^{\ast}g\rangle_{\rho_{\theta}}=0$. From $B_{\theta}g=\lambda g$, we get $\lambda\|g\|_{\mu}^{2}=0$, which contradicts $\lambda\neq 0$ and $g\neq 0$. Thus $f\neq 0$, and $\lambda$ is an eigenvalue of $T_{\theta}$. The correspondences $f\mapsto J_{\theta}f$ and $g\mapsto J_{\theta}^{\ast}g$ restrict to inverse linear isomorphisms between the nonzero eigenspaces, so multiplicities coincide. ∎
Lemma D.4.
Under (A1)(A3), the map $\Theta_{0}\ni\theta\mapsto B_{\theta}$ is continuous in the Hilbert–Schmidt norm:
$\lim_{\vartheta\to\theta}\lVert B_{\theta}-B_{\vartheta}\rVert_{\textrm{HS}}=\lim_{\vartheta\to\theta}\lVert K_{\theta}-K_{\vartheta}\rVert_{L^{2}(\mu\times\mu)}=0.$
where $K_{\theta}(x,y)\coloneqq\sqrt{p_{\theta}(x)}\,k(x,y)\,\sqrt{p_{\theta}(y)}$ is the kernel of $B_{\theta}$.
Proof.
We have
$K_{\theta}(x,y)=\sqrt{p_{\theta}(x)}\,k(x,y)\,\sqrt{p_{\theta}(y)}.$
Since $\lvert k(x,y)\rvert\leq K$ (by (A3)), we obtain
$\lVert K_{\theta}-K_{\vartheta}\rVert_{L^{2}(\mu\times\mu)}\leq M\,\left\lVert\sqrt{p_{\theta}}\otimes\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\otimes\sqrt{p_{\vartheta}}\right\rVert_{L^{2}(\mu\times\mu)},$
where $\otimes$ denotes the tensor product (i.e., $(\sqrt{p_{\theta}}\otimes\sqrt{p_{\theta}})(x,y)=\sqrt{p_{\theta}(x)}\sqrt{p_{\theta}(y)}$). We use the triangle inequality:
$\sqrt{p_{\theta}}\otimes\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\otimes\sqrt{p_{\vartheta}}=\left(\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\right)\otimes\sqrt{p_{\theta}}+\sqrt{p_{\vartheta}}\otimes\left(\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\right).$
Thus, by the properties of the $L^{2}$ norm of tensor products,
$\left\lVert\sqrt{p_{\theta}}\otimes\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\otimes\sqrt{p_{\vartheta}}\right\rVert_{2}$ $\leq\left\lVert(\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}})\otimes\sqrt{p_{\theta}}\right\rVert_{2}+\left\lVert\sqrt{p_{\vartheta}}\otimes(\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}})\right\rVert_{2}$
$=\lVert\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\rVert_{2}\lVert\sqrt{p_{\theta}}\rVert_{2}+\lVert\sqrt{p_{\vartheta}}\rVert_{2}\lVert\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\rVert_{2}$
$=\left(\lVert\sqrt{p_{\theta}}\rVert_{2}+\lVert\sqrt{p_{\vartheta}}\rVert_{2}\right)\lVert\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\rVert_{2}.$
By (A1), $\lVert\sqrt{p_{\theta}}\rVert_{2}^{2}=\int p_{\theta}\mathop{}\!\mathrm{d}\mu=1$ (and similarly for $\vartheta$). Thus, the term in parentheses is $1+1=2$.
$\left\lVert\sqrt{p_{\theta}}\otimes\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\otimes\sqrt{p_{\vartheta}}\right\rVert_{2}\leq 2\,\lVert\sqrt{p_{\theta}}-\sqrt{p_{\vartheta}}\rVert_{2}.$
By assumption (A2) (Hellinger continuity), the right-hand side converges to 0 as $\vartheta\to\theta$. Hence $\lVert K_{\theta}-K_{\vartheta}\rVert_{2}\to 0$, which yields $\lVert B_{\theta}-B_{\vartheta}\rVert_{\textrm{HS}}\to 0$. ∎
D.3 Gaussian kernel under Gaussian measures: eigenvalues and a uniform gap rate
We now examine the divergence rate of $P_{n}$ necessary to fulfill the uniform growing gap condition for the Gaussian kernel under the family of Gaussian measures. As established in Proposition D.2, the eigenvalues remain invariant with respect to mean shifts. Consequently, we fix the mean $\mu$ and define our parameter set solely by the variance.
For $\gamma>0$, define the Gaussian kernel as
$k_{\gamma}(x,y)\coloneqq\exp(-\gamma^{-2}(x-y)^{2}),\qquad x,y\in\mathbb{R}.$
For mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}\in[\sigma_{\min}^{2},\sigma_{\max}^{2}]\subset\left(0,\infty\right)$, let $\nu_{\mu,\sigma}=\mathcal{N}(\mu,\sigma^{2})$ and define the integral operator
$\left(T_{\mu,\sigma}f\right)(x)\coloneqq\int_{\mathbb{R}}k_{\gamma}(x,y)f(y)\mathop{}\!\mathrm{d}\nu_{\mu,\sigma}(y),\qquad T_{\mu,\sigma}:L^{2}\left(\nu_{\mu,\sigma}\right)\to L^{2}(\nu_{\mu,\sigma}).$
Since $k_{\gamma}$ is bounded and continuous, $T_{\mu,\sigma}$ is compact and self-adjoint; thus its nonzero spectrum is a decreasing sequence $\lambda_{1}(\sigma)\geq\lambda_{2}(\sigma)\geq\cdots\downarrow 0$ (independent of $\mu$; see Proposition D.2). For $P\in\mathbb{N}$, write the spectral gap
$\gamma_{P}(\sigma)\coloneqq\lambda_{P}(\sigma)-\lambda_{P+1}(\sigma),\qquad c_{P}\coloneqq\inf_{\sigma\in\left[\sigma_{\min},\sigma_{\max}\right]}\gamma_{P}(\sigma).$
It is classical (see, e.g., Fasshauer and McCourt, 2012, Section 3) that, after the identification $L^{2}\left(\nu_{0,\sigma}\right)\cong L^{2}\left(\rho_{\alpha}\right)$ with
$\rho_{\alpha}\left(\mathop{}\!\mathrm{d}x\right)=\frac{\alpha}{\sqrt{\pi}}\,e^{-\alpha^{2}x^{2}}\,\mathop{}\!\mathrm{d}x,\qquad\alpha=\frac{1}{\sqrt{2}\,\sigma},$
the eigenfunctions are Hermite functions and the eigenvalues form a geometric sequence:
$\lambda_{n}(\sigma)=\lambda_{1}(\sigma)\,r(\sigma)^{\,n-1},\qquad n\geq 1,$ (A.28)
with ratio $r(\sigma)\in(0,1)$ given explicitly by
$r(\sigma)=\frac{\varepsilon^{2}}{\varepsilon^{2}+\alpha^{2}+\delta^{2}},\qquad\delta^{2}=\frac{\alpha^{2}}{2}\Bigl(\sqrt{1+\left(\tfrac{2\varepsilon}{\alpha}\right)^{2}}-1\Bigr),\quad\alpha=\frac{1}{\sqrt{2}\,\sigma},$ (A.29)
and
$\lambda_{1}(\sigma)=\sqrt{\frac{\alpha^{2}}{\varepsilon^{2}+\alpha^{2}+\delta^{2}}}.$ (A.30)
By compactness of the variance interval and continuity of (A.29)–(A.30) in $\sigma$, there exist
$r_{\ast}\coloneqq\sup_{\sigma\in\left[\sigma_{\min},\sigma_{\max}\right]}r(\sigma)\in(0,1),\qquad c_{\ast}\coloneqq\inf_{\sigma\in\left[\sigma_{\min},\sigma_{\max}\right]}\lambda_{1}(\sigma)\left(1-r(\sigma)\right)>0.$ (A.31)
From (A.28),
$\gamma_{P}(\sigma)=\lambda_{P}(\sigma)-\lambda_{P+1}(\sigma)=\lambda_{1}(\sigma)\left(1-r(\sigma)\right)\,r(\sigma)^{\,P-1}.$ (A.32)
Taking the infimum over $\sigma\in\left[\sigma_{\min},\sigma_{\max}\right]$ and using (A.31) yields
$\gamma_{P}\geq c_{\ast}\,r_{\ast}^{\,P-1}.$ (A.33)
Theorem D.5.
Let $P_{n}\to\infty$ be integers. Then $\sqrt{n}\,c_{P_{n}}\to\infty$ holds provided
$r_{\ast}^{\,P_{n}}\ \gg\ n^{-1/2}\quad\text{equivalently}\quad P_{n}\ \ll\ \frac{\tfrac{1}{2}\log n}{\log\left(1/r_{\ast}\right)}.$
In particular, for any fixed $\delta\in\left(0,\tfrac{1}{2}\right)$,
$P_{n}\ \coloneqq\ \left\lfloor\frac{\left(\tfrac{1}{2}-\delta\right)\log n}{\log\left(1/r_{\ast}\right)}\right\rfloor\quad\Longrightarrow\quad\sqrt{n}\,c_{P_{n}}\geq c_{\ast}\,r_{\ast}^{-1}\,n^{\delta}\ \xrightarrow[n\to\infty$]{}\ $\infty$. (A.34)
Conversely, if $P_{n}\sim\frac{\tfrac{1}{2}\log n}{\log\left(1/r_{\ast}\right)}$, then $\sqrt{n}\,\gamma_{P_{n}}$ stays of constant order; if $P_{n}$ grows faster, then $\sqrt{n}\,c_{P_{n}}\to 0$.
Proof.
Inequality (A.33) gives $\sqrt{n}\,\gamma_{P_{n}}\geq c_{\ast}\,r_{\ast}^{\,P_{n}-1}\sqrt{n}=c_{\ast}\,r_{\ast}^{-1}\,\sqrt{n}\,r_{\ast}^{\,P_{n}}.$ The stated conditions are precisely those that force the right-hand side to diverge. The borderline and super-logarithmic cases follow similarly. ∎
Appendix E KRR-type CME: sufficient conditions for Assumptions 4.2 and 4.3
We derive explicit convergence rates for the KRR-type CME estimator of $\mu_{Y\mid Z}$ and use them to furnish sufficient conditions for Assumptions 4.2 and 4.3. For, $\mu_{X\mid Z}$ In this section, for simplicity of notation, we assume that $I_{1}=\{1,\dots,n\}$.
Define the population operators
$C_{ZZ}\coloneqq\operatorname{\mathbb{E}}[\phi_{\mathcal{Z}}(Z)\otimes\phi_{\mathcal{Z}}(Z)]\in\mathcal{H}_{\mathcal{Z}}\otimes\mathcal{H}_{\mathcal{Z}},\qquad C_{YZ}\coloneqq\operatorname{\mathbb{E}}[\phi_{\mathcal{Y}}(Y)\otimes\phi_{\mathcal{Z}}(Z)]\in\mathcal{H}_{\mathcal{Y}}\otimes\mathcal{H}_{\mathcal{Z}},$
and their empirical counterparts
$\hat{C}_{ZZ}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\phi_{\mathcal{Z}}(Z_{i})\otimes\phi_{\mathcal{Z}}(Z_{i}),\qquad\hat{C}_{YZ}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\phi_{\mathcal{Y}}(Y_{i})\otimes\phi_{\mathcal{Z}}(Z_{i}).$
For $\lambda>0$, the HS-ridge (KRR-type) estimator and its population analogue are
$\hat{C}_{Y\mid Z,\lambda}=\hat{C}_{YZ}\left(\hat{C}_{ZZ}+\lambda I\right)^{-1},\qquad C_{Y\mid Z,\lambda}=C_{YZ}\left(C_{ZZ}+\lambda I\right)^{-1},$
and the associated CME maps are $\hat{\mu}_{Y\mid Z}(z)=\hat{C}_{Y\mid Z,\lambda}\,\phi_{\mathcal{Z}}(z)$ and $\mu_{Y\mid Z,\lambda}(z)=C_{Y\mid Z,\lambda}\,\phi_{\mathcal{Z}}(z)$.
Assumptions. We impose the following conditions with respect to $C_{ZZ}$ and the target operator $F^{\ast}\in S_{2}(\mathcal{H}_{\mathcal{Z}},\mathcal{H}_{\mathcal{Y}})$ that solves $C_{YZ}=F^{\ast}C_{ZZ}$ on $\overline{\operatorname{range}\left(C_{ZZ}\right)}$ (equivalently $F^{\ast}=C_{YZ}C_{ZZ}^{\dagger}$ there).
Assumption E.1 (Smoothness (source) condition).
There exists $\beta\in(1,2]$ such that
$\left\|F^{\ast}\right\|_{\beta}^{2}\coloneqq\sum_{j=1}^{\infty}\mu_{j}^{-\beta}\,\left\|F^{\ast}e_{j}\right\|^{2}<\infty,$
where $C_{ZZ}e_{j}=\mu_{j}e_{j}$ with $\{e_{j}\}_{j\geq 1}\subset\mathcal{H}_{\mathcal{Z}}$ orthonormal and $\mu_{1}\geq\mu_{2}\geq\cdots\geq 0$.
Assumption E.2 (Eigenvalue decay).
Let $C_{ZZ}:\mathcal{H}_{\mathcal{Z}}\to\mathcal{H}_{\mathcal{Z}}$ be the integral operator with spectral decomposition $C_{ZZ}e_{j}=\mu_{j}e_{j}$, where $\{e_{j}\}_{j\geq 1}\subset\mathcal{H}_{\mathcal{Z}}$ is orthonormal and $\mu_{1}\geq\mu_{2}\geq\cdots\geq 0$. There exist constants $p\in(0,1]$ and $c_{2}>0$ such that
$\mu_{j}\leq c_{2}\,j^{-1/p}\qquad\text{for all }j\in\mathbb{N}.$
Equivalently (up to universal constants), there exists $Q>0$ such that the effective dimension
$N_{Z}(\lambda)\coloneqq\mathrm{tr}\left(C_{ZZ}\left(C_{ZZ}+\lambda I\right)^{-1}\right)\leq Q\,\lambda^{-p}\qquad\text{for all }\lambda\in(0,1].$
Embedding (EMB). Since kernels on $\mathcal{Z}$ and $\mathcal{Y}$ are bounded, evaluation functionals are uniformly bounded by constants $K_{Z},K_{Y}<\infty$. In particular, the embedding $\mathcal{H}_{\mathcal{Z}}]^{\alpha}\hookrightarrow L^{1}(\pi_{Z})$ holds with $\alpha=1$ under boundedness, and we adopt $\alpha=1$ throughout.
Under Assumptions E.1 and E.2, Li et al. (2022) establish that choosing
$\lambda_{n}\asymp n^{-1/(\beta+p)}$
balances the bias term $\lambda^{\beta}$ and the variance term $N_{Z}(\lambda)/n$ and yields the population rate
$\left\|\left(\hat{C}_{Y\mid Z,\lambda_{n}}-F^{\ast}\right)C_{ZZ}^{1/2}\right\|_{\textrm{HS}}^{2}=O_{p}\left(n^{-\beta/(\beta+p)}\right).$ (A.35)
Any choice $\lambda_{n}$ of this order attains the same rate.
Consequently, to bridge Assumptions 4.2 and 4.3 with (A.35), it suffices to show that the gap between the empirical and population $L^{2}$ errors is asymptotically negligible at the target rate; concretely,
$\left|R_{n}-R\right|=o_{p}(R),$
where $R=\left\|\left(\hat{C}_{Y\mid Z,\lambda}-C_{Y\mid Z}\right)C_{ZZ}^{1/2}\right\|_{\textrm{HS}}^{2}$ and $R_{n}=\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\mu}_{Y\mid Z}(Z_{i})-\mu_{Y\mid Z}(Z_{i})\right\|^{2}$.
Proposition E.1.
Assume Assumption 2.1, 2.2 (so that evaluation functionals are uniformly bounded by constants $K_{Z},K_{Y}$) and Assumptions E.1, E.2. Define
$B_{U}\coloneqq\hat{C}_{Y\mid Z,\lambda}-C_{Y\mid Z}\in\mathrm{HS}\left(\mathcal{H}_{\mathcal{Z}},\mathcal{H}_{\mathcal{Y}}\right),\qquad C_{ZZ,\lambda}\coloneqq C_{ZZ}+\lambda I,$
and the population and empirical risks
$R\coloneqq\left\|B_{U}\,C_{ZZ}^{1/2}\right\|_{\mathrm{HS}}^{2}=\operatorname{\mathbb{E}}_{Z}\left\|\hat{\mu}_{Y\mid Z}(Z)-\mu_{Y\mid Z}(Z)\right\|^{2},\quad R_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\mu}_{Y\mid Z}(Z_{i})-\mu_{Y\mid Z}(Z_{i})\right\|^{2}.$
Then for any $\tau\geq 1$, with probability at least $1-c\,e^{-\tau}$,
$R\leq C\left\{\lambda^{\beta}+\frac{N_{Z}(\lambda)}{n}+\frac{1}{n^{2}\lambda}\right\},$
and
$\left|R_{n}-R\right|\leq C\left(\sqrt{\frac{N_{Z}(\lambda)\,\tau}{n}}+\frac{K_{Z}^{2}\,\tau}{\lambda n}\right)\,\left(R+\lambda\|B_{U}\|_{\mathrm{HS}}^{2}\right).$
In particular, choosing $\lambda_{n}\asymp n^{-1/(\beta+p)}$ yields
$R=O_{p}\left(n^{-\beta/(\beta+p)}\right),\qquad\left|R_{n}-R\right|=o_{p}(R)\quad\text{and hence}\quad R_{n}=O_{p}\left(n^{-\beta/(\beta+p)}\right).$
For KRR-type CME estimators with regularization choices $\lambda_{X,n}\asymp n^{-1/(\beta_{X}+p)}$ and $\lambda_{Y,n}\asymp n^{-1/(\beta_{Y}+p)}$, where $\beta_{X},\beta_{Y}>1$ are source exponents and $p\in(0,1]$ is the capacity exponent of $C_{ZZ}$, the empirical mean-squared errors satisfy
$R_{X,n}$ $\coloneqq\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\mu}_{X\mid Z}\left(Z_{i}\right)-\mu_{X\mid Z}\left(Z_{i}\right)\right\|^{2}=O_{p}\left(n^{-\alpha}\right)$
$R_{Y,n}$ $\coloneqq\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\mu}_{Y\mid Z}\left(Z_{i}\right)-\mu_{Y\mid Z}\left(Z_{i}\right)\right\|^{2}=O_{p}\left(n^{-\beta}\right),$
with
$\alpha=\frac{\beta_{X}}{\beta_{X}+p},\qquad\beta=\frac{\beta_{Y}}{\beta_{Y}+p}.$
Using the previously established bound for the cross-term, the contribution of CME estimation errors to the $\sqrt{n}$-normalized statistic obeys
$n\,\left\|\frac{1}{n}\sum_{i=1}^{n}\left(\hat{\mu}_{X\mid Z}\left(Z_{i}\right)-\mu_{X\mid Z}\left(Z_{i}\right)\right)\otimes\left(\hat{\mu}_{Y\mid Z}\left(Z_{i}\right)-\mu_{Y\mid Z}\left(Z_{i}\right)\right)\right\|_{\textrm{HS}}^{2}=O_{p}\left(n^{\,1-(\alpha+\beta)}\right).$
Asymptotic negligibility $\left(o_{p}(1)\right)$ requires a strictly negative exponent, i.e.,
$\alpha+\beta>1\quad\Longleftrightarrow\quad\frac{\beta_{X}}{\beta_{X}+p}+\frac{\beta_{Y}}{\beta_{Y}+p}>1.$
Under this condition, Assumption 4.2 with rates implies Assumption 4.3; the estimation-error interaction does not affect the limiting distribution, and Assumptions 4.2, 4.3 are verified by concrete requirements on $\beta_{X},\beta_{Y}$, and $p$. Moreover, invoking Proposition B.4, it is straightforward to verify that, in the setting of Pogodin et al. (2025), the CME bias is asymptotically negligible for the null distribution under the above assumptions.
Proof of Proposition E.1.
By the standard rank-one expansion,
$R=\mathrm{Tr}\left(B_{U}\,C_{ZZ}\,B_{U}^{\ast}\right)=\left\|B_{U}\,C_{ZZ}^{1/2}\right\|_{\textrm{HS}}^{2},\qquad R_{n}-R=\mathrm{Tr}\left(B_{U}\left(\hat{C}_{ZZ}-C_{ZZ}\right)B_{U}^{\ast}\right).$
Insert $C_{ZZ,\lambda}^{1/2}C_{ZZ,\lambda}^{-1/2}$ on both sides of $\hat{C}_{ZZ}-C_{ZZ}$ and set
$E_{\lambda}\coloneqq C_{ZZ,\lambda}^{-1/2}\left(\hat{C}_{ZZ}-C_{ZZ}\right)C_{ZZ,\lambda}^{-1/2},\qquad A_{\lambda}\coloneqq B_{U}\,C_{ZZ,\lambda}^{1/2}.$
Then $E_{\lambda}$ is self-adjoint and
$R_{n}-R=\mathrm{Tr}\left(A_{\lambda}E_{\lambda}A_{\lambda}^{\ast}\right).$
By Loewner ordering and the trace inequality $\left|\mathrm{Tr}(XY)\right|\leq\left\|X\right\|_{\textrm{HS}}\left\|Y\right\|_{\textrm{HS}}$,
$\left|R_{n}-R\right|\leq\left\|E_{\lambda}\right\|_{\textrm{op}}\,\mathrm{Tr}\left(A_{\lambda}A_{\lambda}^{\ast}\right)=\left\|E_{\lambda}\right\|_{\textrm{op}}\,\left\|A_{\lambda}\right\|_{\textrm{HS}}^{2}=\left\|E_{\lambda}\right\|_{\textrm{op}}\left(R+\lambda\|B_{U}\|_{\textrm{HS}}^{2}\right).$ (A.36)
Define $h(z)\coloneqq C_{ZZ,\lambda}^{-1/2}\phi_{Z}(z)$. Using bounded kernels and the bound $\left\|C_{ZZ,\lambda}^{-1}\right\|_{\textrm{op}}\leq\lambda^{-1}$,
$\sup_{Z}\left\|h(z)\right\|^{2}=\sup_{Z}\left\langle\phi_{Z}(z),\,C_{ZZ,\lambda}^{-1}\phi_{Z}(z)\right\rangle\leq\lambda^{-1}\sup_{Z}\left\|\phi_{Z}(z)\right\|^{2}\leq\frac{K_{Z}^{2}}{\lambda},$
and, by definition of the effective dimension,
$\operatorname{\mathbb{E}}\left\|h(Z)\right\|^{2}=\mathrm{Tr}\left(C_{ZZ,\lambda}^{-1/2}C_{ZZ}C_{ZZ,\lambda}^{-1/2}\right)=N_{Z}(\lambda).$
A Hilbert-space Bernstein inequality then gives, for any $\tau\geq 1$, with probability at least $1-2e^{-\tau}$,
$\left\|E_{\lambda}\right\|_{\textrm{op}}\leq C\left(\sqrt{\frac{N_{Z}(\lambda)\,\tau}{n}}+\frac{K_{Z}^{2}\,\tau}{\lambda n}\right).$ (A.37)
Decompose $B_{U}=B_{\mathrm{est}}+B_{\mathrm{bias}}$. For the estimation part one has, with the same probability level,
$\left\|B_{\mathrm{est}}\,C_{ZZ}^{1/2}\right\|_{\textrm{HS}}^{2}\leq\frac{C}{n}\left(K_{Y}^{2}N_{Z}(\lambda)+\frac{K_{Z}^{2}}{\lambda}\,\left\|F^{\ast}-F_{\lambda}\right\|_{L^{2}(\pi_{Z};\mathcal{H}_{Y})}^{2}+\frac{K_{Z}^{2}}{n\lambda}\right),$ (A.38)
while the source condition $F^{\ast}\in[G]^{\beta}$ with $\beta\in(1,2]$ yields the bias bound
$\left\|B_{\mathrm{bias}}\,C_{ZZ}^{1/2}\right\|_{\textrm{HS}}^{2}=\left\|F_{\lambda}-F^{\ast}\right\|_{L^{2}(\pi_{Z};\mathcal{H}_{Y})}^{2}\leq C\,\lambda^{\beta}.$ (A.39)
Combining (A.38) and (A.39) gives
$R=\left\|B_{U}\,C_{ZZ}^{1/2}\right\|_{\textrm{HS}}^{2}\leq C\left\{\lambda^{\beta}+\frac{N_{Z}(\lambda)}{n}+\frac{1}{n^{2}\lambda}\right\}.$ (A.40)
Under the capacity condition $N_{Z}(\lambda)\leq Q\,\lambda^{-p}$ with $p\in(0,1]$, balancing $\lambda^{\beta}$ and $N_{Z}(\lambda)/n$ gives $\lambda_{n}\asymp n^{-1/(\beta+p)}$, hence
$R=O_{p}\left(n^{-\beta/(\beta+p)}\right),\qquad\frac{1}{n^{2}\lambda_{n}}=n^{-2+1/(\beta+p)}=o\left(n^{-\beta/(\beta+p)}\right),$
since $\beta>1$ implies $2+1/(\beta+p)<-\beta/(\beta+p)$.
Next, using (A.38) together with the interpolation bound $\left\|B_{\mathrm{bias}}\right\|_{\textrm{HS}}^{2}\leq C\,\lambda^{\beta-1}$ (valid for $\beta>1$),
$\lambda\|B_{U}\|_{\textrm{HS}}^{2}\leq C\left(\frac{\lambda N_{Z}(\lambda)}{n}+\frac{\lambda^{\beta}}{n}+\frac{1}{n^{2}}\right)+C\,\lambda^{\beta}=o_{p}\left(\lambda^{\beta}+\frac{N_{Z}(\lambda)}{n}\right)=o_{p}(R),$
where the last equality uses (A.40). Combining (A.37) with (A.36) and using (A.40), fix $\varepsilon\in(0,1)$ and set $\tau_{\varepsilon}\coloneqq\log(c/\varepsilon)$. Define
$\psi_{n}\left(\tau_{\varepsilon}\right)\coloneqq C\left(\sqrt{\frac{N_{Z}(\lambda)\,\tau_{\varepsilon}}{n}}+\frac{K_{Z}^{2}\,\tau_{\varepsilon}}{\lambda n}\right),\qquad\Delta_{n}\coloneqq\lambda\|B_{U}\|_{\textrm{HS}}^{2}.$
Then, for all $n$, the event
$\mathcal{E}_{n}\coloneqq\left\{\,\left|R_{n}-R\right|\leq\psi_{n}\left(\tau_{\varepsilon}\right)\left(R+\Delta_{n}\right)\,\right\}$
satisfies $\mathbb{P}(\mathcal{E}_{n})\geq 1-\varepsilon$. With $\lambda=\lambda_{n}\asymp n^{-1/(\beta+p)}$ one has $\psi_{n}(\tau_{\varepsilon})\to 0$ and $\Delta_{n}=o_{p}(R)$. Hence, for any fixed $\eta\in(0,1)$,
$\mathbb{P}\left(\frac{|R_{n}-R|}{R}>\eta\right)\leq\underbrace{\mathbb{P}\left(\mathcal{E}_{n}^{\mathrm{c}}\right)}_{\leq\,\varepsilon}+\mathbb{P}\left(\psi_{n}(\tau_{\varepsilon})>\frac{\eta}{2}\right)+\mathbb{P}\left(\frac{\Delta_{n}}{R}>\frac{\eta}{2}\right)\xrightarrow[n\to\infty]{}\varepsilon.$
Since $\varepsilon\in(0,1)$ is arbitrary, it follows that $|R_{n}-R|/R\to 0$ in probability, i.e.
$\left|R_{n}-R\right|=o_{p}(R)\quad\text{and}\quad R_{n}=R+o_{p}(R)=O_{p}\left(n^{-\beta/(\beta+p)}\right).$
Appendix F Power of Regression-Based CI Tests
Assume that the random vector $(X,Y,Z)$ takes values in $\mathbb{R}^{d_{X}}\times\mathbb{R}^{d_{Y}}\times\mathbb{R}^{d_{Z}}$. Let $\mathcal{E}_{XYZ}$ be the set of all such distributions.
The general alternative space $\mathcal{P}_{1}$, representing conditional dependence, is consequently the set of distributions where this condition is not met. That is, $\mathcal{P}_{1}$ consists of distributions for which
$\operatorname{\mathbb{E}}\left[\left(f(X,Z)-\operatorname{\mathbb{E}}[f(X,Z)\mid Z]\right)\left(g(Y)-\operatorname{\mathbb{E}}[g(Y)\mid Z]\right)\mid Z\right]\neq 0\quad\text{a.s.-}P_{Z}$
holds for some functions $f\in L^{2}_{XZ}$ and $g\in L^{2}_{Y}$. In this work, we study the extent to which three specific regression-based CI tests can reject distributions in $\mathcal{P}_{1}$ with non-trivial power.
It is worth noting that this regression-based perspective is quite general. For instance, kernel-based CI tests, which estimate CMEs in a RKHS, can also be formulated as regression problems. Viewed this way, these kernel methods can be understood as generalizations of methods such as the Generalized Covariance Measure (GCM) and the Weighted Generalized Covariance Measure (WGCM). Specifically, the kernel $k_{\mathcal{Z}}$ associated with the conditioning variable $Z$ in the CME approach serves a role analogous to the weight function in WGCM. A detailed exploration of this connection, however, is omitted here, as our investigation focuses on regression-based CI tests that do not explicitly utilize the kernel framework.
F.1 Three Population Functionals
Here, we define the population versions of the statistics employed by these three tests. We begin by defining $\gamma(Q)$:
$\gamma(Q)\coloneqq\operatorname{\mathbb{E}}_{P}\left[\operatorname{Cov}(X,Y\mid Z)\right]=\operatorname{\mathbb{E}}[(Y-\operatorname{\mathbb{E}}[Y\mid Z])\otimes(X-\operatorname{\mathbb{E}}[X\mid Z])],$
whose sample analogue appears in Shah and Peters (2020).
Next, we consider $\Gamma(Q)(z)$:
$\Gamma(Q)(z)\coloneqq\operatorname{Cov}_{Q}(X,Y\mid Z=z),$
estimated by the WGCM of Scheidegger, Hörrmann and Bühlmann (2022); replacing the weight function by a kernel function yields the Kernelized GCM, as described in, e.g., Fernández and Rivera (2024).
Finally, we introduce $\Delta(P)$:
$\Delta(Q)\coloneqq\sup_{f\in L^{2}_{X},g\in L^{2}_{Y}}\left|\operatorname{\mathbb{E}}_{Q}\left[\operatorname{Cov}\bigl(f(X),g(Y)\mid Z\bigr)\right]\right|,$
whose empirical version is implemented in Shi et al. (2021).  [^1] Following Shi et al. (2021), we call the resulting statistic the Maximum GCM (MGCM) to emphasize the maximization over function classes. The relationships among the null and alternative hypotheses for these three tests are summarized in Table 5.
Table 5: Regression-based tests and their detectable alternatives (all $\mathcal{H}_{1}^{\mathrm{T}}\subset\mathcal{P}_{1}$). The symbol $\equiv$ denotes equality almost surely (a.s.), and $\not\equiv$ indicates that the values differ with positive probability.
Test Functional Null class $\mathcal{H}_{0}^{\mathrm{T}}$ Alternative $\mathcal{H}_{1}^{\mathrm{T}}$
GCM (Shah and Peters, 2020) $\gamma$ $\gamma(Q)=0$ $\gamma(Q)\neq 0$
WGCM (Scheidegger, Hörrmann and Bühlmann, 2022) $\Gamma$ $\Gamma(Q)\equiv 0$ $\Gamma(Q)\not\equiv 0$
MGCM (Shi et al., 2021) $\Delta$ $\Delta(Q)=0$ $\Delta(Q)\neq 0$
Each test’s null class, $\mathcal{H}_{0}^{\mathrm{T}}$, is defined by the condition under which its corresponding population functional is zero. The null class for MGCM, characterized by $\Delta(Q)=0$, is known as weak conditional independence (Daudin, 1980). The other nulls also represent strictly weaker conditions. Specifically, the null for WGCM requires $\Gamma(Q)\equiv 0$ almost surely, while the null for GCM only requires that the conditional covariance is zero on average, $\gamma(Q)=0$.
Consequently, the alternative hypothesis classes $\mathcal{H}_{1}$ detectable by WGCM and MGCM are strictly broader than that of GCM. This hierarchy is given by:
$\mathcal{H}_{1}^{\text{GCM}}\subsetneq\mathcal{H}_{1}^{\text{WGCM}}\kern 5.0pt\text{and}\kern 5.0pt\mathcal{H}_{1}^{\text{GCM}}\subsetneq\mathcal{H}_{1}^{\text{MGCM}}$
However, the functionals $\Gamma$ and $\Delta$ are incomparable, which means that neither of their corresponding alternative classes is a subset of the other. The following examples are chosen to illustrate this incomparability.
F.2 Illustrative examples
Example F.1 (Even-moment dependence: WGCM powerless, MGCM powerful).
Take $Z=c$ (constant), $X\sim\mathcal{N}(0,1)$ and $Y=X^{2}-1$.
$\Gamma(Q)(c)=\operatorname{Cov}(X,Y\mid Z=c)=\operatorname{\mathbb{E}}[XY]-\operatorname{\mathbb{E}}[X]\operatorname{\mathbb{E}}[Y]=\operatorname{\mathbb{E}}[X(X^{2}-1)]-0\times 0=0,$
so $\Gamma(Q)\equiv 0$ and the WGCM is powerless against this alternative. Choose $f(x)=x^{2}$ and $g(y)=y$; then
$\operatorname{Cov}(f(X),g(Y)\mid Z=c)=\operatorname{\mathbb{E}}[(X^{2}-1)(Y-0)]=\operatorname{\mathbb{E}}[(X^{2}-1)^{2}]=\operatorname{Var}(X^{2})=3-1=2,$
so $\Delta(Q)\geq 2>0$ and MGCM does have power.
Example F.2 (Binary XOR dependence: MGCM powerless, WGCM powerful).
Let $X,Y,Z\in\{0,1\}$ and the conditional probabilities $\mathbb{P}(X=x,Y=y\mid Z=z)$ for $x,y,z\in\{0,1\}$ are given by:
[…]
As discussed later, the binary nature of $X$ and $Y$ in this example allows WGCM to guarantee power. In contrast, MGCM cannot do so for this data-generating process, as shown by Shi et al. (2021).
F.3 Special Strength of WGCM for binary $X,Y$
If $X$ and $Y$ are binary (or more generally categorical), $\Gamma(Q)\equiv 0$ is equivalent to CI (Scheidegger, Hörrmann and Bühlmann, 2022, Section 1.1.1). Hence
$\mathcal{H}_{1}^{\text{WGCM}}=\mathcal{P}_{1}\quad\text{when }X,Y\text{ are binary},$
so WGCM enjoys full asymptotic power for this important case.
In summary, GCM is simple but blind to alternatives with $\operatorname{\mathbb{E}}[\operatorname{Cov}(X,Y\mid Z)]=0$. WGCM and MGCM extend power in distinct and complementary directions: WGCM thrives when conditional covariance cancels on average (and dominates for binary data), while MGCM uncovers higher-order nonlinear dependence. Neither test alone exhausts the alternative space $\mathcal{P}_{1}$, underscoring the value of combining weighting and maximization strategies in regression-based CI testing.
F.4 In our simulation settings
We work under the low-dimensional designs in Section 6.11: let $Z\sim\mathcal{N}(0,1)$ and $f_{a}(z)=\exp\left(-z^{2}/2\right)\sin\left(az\right)$ for $a\in\{2,4,6\}$. The baseline model is
$X=f_{a}(Z)+0.3\,\varepsilon_{X},\qquad Y=f_{a}(Z)+0.3\,\varepsilon_{Y},$
where the noises $\varepsilon_{X},\varepsilon_{Y}$ are mean zero, independent of each other and of $Z$. Throughout this subsection, we study
$\Gamma(Q)(z)\coloneqq\operatorname{Cov}\left(X,Y\mid Z=z\right),\qquad\gamma(Q)\coloneqq\operatorname{\mathbb{E}}\left[\,\Gamma(Q)(Z)\right],$
in order to explain, case by case, why GCM (which targets $\gamma$) and WGCM (which targets $\Gamma$) do or do not enjoy guaranteed power under our alternatives.
Baseline (null). For each fixed $z$,
$\Gamma(Q)(z)=\operatorname{Cov}\left(f_{a}(z)+0.3\varepsilon_{X},f_{a}(z)+0.3\varepsilon_{Y}\middle|Z=z\right)=0.09\operatorname{Cov}\left(\varepsilon_{X},\varepsilon_{Y}\mid Z=z\right)=0,$
hence $\gamma(Q)=\operatorname{\mathbb{E}}[\Gamma(Q)(Z)]=0$. Both GCM and WGCM correctly retain the null.
DPG1-1 (linear perturbation). Replace $Y$ by $Y^{\mathrm{new}}=Y+0.2\,X$. Then
$Y^{\mathrm{new}}=(1+0.2)\,f_{a}(Z)+0.3\,\varepsilon_{Y}+0.06\,\varepsilon_{X}.$
For each $z$,
$\Gamma(Q)(z)$ $=\operatorname{Cov}\left(f_{a}(z)+0.3\,\varepsilon_{X},\ (1+0.2)\,f_{a}(z)+0.3\,\varepsilon_{Y}+0.06\,\varepsilon_{X}\ \middle|\ Z=z\right)$
[…]
so that $\gamma(Q)=\operatorname{\mathbb{E}}[\Gamma(Q)(Z)]=0.018\,\operatorname{Var}(\varepsilon_{X})>0$. Hence both WGCM and GCM have nontrivial power.
DPG1-2 (latent even-odd structure; no guarantees for GCM/WGCM). Introduce a latent $U\sim\mathcal{N}(0,1)$, independent of $Z,\varepsilon_{X},\varepsilon_{Y}$, and define
$X^{\mathrm{new}}=f_{a}(Z)+0.3\,\varepsilon_{X}+0.3\left(U^{2}-1\right),\qquad Y^{\mathrm{new}}=f_{a}(Z)+0.3\,\varepsilon_{Y}+0.3\,U.$
For each $z$,
$\Gamma(Q)(z)$ $=\operatorname{Cov}\left(0.3\,\varepsilon_{X}+0.3\left(U^{2}-1\right),\ 0.3\,\varepsilon_{Y}+0.3\,U\ \middle|\ Z=z\right)$
$=0.09\,\operatorname{Cov}\left(U^{2}-1,\,U\mid Z=z\right).$
Because $U\mid Z=z\sim\mathcal{N}(0,1)$, we have $\operatorname{\mathbb{E}}[U\mid Z=z]=0$ and $\operatorname{\mathbb{E}}[U^{3}\mid Z=z]=0$, whence $\operatorname{Cov}\left(U^{2}-1,\,U\mid Z=z\right)=0$. Thus $\Gamma(Q)(z)=0$ for all $z$ and $\gamma(Q)=0$. Neither WGCM nor GCM can guarantee power against this alternative in our construction. Intuitively, the even component $U^{2}-1$ in $X$ is orthogonal to the odd component $U$ in $Y$ by symmetry.
DPG1-3 ($Z$-coupled latent; WGCM retains guarantees, GCM does not). As in Case 2, but now let $U=f_{1}(Z)+\varepsilon_{U}$ with $\varepsilon_{U}\sim\mathcal{N}(0,1)$ independent of $Z,\varepsilon_{X},\varepsilon_{Y}$. For $z\in\mathbb{R}$ write $\mu=f_{1}(z)$. Using
$\operatorname{\mathbb{E}}[U\mid Z=z]=\mu,\quad\operatorname{\mathbb{E}}[U^{2}\mid Z=z]=\mu^{2}+1,\quad\operatorname{\mathbb{E}}[U^{3}\mid Z=z]=\mu^{3}+3\mu,$
we obtain
$\operatorname{Cov}(U^{2}-1,\,U\mid Z=z)$ $=\operatorname{\mathbb{E}}[(U^{2}-1)U\mid Z=z]-\operatorname{\mathbb{E}}[U^{2}-1\mid Z=z]\operatorname{\mathbb{E}}[U\mid Z=z]$
$=(\mu^{3}+2\mu)-\mu^{3}=2\mu.$
Therefore
$\Gamma(Q)(z)=\operatorname{Cov}\left(X^{\mathrm{new}},Y^{\mathrm{new}}\mid Z=z\right)=0.09\cdot 2\,\mu=0.18\,f_{1}(z),$
which is nonzero on a set of $z$ of positive probability. Consequently WGCM, which targets $\Gamma$, has asymptotic power. In contrast,
$\gamma(Q)=\operatorname{\mathbb{E}}[\Gamma(Q)(Z)]=0.18\,\operatorname{\mathbb{E}}\left[f_{1}(Z)\right]=0,$
since $f_{1}(z)=\exp\left(-z^{2}/2\right)\sin(z)$ is odd and $Z$ is symmetric. Hence GCM, which aggregates $\Gamma$ by averaging over $Z$, does not admit a power guarantee here.
Summary. Case 1: $\Gamma>0$ and $\gamma>0$ so both WGCM and GCM have power. Case 2: $\Gamma\equiv 0$ and $\gamma=0$ so neither has a guarantee. Case 3: $\Gamma\not\equiv 0$ but $\gamma=0$, hence WGCM retains a guarantee whereas GCM does not.
Appendix G Explicit constructions of the empirical eigenfunctions
In this section, we describe an explicit construction of the empirical eigenfunctions $\hat{e}_{p}$ and $\hat{f}_{q}$. Since the arguments for $X$ and $Y$ are identical, we focus on $\hat{C}_{XX}$ and $\hat{e}_{p}$. Let $K_{X}\in\mathbb{R}^{n_{2}\times n_{2}}$ be the Gram matrix with entries $K_{X,ij}=k_{\mathcal{X}}(X_{i},X_{j})$. For any eigenpair $(\kappa_{p},u_{p})$ of $K_{X}$ with $\kappa_{p}>0$ and $\|u_{p}\|_{2}=1$, the corresponding normalized empirical eigenfunction is given explicitly by
$\hat{e}_{p}(\cdot)=\frac{1}{\sqrt{\kappa_{p}}}\sum_{i\in I_{2}}u_{p}[i]\;k_{\mathcal{X}}(X_{i},\cdot),$
with corresponding eigenvalue $\hat{\lambda}_{p}=\kappa_{p}/n_{2}$. The same construction applied to the Gram matrix of $\{Y_{i}\}_{i\in I_{2}}$ gives the explicit form of $\hat{f}_{q}$.
We now briefly justify this representation. By definition,
$\hat{C}_{XX}=\frac{1}{n_{2}}\sum_{i\in I_{2}}k_{\mathcal{X}}(X_{i},\cdot)\otimes k_{\mathcal{X}}(X_{i},\cdot),$
so every element in the range of $\hat{C}_{XX}$ is a finite linear combination of $\{k_{\mathcal{X}}(X_{i},\cdot):i\in I_{2}\}$. In particular, any eigenfunction $\hat{e}_{p}$ with nonzero eigenvalue $\hat{\lambda}_{p}$ can be written as
$\hat{e}_{p}(\cdot)=\sum_{i\in I_{2}}\alpha_{p}[i]\;k_{\mathcal{X}}(X_{i},\cdot)$
for some coefficient vector $\alpha_{p}\in\mathbb{R}^{n_{2}}$.
Substituting this representation into the eigenvalue equation $\hat{C}_{XX}\hat{e}_{p}=\hat{\lambda}_{p}\hat{e}_{p}$ and using the reproducing property of the kernel, one obtains the finite-dimensional eigenproblem
$K_{X}\,\alpha_{p}=n_{2}\,\hat{\lambda}_{p}\,\alpha_{p}.$
Thus $\alpha_{p}$ must be an eigenvector of $K_{X}$. Writing $\kappa_{p}>0$ for an eigenvalue of $K_{X}$ and $u_{p}$ for a corresponding unit eigenvector, the choice $\alpha_{p}=u_{p}/\sqrt{\kappa_{p}}$ yields the normalized eigenfunction
$\hat{e}_{p}(\cdot)=\frac{1}{\sqrt{\kappa_{p}}}\sum_{i\in I_{2}}u_{p}[i]\;k_{\mathcal{X}}(X_{i},\cdot),$
with eigenvalue $\hat{\lambda}_{p}=\kappa_{p}/n_{2}$, as claimed.
Appendix H Additional Simulations
H.1 Bootstrap Multipliers for a $V$-statistic wild bootstrap
We consider a wild bootstrap for the NCCM statistic with V-statistic form. The multiplier acts on the summands: with i.i.d. multipliers $W_{1}^{\ast},\dots,W_{n}^{\ast}$ independent of the data,
$T_{n}^{\ast}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}W_{i}^{\ast}W_{j}^{\ast}\langle u_{i},u_{j}\rangle_{\mathcal{H}}.$
We impose only the standard moment conditions
$\operatorname{\mathbb{E}}[W^{\ast}]=0,\qquad\operatorname{\mathbb{E}}\left[(W^{\ast})^{2}\right]=1,\qquad\mu_{4}\coloneqq\operatorname{\mathbb{E}}\left[(W^{\ast})^{4}\right]<\infty,$
which are satisfied by Gaussian $\mathcal{N}(0,1)$, Rademacher $\mathbb{P}(W^{\ast}=\pm 1)=\tfrac{1}{2}$, and the Mammen two-point law
$W^{\ast}=\begin{cases}\dfrac{1-\sqrt{5}}{2},&\text{with probability }\dfrac{\sqrt{5}+1}{2\sqrt{5}},\\ \dfrac{1+\sqrt{5}}{2},&\text{with probability }\dfrac{\sqrt{5}-1}{2\sqrt{5}},\end{cases}$
for which $\operatorname{\mathbb{E}}\left[(W^{\ast})^{3}\right]=1$ and $\operatorname{\mathbb{E}}\left[(W^{\ast})^{4}\right]=2$.
Fix the sample $\{u_{i}\}_{i=1}^{n}\subset\mathcal{H}$. A direct fourth-order moment factorization yields
$\operatorname{\mathbb{E}}\left[T_{n}^{\ast}\mid u_{1:n}\right]$ $=\frac{1}{n^{2}}\sum_{i=1}^{n}\|u_{i}\|_{\mathcal{H}}^{2},$
$\operatorname{\mathbb{E}}\left[(T_{n}^{\ast})^{2}\mid u_{1:n}\right]$ $=\frac{1}{n^{4}}\left\{\left(\sum_{i=1}^{n}\|u_{i}\|_{\mathcal{H}}^{2}\right)^{2}+2\sum_{i=1}^{n}\sum_{j=1}^{n}\left\langle u_{i},u_{j}\right\rangle_{\mathcal{H}}^{2}+\left(\mu_{4}-3\right)\sum_{i=1}^{n}\|u_{i}\|_{\mathcal{H}}^{4}\right\}.$
Writing
$A(u_{1:n})\coloneqq\frac{1}{n^{4}}\left\{\left(\sum_{i=1}^{n}\|u_{i}\|_{\mathcal{H}}^{2}\right)^{2}+2\sum_{i=1}^{n}\sum_{j=1}^{n}\left\langle u_{i},u_{j}\right\rangle_{\mathcal{H}}^{2}\right\},\qquad B(u_{1:n})\coloneqq\frac{1}{n^{4}}\sum_{i=1}^{n}\|u_{i}\|_{\mathcal{H}}^{4}\ \geq 0,$
we have the affine dependence
$\operatorname{\mathbb{E}}\left[(T_{n}^{\ast})^{2}\,\mid\,u_{1:n}\right]=A(u_{1:n})+\left(\mu_{4}-3\right)\,B(u_{1:n}).$
Hence, for fixed data, the conditional second moment (and therefore the conditional variance) of $T_{n}^{\ast}$ is monotone nondecreasing in $\mu_{4}$. Any moment-based upper-tail bound (e.g., Cantelli/Markov) then implies that the bootstrap upper quantiles—and thus the critical values—are nondecreasing in $\mu_{4}$. Since
$\mu_{4}\left(\text{Gaussian}\right)=3>\mu_{4}\left(\text{Mammen}\right)=2>\mu_{4}\left(\text{Rademacher}\right)=1,$
multipliers with larger $\mu_{4}$ are more conservative in finite samples (Gaussian $\succ$ Mammen $\succ$ Rademacher).
H.2 Empirical summary and choice
Table 6 accords with the second-moment prediction. Under the null, Gaussian yields the smallest Type I error across DGPs and $n$, whereas Rademacher can exceed the nominal level in small samples; under alternatives, Rademacher attains the highest power. We therefore adopt Gaussian multipliers to prioritize small-sample level control.
H.3 Other bootstrap routes
As also noted by Lee, Zhang and Shao (2020), alternative bootstrap designs exist (e.g., response-perturbation wild bootstraps such as Patilea, Sánchez-Sellero and Saumard (2016)), but in our setting they require recomputing the full statistic at each replicate, so regressions must be repeatedly refit and the computational cost is high.
Table 6: Empirical rejection rates for SGCM with different bootstrap multipliers $\left(\alpha=0.05\right)$. Gaussian is most conservative; Rademacher attains the highest power.
DGP Multiplier Sample Size $\left(n\right)$
Scenario 100 200 300 400
(a2) Size (Null) Gaussian 0.053 0.044 0.070 0.067
Mammen 0.071 0.053 0.079 0.071
Rademacher 0.092 0.064 0.090 0.074
Power (DGP 1-1) Gaussian 0.212 0.480 0.700 0.820
Mammen 0.244 0.508 0.718 0.836
Rademacher 0.295 0.543 0.738 0.844
Power (DGP 1-2) Gaussian 0.143 0.337 0.490 0.675
Mammen 0.173 0.373 0.526 0.696
Rademacher 0.216 0.418 0.548 0.723
Power (DGP 1-3) Gaussian 0.295 0.691 0.903 0.974
Mammen 0.335 0.714 0.916 0.975
Rademacher 0.394 0.754 0.923 0.978
(a4) Size (Null) Gaussian 0.060 0.071 0.088 0.073
Mammen 0.072 0.082 0.094 0.079
Rademacher 0.098 0.090 0.099 0.087
Power (DGP 1-1) Gaussian 0.280 0.569 0.762 0.868
Mammen 0.327 0.601 0.779 0.875
Rademacher 0.402 0.623 0.798 0.883
Power (DGP 1-2) Gaussian 0.145 0.310 0.470 0.626
Mammen 0.179 0.340 0.503 0.658
Rademacher 0.237 0.378 0.544 0.681
Power (DGP 1-3) Gaussian 0.233 0.527 0.780 0.910
Mammen 0.273 0.559 0.807 0.920
Rademacher 0.335 0.610 0.821 0.934
H.4 Results for DGP (a6)
In addition to the low-dimensional settings (a2) and (a4) in Section 6.1.1, we also conduct simulations for the more challenging case (a6), obtained by the same construction with $a=6$ (thus further increasing the oscillation frequency of $f_{a}$ and the difficulty of the problem).
Table 7 summarizes empirical size and power for DGP (a6). None of the methods attains the nominal size $\alpha=0.05$; inflation is severe for KCI and CDCOV, while GCM, WGCM, and SGCM are comparatively better calibrated, with SGCM closest to the nominal level.
For power, the apparent advantages of GCM and WGCM at $n\in\{100,200\}$ largely reflect their size inflation in this difficult setting. Accounting for calibration, the overall power of SGCM is comparable to that of GCM and WGCM, and by $n=400$ SGCM approaches WGCM. In short, among the relatively better-calibrated procedures, SGCM provides the most balanced size-power trade-off under (a6).
Table 7: Empirical size ($\alpha=0.05$) and power for DGP (a6).
DGP Method 100 200 300 400
Size gcm 0.232 0.158 0.134 0.123
wgcm 0.388 0.250 0.170 0.140
cdcov 1.000 1.000 1.000 1.000
kci 1.000 1.000 1.000 1.000
mmdci 1.000 1.000 1.000 1.000
ccit 0.454 0.444 0.430 0.401
sgcm 0.102 0.095 0.123 0.104
Power (DGP1-1) gcm 0.788 0.911 0.973 0.990
wgcm 0.832 0.852 0.896 0.929
cdcov
kci
mmdci
ccit
sgcm 0.424 0.643 0.829 0.910
[…]
References for Supplementary Material
[…]